### v1

# The Internal Model Principle of Time-Varying Optimization

Abstract
Time-varying optimization problems are central to many engineering applications where performance metrics and system constraints evolve dynamically with time. A number of algorithms have been proposed in recent years to solve such problems; a common feature of all these methods is that they implicitly require precise knowledge of the temporal variability of the solutions in order to exactly track the optimizers. In this paper, we seek to lift these stringent assumptions. Our main result is a fundamental characterization showing that an algorithm can track an optimal trajectory if and only if it contains a model of the temporal variability of the problem. We refer to this concept to as the internal model principle of time-varying optimization. By recasting the optimization objective as a nonlinear regulation problem and using tools from center manifold theory, we provide necessary and sufficient conditions both for an optimization algorithm to achieve exact asymptotic tracking and for such an algorithm to exist. We illustrate the applicability of the approach numerically on both synthetic problems as well as practical problems in transportation.
I Introduction
Time-varying optimization problems play a central role in several scientific domains, as they underpin many important contemporary engineering problems. Examples include training in Machine Learning [1, 2], dynamic signal estimation in Signal Processing [3], trajectory tracking in Robotics [4], system optimization in Industrial Control [5], and much more. Historically, discrete-time algorithms for time-varying optimization have been proposed and studied first since they emerge as a natural extension of their time-invariant counterparts, allowing for cost functions and constraints that may change over time [6]. These approaches build on the classical perspective on mathematical optimization, which seeks to construct methods to determine optimizers and consist of iterative procedures implemented on digital devices. Recently, there has been a renewed interest in the use and development of continuous-time dynamics for optimization purposes, due to the possibility of utilizing tools from dynamical systems for their analysis [7, 8] and motivated by practice, where optimization is increasingly used to control physical systems [9].
Motivated by these recent developments, in this paper we study time-varying convex optimization problems and focus on the use of continuous-time dynamics to track exactly optimal solutions. Although, by now, several approaches have been developed for this purpose [10, 11, 12, 13], all these techniques implicitly require full knowledge of the temporal variability of the problem [13]. Unfortunately, in most practical applications, having such knowledge is impractical, either because the temporal variability enters the optimization in the form of exogenous disturbances that are unknown and cannot be measured (see, e.g., [9, 14]), or simply because it is unrealistic to ask for a noiseless model of how the problem changes with time. Departing from this, in this paper we pose the following question: is it possible to track (exactly and asymptotically) a minimizer of a time-varying optimization problem without any knowledge of the temporal variability of the optimization? Interestingly, we prove a fundamental result showing that tracking can be achieved if and only if the temporal variability of the problem can be ‘observed’ by the algorithm, and the latter incorporates a suitably reduplicated model of such a variability. We refer to this conclusion as the internal model principle of time-varying optimization, akin to its control-theoretic counterpart [15, 16]. Our approach relies on reinterpreting the optimization algorithm design as a nonlinear, multivariable regulation problem [17], and our analysis uses tools from center manifold theory [18, 16]. Figure 1 illustrates the architecture of the gradient-feedback design scheme studied in this paper.
Figure 1: Architecture of the gradient-feedback design scheme studied in this work. An optimization algorithm is to be designed (bottom block), having access only to gradient evaluations of the loss function to be minimized (top right block), and generating a sequence of exploration points $x(t)$ at which the gradient shall be evaluated. The loss function to be optimized varies with time, where the temporal variability $\theta(t)$ is assumed to be unmeasurable and generated by an exosystem (top left block). Shaded blocks emphasize the presence of dynamics.
Related works. The literature on methods for time-varying optimization is mainly divided into two classes of solutions. The first class consists of methods that do not utilize any model of the temporal variability of the problem [19, 20, 21, 22]; instead, they seek to solve a sequence of static problems. Several established approaches belong to this class, including the online gradient descent method [23] and the online Newton step algorithm [24] – see [25] and references therein. Because the temporal variability of the problem is unknown (or ignored), algorithms in this class can make a decision only after each variation has been observed, thus incurring a certain ‘regret.’ Mathematically, these techniques are capable of reaching only a neighborhood of an optimizer [13], and exact tracking is out of reach in general for these approaches. In contrast, the second class of methods uses a model of the temporal evolution of the problem to seek to exactly track optimal trajectories [26, 2, 27, 28]. Particularly celebrated is the prediction-correction algorithm (see [10] and the recent work [11]), whereby at each time a prediction step is used to anticipate how the optimizer will evolve over time, and a correction step is used to seek a solution to each instantaneous optimization problem. We refer to [13] for a recent overview of the topic. Recent years have witnessed a growing interest in this problem: [4] uses contraction to study these methods; [29] uses sampling to estimate the temporal variability of the problem; a recent survey on analyzing optimization algorithms using tools from control has appeared in [30]; constrained optimization problems are studied in [31]. Despite these efforts, the fundamental question of what is the basic algorithm structure that enables exact asymptotic tracking remains unanswered.
Contributions. This paper makes four main contributions. First, we recast the problem of designing a time-varying optimization algorithm as a nonlinear multivariable regulation problem, whereby an algorithm is to be designed to regulate the gradient of a certain function to zero. We leverage this formulation to characterize the class of optimization algorithms that can achieve tracking. In a net departure from existing approaches (e.g., [23, 13, 11, 12]), our characterization is general and allows us to study not only a single optimization method, but an entire class, which enables us to derive fundamental results for all algorithms in the class. Second, by harnessing tools from center manifold theory [18, 16], we provide necessary and sufficient conditions for an optimization algorithm to ensure tracking. Interestingly, these conditions depend on the properties of the loss function (through a gradient invertibility-type condition) and on the inner model describing the temporal variability of the problem. This property allows us to prove the internal model principle of time-varying optimization, which states that for an optimization algorithm to achieve asymptotic tracking, it must incorporate a reduplicated model of the temporal variability of the problem. This feature is implicit in all existing approaches for time-varying optimization [13] but, to the best of the authors’ knowledge, lacked a formal proof until now. Third, we derive necessary and sufficient conditions for the existence of a tracking algorithm. Fourth, we use our characterizations to design algorithms for time-varying optimization. With respect to the existing literature, our algorithm does not require one to know or measure exactly the temporal variability of the problem, and thus uses less stringent assumptions. Finally, we illustrate the applicability of the approach numerically on both synthetic problems as well as practical problems in transportation.
Organization. Section II presents the problem of interest, Section III studies the simpler problem of parameter-feedback optimization, which allows us to derive the necessary framework, and Section IV tackles the time-varying optimization design problem in its full generality and contains our main results. Section V discusses the tracking accuracy in relationship to the fidelity of the internal model, Section VI presents some extensions to constrained optimization, Section VII validates numerically the results, and Section VIII illustrates our conclusions. Finally, in the Appendix we summarize basic facts on center manifold theory that are used extensively throughout the paper.
[…]
II Problem setting
II-A Problem statement
We consider the time-varying optimization problem:
$\min_{x\in{}^{n}}f(x,\theta(t)),$ (1)
where $t\in\mathbb{R}_{\geq 0}$ denotes time and $f:{}^{n}\times\Theta\rightarrow\real,$ $\Theta\subseteq{}^{p},$ is a loss function that is parametrized by the time-varying parameter vector $\theta:\mathbb{R}_{\geq 0}\rightarrow\Theta.$ We make the following assumptions throughout.
Assumption 1 (Properties of the objective function).
The map $x\mapsto f(x,\theta)$ is convex and $x\mapsto\nabla_{x}f(x,\theta)$ is Lipschitz continuous in n, for each $\theta\in\Theta.$square
Assumption 2 (Existence of an exosystem).
There exists a smooth (i.e., $C^{\infty}$) vector field $s:\Theta\rightarrow{}^{p}$ and $\theta(0)\in\Theta$ such that the parameter vector $\theta(t)$ satisfies
$\dot{\theta}(t)=s(\theta(t))$ (2)
for all $t\in\mathbb{R}_{\geq 0}.$square
Convexity and smoothness are standard assumptions in optimization [25]. Assumption 2 ensures the existence of an autonomous system, called exosystem, describing the class of temporal variabilities of the cost taken into consideration.
In what follows, we will say that $x^{\star}:\mathbb{R}_{\geq 0}\rightarrow{}^{n}$ is a critical trajectory of (1) if it satisfies:
$0=\nabla_{x}f(x^{\star}(t),\theta(t)),\quad\forall t\in\mathbb{R}_{\geq 0}.$
We will assume existence of a critical trajectory and that any critical trajectory is continuous.
Remark 1.
The time variation in (1) is captured implicitly through the parameter vector $\theta(t)$. A related, yet slightly more general, problem is as follows:
$\min_{x\in{}^{n}}f_{0}(x,t),$ (3)
where the dependency on time is explicit. Problem (1) can be recast uniquely as in (3) by letting $f_{0}(x,t)=f(x,\theta(t))$ for all $x$ and $t$. On the other hand, in general, there exists an infinite number of ways to parametrize (3) as in (1), thus leading to possible ambiguities. For instance, any function $f_{0}(x,t)$ may be parametrized by $\theta(t)=t$ (so that $f_{0}\equiv f$), although this is not always compatible with our assumptions (see Assumption 3 and Remark 2 below). $\square$
In this work, we are interested in the problem of designing an optimization algorithm that is capable of determining and tracking a critical trajectory of (1). We are interested in doing so with as little knowledge as possible on $\theta(t)$ and its temporal variability. From an optimization perspective, three types of assumptions may be considered (presented in increasing order of restrictiveness):
  1. (A1)
    The parameter signal $\theta(t)$ is known or measurable at each $t\in\mathbb{R}_{\geq 0},$ but no knowledge of $s(\theta)$ is available.
  2. (A2)
    The vector field $s(\theta)$ is known, but $\theta(0)$ (and hence also $\theta(t)$) is unknown.
  3. (A3)
    The vector field $s(\theta)$ as well as $\theta(0)$ are known.
Under these assumptions, different methods have been developed in the literature, each with corresponding guarantees:
  1. (M1)
    Under (A1), all methods from static optimization (e.g., [23, 24, 19, 20, 21, 22, 13]) can be applied directly. All these methods can achieve convergence only to a neighborhood of a critical trajectory.
  2. (M2)
    The approach of [12] is applicable in this case, which is however limited to strongly convex quadratic optimization problems with linear temporal variablities.
  3. (M3)
    The class of prediction correction methods [10] and their variations [11] take advantage of this knowledge and ensures asymptotic tracking of $x^{\star}(t)$.
A natural question emerges directly from the above discussion: is (A3) necessary to ensure exact tracking and, if not, can it be relaxed (e.g., to (A1) or (A2))? With this motivation, in this paper we are driven by the following question.
Problem 0 (Minimal knowledge for exact tracking).
What is the least-restrictive set of assumptions (cf. (A1)–(A3)) under which there exists an optimization algorithm that achieves exact asymptotic tracking of the critical points of (1)?  $\square$
Our objective is to seek an optimization algorithm that assumes no access to $\theta(t)$; instead, as common in first-order optimization approaches [25], we will assume only the availability of functional evaluations of the gradient function $\nabla_{x}f(x,\theta(t))$ at points $x\in{}^{n},$ selected by the algorithm. Formally, the optimization algorithm will be described by an internal state $z(t),$ which takes values on an open subset $\mathcal{Z}\subseteq{}^{n_{c}},n_{c}\in\mathbb{N}_{>0}$; the optimization algorithm generates a sequence of points $x(t)\in{}^{n}$ (called exploration signal) at which the gradient shall be evaluated, and processes functional evaluations of the gradient at these points, $y(t)=\nabla_{x}f(x(t),\theta(t))$ (called gradient feedback signal). Mathematically, the optimization algorithm is:
$\dot{z}(t)$ $=F_{c}(z(t),y(t)),$ $x(t)$ $=G_{c}(z(t)),$ (4a)
together with the gradient-feedback signal:
$y(t)=\nabla_{x}f(x(t),\theta(t)),$ (4b)
where $F_{c}:\mathcal{Z}\times{}^{n}\rightarrow{}^{n_{c}}$ and $G_{c}:\mathcal{Z}\rightarrow{}^{n}$ are functions to be designed. In the remainder, we refer to (4) as a dynamic gradient-feedback optimization algorithm. The architecture of the studied gradient-feedback algorithm is illustrated in Figure 1. Notice that the dynamics of the optimization algorithm (4), coupled with the time-variability generator (2), have the form of a nonlinear autonomous system:
$\dot{z}(t)$ $=F_{c}(z(t),y(t)),$
$y(t)$ $=\nabla_{x}f(G_{c}(z(t)),\theta(t)),$
$\dot{\theta}(t)$ $=s(\theta(t)).$ (5)
Our objective is to design $F_{c}(z,y),$ $G_{c}(z),$ and $n_{c},$ so that $y(t)\rightarrow 0$ as $t\rightarrow\infty,$ which ensures that $x(t)$ tracks, with zero asymptotic error, a critical trajectory $x^{\star}(t)$ of (1).
We will assume that $\theta(t)$ takes values in a neighborhood of the origin, and thus let $\Theta$ be some neighborhood of the origin of ${}^{p}.$ Note that there is no loss of generality in doing so, because if $\theta(t)$ takes values in the neighborhood of any other point, the former can be shifted to the origin via a change of variables without altering the critical points of (1). In what follows, we will denote by $x^{\star}_{\circ}\in{}^{n}$ a point such that
$0=\nabla_{x}f(x^{\star}_{\circ},0),$ (6)
and assume that $x^{\star}_{\circ}\in{}^{n}$ is locally unique. Moreover, we will let $z^{*}_{\circ}$ be such that $x_{\circ}^{\star}=G_{c}(z^{\star}_{\circ}),$ and assume that $z^{*}_{\circ}$ exists and is locally unique.
Definition 1.
We say that (4) asymptotically tracks a critical trajectory of (1) if, for each initial condition $(z(0),\theta(0))$ in some neighborhood of $(z_{\circ}^{*},0)\in\mathcal{Z}\times\Theta$, the solution of (II-A) satisfies $y(t)\to 0$ as $t\to\infty$. $\square$
Notice that, when an algorithm asymptotically tracks a critical trajectory, $x(t)\rightarrow x^{\star}(t)$ for some critical trajectory $x^{\star}(t);$ namely, critical point is readily given by the exploration signal. We make the objective of this work formal as follows.
Problem 1 (Dynamic gradient-feedback problem).
Find necessary and sufficient conditions (in terms of the loss function $f$) for the existence of a gradient-feedback optimization algorithm that asymptotically tracks a critical trajectory of (1). Derive a method to design such an algorithm when these conditions hold. $\square$
It is important to note that existence conditions for $F_{c}(z,y)$ and $G_{c}(z)$ shall not depend on the temporal variability of the optimization (namely, $s(\theta)$), but only on the properties of the optimization problem. Conversely, if such conditions were to depend on $s(\theta)$, a more general class of optimization algorithms than (4) could be constructed, implying the formulation (4) would not be sufficiently general.
II-B Standing assumptions
We list hereafter the basic assumptions on which our approach to gradient-feedback theory is based.
Assumption 3 (Stability of the exosystem).
The point $\theta=0$ is a stable equilibrium of (2), and there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of 00 with the property that each initial condition $\theta(0)\in\Theta_{\circ}$ is Poisson stable [^1] Recall that, for a nonlinear system of the form (2), an initial condition $\theta^{\circ}$ is said to be Poisson stable if the solution starting from $\theta^{\circ}$ at time $t$ of (2) (here denoted by $\theta_{t}\left(\theta^{\circ}\right)$)) is defined for all $t\in\mathbb{R}_{\geq 0}$ and, for every neighborhood $U$ of $\theta^{\circ}$ and each $T\in\mathbb{R}_{>0}$, there exists a time $t_{1}>T$ such that $\theta_{t_{1}}(\theta^{\circ})\in U$, and a time $t_{2}<-T$ such that $\theta_{t_{2}}(\theta^{\circ})\in U$.. $\square$
[…]
For convenience, we will require that $F_{c}(z^{\star}_{\circ},0)=0.$ This ensures that the optimization algorithm (4) has an equilibrium at $z=z^{\star}_{\circ},$ and the corresponding gradient feedback signal is identically zero: $y(t)=\nabla_{x}f(x_{\circ}^{\star},0)=0.$ Since $\theta(t)$ is unmeasurable, the temporal variability of the cost can only be evaluated through measurements of $y(t);$ hence, we make the following assumption.
Assumption 4 (Detectability of the exosystem).
Let
$Q:=\left[\frac{\partial\nabla_{x}f}{\partial\theta}\right]_{(x,\theta)=(x_{\circ}^{\star},0)},$ $S:=\left[\frac{\partial s}{\partial\theta}\right]_{\theta=0}.$ (7)
The pair $(Q,S)$ is detectable. $\square$
Detectability ensures that all the modes of the exosystem (2) can be reconstructed from the available measurements $y(t).$ Undetectability of $(Q,S)$ corresponds to a redundant description of the exogenous signal: indeed, if some modes of the exosystem do not influence the gradient, then $y(t)$ would be independent of those modes and they can thus be removed without altering the problem.
Example 1 (Detectability).
Suppose the objective function admits the decomposition $f(x,\theta(t))=\hat{f}(x)+x^{\mathsf{T}}\theta_{1}(t),$ where $\theta_{1}(t)=\alpha\cos(\omega t)$ for some $\alpha$ and $\omega$. The parameter signal can be generated by (2) with initial condition $\theta(0)=(\alpha,0)$ and vector field:
$s(\theta)=\begin{bmatrix}0&1\\ -\omega^{2}&0\end{bmatrix}\theta.$
Notice that, although the gradient signal $y(t)=\nabla_{x}\hat{f}(x)+\theta_{1}(t)$ is independent of $\theta_{2}(t),$ both modes of the exosystem affect the gradient (indeed, $\theta_{1}(t)=\alpha\cos(\omega t)$ cannot be generated by a linear system with a single mode). Notice also that the pair $(Q,S)$ is detectable. $\square$
III The parameter-feedback problem
To address Problem 1, we first study a simpler problem that allows us to derive the necessary framework to tackle our objectives in their generality. To this end, in place of the dynamic optimization algorithm (4), we begin by considering an algebraic optimization algorithm of the form:
$x(t)=H_{c}(\theta(t)),$ (8)
where $H_{c}:\Theta\rightarrow{}^{n}$ is a mapping to be designed; we will require that $H_{c}$ is of class $C^{0}$ and satisfies the fixed-point condition $x^{\star}_{\circ}=H_{c}(0).$ Because of the explicit dependence on $\theta(t),$ we will refer to (8) to as a parameter-feedback optimization algorithm. Our objective is to design the map $H_{c}$ so that the composition of (2), (4b), and (8):
$y(t)$ $=\nabla_{x}f(H_{c}(\theta(t)),\theta(t)),$
$\dot{\theta}(t)$ $=s(\theta(t)),$ (9)
tracks, with zero asymptotic error, the critical trajectories of (1). For the framework considered in this section, Problem 1 is reformulated as follows.
Problem 2 (Static parameter-feedback problem).
Find necessary and sufficient conditions (in terms of the loss function $f$) for the existence of a parameter-feedback optimization algorithm that asymptotically tracks a critical trajectory of (1). Derive a method to design such an algorithm when these conditions hold. $\square$
Solvability of the static parameter-feedback problem will depend on the existence of a function that zeros the gradient.
Definition 2 (Mapping zeroing the gradient).
We say that a mapping $H_{c}:\Theta\to{}^{n}$ zeros the gradient at the point $\theta\in\Theta$ if
00 0 $=\nabla_{x}f(H_{c}(\theta),\theta).$ (10)
Moreover, we say that $H_{c}$ zeros the gradient globally if (10) holds for all $\theta\in\Theta$, and locally if (10) holds for all $\theta\in\Theta_{\circ},$ where $\Theta_{\circ}\subseteq{}^{p}$ is some neighborhood of the origin [^2] The neighborhoods in Defn. 2 and Assumption 3 need not be the same.. $\square$
The following result characterizes all parameter-feedback optimization algorithms that achieve asymptotic tracking of a critical trajectory.
Theorem 1 (Parameter-feedback algorithm characterization).
Let Assumptions 1, 2, and 3 hold. The parameter-feedback algorithm (III) asymptotically tracks a critical trajectory of (1) if and only if the mapping $H_{c}$ zeros the gradient locally.
Proof.
(Only if) Suppose $y(t)\to 0$ as $t\to\infty$. We will show that $H_{c}$ zeros the gradient locally. By Poisson stability of the exosystem (cf. Assumption 3), there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin such that, for every $\theta_{\circ}\in\Theta_{\circ},$ every $\varepsilon>0,$ and every $T>0,$ the trajectory of (III) satisfies
$\|\theta(t)-\theta_{\circ}\|<\varepsilon,$
at some $t>T$. Then, $\lim_{t\rightarrow\infty}y(t)=0$ can hold only if (10) holds for all $\theta\in\Theta_{\circ}$.
(If) Suppose $H_{c}$ zeros the gradient on some neighborhood $\Theta_{0}\subset\Theta$ of the origin, i.e., (10) holds for all $\theta\in\Theta_{0}$. By Assumption 3, the point $\theta=0$ is a stable equilibrium of the exosystem, so there exists some other neighborhood $\Theta_{1}\subset\Theta$ of the origin and time $T>0$ such that $\theta(t)\in\Theta_{0}$ for all $t>T$ when $\theta(0)\in\Theta_{1}$. Since $H_{c}$ zeros the gradient on $\Theta_{0}$, this implies that $y(t)\to\infty$ as $t\to\infty$. ∎
Remark 2.
Poisson stability of the exosystem is needed for the gradient zeroing condition to be necessary for asymptotic tracking of a critical trajectory (but is not needed for sufficiency). This is because Poisson stability implies that all points sufficiently close to the origin are revisited infinitely often, so $H_{c}$ must zero the gradient on a neighborhood of the origin. More generally, $H_{c}$ must zero the gradient on all limit points of $\theta(t)$$\square$
The result provides a full characterization of all parameter-feedback algorithms that ensure exact asymptotic tracking. In words, $x=H_{c}(\theta)$ is a parameter-feedback algorithm if and only if it zeros the gradient everywhere in a neighborhood of the critical point. As a byproduct, the result also provides a necessary and sufficient condition for the solvability of the parameter-feedback problem: the problem is solvable if and only if the set of solutions to the system of equations $0=\nabla_{x}f(x,\theta)$ can be expressed, everywhere in a neighborhood of the origin of $\Theta$, as the graph of a function $x=H_{c}(\theta).$ Finally, the theorem provides an explicit form for the desired parameter-feedback algorithm (8): this is given by $x(t)=H_{c}(\theta(t));$ this algorithm ensures that, for all initial states $\theta(0)$ sufficiently close to the origin, $y(t)\rightarrow 0$ as $t\rightarrow\infty.$ In this sense, Theorem 1 provides a complete answer to Problem 2. We illustrate the applicability of the result and the necessity of the provided condition in the following example.
Figure 2: Investigation of the condition (10). (Left) Loss function $f(x,\theta)$ studied in Example 2, plotted for $\theta=0.$ (Right) Gradient $\nabla f(x,\theta).$ The function $f(x,0)$ admits two critical points: $x^{\star}_{\circ,1}=\frac{1}{\sqrt{3}}$ and $x^{\star}_{\circ,2}=-\frac{2}{\sqrt{3}}.$ At $x^{\star}_{\circ,1},$ condition (10) is not satisfied, since with an an upward shift of the graph of $\nabla_{x}f(x,0),$ $x^{\star}_{\circ,1}$ is no longer a critical point of $f(x,0).$ On the other hand, (10) holds for $x^{\star}_{\circ,2},$ since $x^{\star}_{\circ,2}$ varies continuously as $\theta$ is perturbed. See Example 2 for a discussion.
Example 2.
Consider an instance of (1) with $n=p=1,$
$f(x,\theta)=(x-1)^{2}(x+1)^{2}+\frac{8}{3\sqrt{3}}x+\theta x.$ (11)
See Fig. 2(a) for an illustration of this function. For $\theta=0,$ the optimization problem associated with (11) admits two critical points: $x^{\star}_{\circ,1}=\frac{1}{\sqrt{3}}$ and $x^{\star}_{\circ,2}=-\frac{2}{\sqrt{3}};$ indeed, it follows from direct inspection that $\nabla_{x}f(x^{\star}_{\circ,1},0)=0$ and $\nabla_{x}f(x^{\star}_{\circ,2},0)=0.$ At the critical point $x^{\star}_{\circ,1},$ a function $H_{c}(\theta)$ as in (10) does not exist. This can be visualized with the aid of Fig. 2(b): if $\theta=0$ is perturbed to $\theta+\epsilon,$ $\epsilon>0,$ the graph of $\nabla_{x}f(x,0)$ (illustrated in Fig. 2(b)) shifts upward and the equation $\nabla_{x}f(x,\epsilon)=0$ no longer admits a solution in a neighborhood of $x^{\star}_{\circ,1}.$
On the other hand, for the critical point $x^{\star}_{\circ,2},$ any arbitrarily-small upward or downward shift of the graph of $\nabla_{x}f(x,0)$ results in a continuous perturbation of $x^{\star}_{\circ,2}$ (see Fig. 2(b)), thus suggesting existence of $x=H_{c}(\theta)$ as in (10). This graphical observation can be formalized with the aid of the implicit function theorem [32], as described next. Define $F(x,\theta):=\nabla_{x}f(x,\theta)$ and notice that $F$ is continuously differentiable with $F(x^{\star}_{\circ,2},0)=0.$ By the implicit function theorem, there exists a neighborhood $\Theta_{\circ}$ of $x^{\star}_{\circ,2}$ and a function $H_{c}:\Theta_{\circ}\rightarrow{}^{n}$ such that $F(H_{c}(\theta),\theta)=0$ in $\Theta_{\circ},$ provided that $\left.\frac{\partial F}{\partial x}\right|_{(x,\theta)=(x^{\star}_{\circ,2},0)}\neq 0.$ By inspection, it is immediate to see that the latter condition is satisfied. $\square$
Existence of a map $x=H_{c}(\theta)$ as in (10) can be ensured for general problems with the aid of the implicit function theorem [32], as illustrated next.
Proposition 2 (Existence of local parameter-feedback algorithms).
Let Assumptions 1, 2, and 3 hold, and let $X_{\circ}\times\Theta_{\circ}$ be some neighborhood of $(x_{\circ}^{\star},0).$ Further, assume that:
  1. (i)
    the loss function $f$ is $C^{1}$ on $X_{\circ}\times\Theta_{\circ},$
  2. (ii)
    $x\mapsto f(x,\theta)$ is $C^{2}$ on $X_{\circ}$ for each $\theta\in\Theta_{\circ},$ and
  3. (iii)
    the Hessian $\left.\nabla^{2}_{xx}f(x,\theta)\right|_{x=x^{\star}_{\circ},\theta=0}$ is positive definite.
Then, there exits a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient locally. $\square$
Proof.
Define $F(x,\theta)=\nabla_{x}f(x,\theta)$ and note that, under the stated assumptions, $F$ is $C^{0}$ on $X_{\circ}\times\Theta_{\circ}$, the mapping $x\mapsto F(x,\theta)$ is $C^{1}$ on $X_{\circ}$ for all $\theta\in\Theta_{\circ}$, and
$\det\left[\frac{\partial F(x,\theta)}{\partial x}\right]_{(x,\theta)=(x_{\circ}^{\star},0)}\neq 0.$
Then, the result follows from the implicit function theorem (see, e.g., [33, Thm. 1]) applied to the first-order optimality conditions $0=F(x,\theta)$ at the point $(x,\theta)=(x^{\star}_{\circ},0)$. ∎
Intuitively, when the loss function is twice continuously differentiable in the decision variable with a positive definite Hessian, every critical point is a local isolated minimum and existence of $H_{c}(\theta(t))$ is guaranteed by the implicit function theorem. Notice that, differently from the characterization in Theorem 1, the conditions offered by Proposition 2 are sufficient but not necessary.
Although the conditions are immediate to verify, the convergence claims of Theorem 1 and Proposition 2 are of local nature, namely, $y(t)\rightarrow 0$ is ensured provided that $\theta(0)$ is sufficiently close to the origin. The following result provides a sufficient condition for global convergence.
Theorem 3 (Existence of global parameter-feedback algorithms).
Let Assumptions 1, 2, and 3 hold. Further, assume:
  1. (i)
    the loss function $f$ is $C^{1}$ on ×${}^{n}\times\Theta$,
  2. (ii)
    $x\mapsto f(x,\theta)$ is $C^{3}$ on n for each $\theta\in\Theta$,
  3. (iii)
    the Hessian $\nabla^{2}_{xx}f(x,\theta)$ is positive definite on ×${}^{n}\times\Theta$,
  4. (iv)
    the mappings $\nabla_{xx}^{2}f$ and $\frac{\partial\nabla_{xx}^{2}f}{\partial x}$ are $C^{0}$ on ×${}^{n}\times\Theta.$
Then, there exits a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient globally. $\square$
Proof.
Define $F(x,\theta)=\nabla_{x}f(x,\theta)$ and note that, under the stated assumptions, $F$ is $C^{0}$ on ×${}^{n}\times\Theta$, the mapping $x\mapsto F(x,\theta)$ is $C^{2}$ on n for all $\theta\in\Theta$,
$\det\left[\frac{\partial F(x,\theta)}{\partial x}\right]\neq 0,\quad\forall(x,\theta)\in{}^{n}\times\Theta,$
and the mappings $\frac{\partial F}{\partial x}$ and $\frac{\partial^{2}F}{\partial x^{2}}$ are $C^{0}$ in ×${}^{n}\times\Theta$. Hence, assumptions (B1)–(B5) of a global version of the implicit function theorem [33, Thm. 2] (see also [34, Thm. 6]) are satisfied for the first-order optimality conditions $0=F(x,\theta)$, and the claim follows. ∎
Theorem 3 shows that, under additional continuity assumptions on the loss function (cf. conditions (i), (ii), and (iv)) and when the Hessian of the loss is positive definite everywhere, convergence of (III) is ensured globally. More formally, under the assumptions of Theorem 3, the parameter-feedback algorithm $x(t)=H_{c}(\theta(t))$ ensures that $y(t)\rightarrow 0$ as $t\rightarrow\infty$ for all initial conditions $\theta(0)\in\Theta$. This result is particularly relevant, as it ensures the existence of parameter-feedback algorithms for the (very important in practice) class of strictly convex loss functions (studied in, e.g., [11, 12]).
We conclude this section by illustrating the design procedure for parameter-feedback algorithms on a quadratic problem.
Example 3.
Consider an instance of (1) with quadratic cost and time-variability that depends linearly on $\theta(t)$ (which has been investigated in [12]):
$f(x(t),\theta(t))=\tfrac{1}{2}x(t)^{\mathsf{T}}Rx(t)+x(t)^{\mathsf{T}}Q\theta(t),$ (12)
with matrices $R\in\mathbb{S}^{n}$ and $Q\in{}^{n\times p}$. In this case, the signal we wish to regulate to zero is: $y(t)=\nabla_{x}f(x(t),\theta(t))=Rx(t)+Q\theta(t).$ For arbitrary $\theta,$ this problem admits a critical point if and only if $\operatorname{Im}Q\subseteq\operatorname{Im}R$, in which case $x_{\circ}^{\star}$ is unique. Applying Theorem 1 amounts to finding a linear transformation $H_{c}\in{}^{n\times p}$ such that $0=(RH_{c}+Q)\theta$ for all $\theta$ in a neighborhood of the origin. Assuming $\operatorname{Im}Q\subseteq\operatorname{Im}R$, we can choose $H_{c}=-R^{\dagger}Q$, where $R^{\dagger}$ is the pseudo-inverse of $R$. Note that, by substituting into (III), we have
$y(t)=RH_{c}\theta(t)+Q\theta(t)=0,\quad\forall t\in\mathbb{R}_{\geq 0}.$
Namely, the gradient is identically zero at all times. We also note that this is a particular feature of the parameter feedback problem, which can be achieved because $\theta(t)$ is assumed to be known when implementing (8). As we will see shortly below (cf. Section IV), $y(t)=0$ can be achieved only as $t\rightarrow\infty$ for the gradient-feedback problem.
We conclude by noting that $y(t)=0$ for all $t\in\mathbb{R}_{\geq 0}$ holds for all initial conditions $\theta(0)$ because the mapping $H_{c}(\theta)$ derived here zeros the gradient globally (and not only in some neighborhood of the origin). See also Theorem 3. $\square$
IV The dynamic gradient-feedback problem
In this section, we will build on the framework derived for the parameter-feedback problem to address the main objectives of this work: Problem 1.
IV-A Fundamental results
We begin with the following instrumental characterization.
Theorem 4 (Gradient-feedback algorithm characterization).
Suppose Assumptions 1, 2, and 3 hold, and assume that $F_{c}(z,y)$ and $G_{c}(z)$ are such that the equilibrium $z=z^{\star}_{\circ}$ of
$\dot{z}(t)$ $=F_{c}(z(t),\nabla_{x}f(G_{c}(z(t)),\theta(t))),$
is exponentially stable. Then, the gradient-feedback optimization algorithm (II-A) asymptotically tracks a critical trajectory of (1) if and only if there exists a $C^{2}$ mapping $z=\sigma(\theta)$ with $\sigma(0)=z_{\circ}^{\star}$, defined in a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin, that satisfies the following for all $\theta\in\Theta_{\circ}$:
$\frac{\partial\sigma(\theta)}{\partial\theta}s(\theta)$ $=F_{c}(\sigma(\theta),0),$ (13a)
00 0 $=\nabla_{x}f(G_{c}(\sigma(\theta)),\theta).$ (13b)
Proof.
(Only if). We first prove that $\lim_{t\rightarrow\infty}y(t)=0$ implies (13). The coupled dynamics (II-A) have the form:
$\dot{z}$ $=(A_{c}+B_{c}RM)z+B_{c}Q\theta+\chi(x,\theta),$
$\dot{\theta}$ $=S\theta+\psi(\theta).$ (14)
for some mappings $\chi(x,\theta)$ and $\psi(\theta)$ that vanish at the origin along with their first-order derivatives, and where $Q$ and $S$ are defined in (7) and
$A_{c}$ $=\left[\frac{\partial F_{c}}{\partial z}\right]_{(z,y)=(z_{\circ}^{\star},0)},$ $B_{c}$ $=\left[\frac{\partial F_{c}}{\partial y}\right]_{(z,y)=(z_{\circ}^{\star},0)},$
$R$ $=\left[\frac{\partial\nabla_{x}f}{\partial x}\right]_{(x,\theta)=(x_{\circ}^{\star},0)},$ $M$ $=\left[\frac{\partial G_{c}}{\partial z}\right]_{z=z_{\circ}^{\star}}.$
By assumption, the eigenvalues of the matrix $A_{c}+B_{c}RM$ are in $\mathbb{C}^{-}$ and those of $S$ on the imaginary axis. By Theorem 8, the system (IV-A) has a center manifold at (0,0): the graph of a mapping $z=\sigma(\theta),$ with $\sigma(\theta)$ satisfying (see (31))
$\frac{\partial\sigma(\theta)}{\partial\theta}s(\theta)=F_{c}(\sigma(\theta),\nabla_{x}f(G_{c}(\sigma(\theta)),\theta)).$
By Assumption 3, no trajectory on this manifold converges to zero, so $\lim_{t\rightarrow\infty}y(t)=0$ can hold only if (13b) holds, in which case the above equation reduces to (13a).
(If). We now prove that (13) implies ${\lim_{t\rightarrow\infty}y(t)=0}$. By Theorem 9, the center manifold $z=\sigma(\theta)$ is locally attractive; namely, $z(t)\rightarrow\sigma(\theta(t))$ as $t\rightarrow\infty.$ Then, the fulfillment of (13b) guarantees that $y(t)\rightarrow 0.$
The two conditions in (13) fully characterize the class of optimization algorithms that achieve asymptotic tracking of a critical trajectory. In words, (4) tracks a critical trajectory if and only if, for some mapping $\sigma,$ the composite function $G_{c}\circ\sigma$ zeros the gradient locally (see (13b)), and the controller $F_{c}(z,y)$ is algebraically related to the exosystem $s(\theta)$ as given by (13a). Notice that, by Theorem 1, the former condition implies that
$x(t)=G_{c}(\sigma(\theta)),$ (15)
is a parameter-feedback optimization algorithm for (1).
Remark 3 (The internal model principle).
We interpret condition (13a) as the internal model principle of time-varying optimization, as it expresses the requirement that any optimization algorithm that achieves asymptotic tracking must include an internal model of the exosystem. Note that the use of a copy of the temporal variability of the optimization problem is explicit in the prediction-correction algorithm [10] (precisely, through the term $\nabla_{xt}f_{0}(x(t),t)$ – see Remark 1 for notation). $\square$
It is important to note that, by Theorem 4, the exosystem state $\theta$ and that of the optimization $z$ must be related, everywhere in $\Theta_{\circ},$ by the relationship:
$z(t)=\sigma(\theta(t)).$ (16)
Intuitively, (16) is interpreted as the existence of a change of coordinates between the state of the exosystem and that of the optimization (we discuss the invertibility properties of $\sigma(\theta)$ shortly below; see Section IV-B and Remark 4).
Remark 4.
An important special case is obtained when $\sigma$ is the identity operator on $\Theta$; in this case, (13) simplifies to:
$s(\theta)$ $=F_{c}(\theta,0),$
which states that the controller vector field $F_{c}(z,y)$ must coincide with that of the exosystem $s(\theta)$ in $\Theta_{\circ}.$ In this case, (16) gives $z(t)=\theta(t);$ namely, the controller state $z(t)$ and that of the exosystem $\theta(t)$ coincide in $\Theta_{\circ}.$square
While Theorem 4 provides a full characterization of all gradient-feedback algorithms that achieve tracking, it remains to address under what conditions on the loss $f(x,\theta)$ such an algorithm is guaranteed to exist. This question is addressed by the following result.
Theorem 5 (Existence of gradient-feedback algorithms).
Suppose Assumptions 1, 2, 3, and 4 hold. There exists a gradient-feedback optimization algorithm that solves Problem 1 if and only if there exists a mapping $H_{c}:\Theta\to{}^{n}$ that zeros the gradient locally. $\square$
Proof.
(Only if) By Theorem 4, there exists a mapping $z=\sigma(\theta)$ such that (13b) holds. Then, (10) holds immediately by letting $H_{c}(\theta)=G_{c}(\sigma(\theta))$.
(If) We will prove this claim by constructing a gradient-feedback algorithm that achieves $y(t)\rightarrow 0$ as $t\to\infty$.
First, notice that by Assumption 4, there exists a matrix $L$ such that $S-LQ$ has eigenvalues in $\mathbb{C}^{-}.$ Consider the algorithm (4) with $n_{c}=p$ and [^3] Note that the second argument of the gradient is evaluated at the algorithm state $z$ instead of the parameter vector $\theta$ as the latter is unknown.
$F_{c}(z,y)$ $=s(z)+L(y-\nabla_{x}f(H_{c}(z),z)),$
$G_{c}(z)$ $=H_{c}(z),$
where $H_{c}(z)$ is as in (10). The Jacobian of $\dot{z}(t)=F_{c}(z(t),y(t))$ with respect to $z$ is given by $S-LQ$ (notice that two terms of the form $LRM$ with opposite sign cancel out in forming the Jacobian: one from the gradient term $\nabla_{x}f(H_{c}(z),z)$ and the other one from $y$). The claim thus follows by application of Theorem 4 with $\sigma$ the identity operator on $\Theta$. ∎
Interestingly, the conditions for existence of a gradient-feedback algorithm and those for existence of a parameter-feedback algorithm are identical. This is not surprising, since a parameter-feedback algorithm is assumed to have access to $\theta(t)$ while a gradient-feedback algorithm needs to measure $\theta(t)$ indirectly through $y(t).$ More precisely, as stated by (16), the dynamic state of the controller $z(t)$ acts as alternative representation (i.e., in different coordinates) of the exosystem state $\theta(t),$ while the exploration signal $x(t)$ acts as a parameter-feedback algorithm (see (15)).
The proof of Theorem 5 is constructive, as it provides a design procedure to construct $F_{c}(z,y)$ and $G_{c}(z)$ that constitute a gradient-feedback optimization algorithm. Such a procedure is presented in Algorithm 1, where a Luenberger observer is used to estimate the exosystem state $\theta(t)$ (see line $4$), and a parameter feedback algorithm is then applied to the estimated exosystem state to regulate the gradient to zero (precisely, $G_{c}(z)$ is designed following the approach of Theorem 1 – see line $3$).
Data: $s(\theta),$ $\nabla_{x}f(x,\theta),$ $H_{c}(\theta)$ satisfying (13), Jacobian matrices $Q$ and $S$ in (7)
1 $n_{c}\leftarrow n$;
2 $L\leftarrow$ any matrix such that $S-LQ$ is Hurwitz;
3 $G_{c}(z)\leftarrow H_{c}(z)$;
4 $F_{c}(z,y)\leftarrow s(z)+L(y-\nabla_{x}f(H_{c}(z),z))$;
Result: $F_{c}(z,y)$, $G_{c}(z)$, and $n_{c}$ that solve Problem 1
Algorithm 1 Gradient-feedback algorithm design
Remark 5.
Instead of a Luenberger observer, alternative dynamic observers could be considered in Line 4 of the algorithm to achieve different asymptotic or transient properties of the resulting gradient-feedback optimization algorithm; we leave the investigation of alternative state observer algorithms as the scope of future works. $\square$
We illustrate the applicability of (13) on a quadratic problem in the following example.
Example 4.
Consider the quadratic problem studied in Example 3, and assume that the exosystem follows the linear model $\dot{\theta}=S\theta$ for some matrix $S\in{}^{p\times p}.$ According to Theorem 5, an optimization algorithm given by
$\dot{z}=A_{c}z+B_{c}y,$ $x=G_{c}z,$ $y=Rx+Q\theta,$ (17)
where $A\in{}^{n_{c}\times n_{c}},$ $B_{c}\in{}^{n_{c}\times n},$ $G_{c}\in{}^{n\times n_{c}},$ achieves asymptotic tracking if and only if there exists a linear transformation $\Sigma\in{}^{n_{c}\times p}$ such that:
$\Sigma S$ $=A_{c}\Sigma,$ (18a)
00 0 $=(RG_{c}\Sigma+Q),$ (18b)
for all $\theta$ in a neighborhood of the origin. When this condition holds, the application of Algorithm 1 gives
$A_{c}=S,$ $B_{c}=L,$ $G_{c}=-R^{\dagger}Q,$
where $L$ is any matrix such that $S-LQ$ is Hurwitz; notice that this choice satisfies (18) with $\Sigma=I.$square
IV-B The optimization algorithm as a copy of the exosystem
As discussed after (16), the state of any optimization algorithm that achieves asymptotic tracking and that of the exosystem must be related by a change of coordinates. In the following, we show that $\sigma(\theta)$ is indeed an injective map, and it can thus be interpreted as a change of coordinates. We begin with the following result.
Proposition 6.
Let the assumptions of Theorem 5 hold and let $F_{c}(z,y)$ and $G_{c}(z)$ be obtained by Algorithm 1. Then, there exists an injective map $\sigma(\theta)$ such that (16) holds. $\square$
Proof.
The claim follows by noting that $F_{c}(z,y)$ and $G_{c}(z)$ obtained by Algorithm 1 satisfy (13) with $\sigma$ the identity operator on $\Theta$. ∎
The interpretation of Proposition 6 is that, for any solution returned by Algorithm 1, $F_{c}(z,y)$ incorporates a copy of $s(\theta);$ precisely,
$F_{c}(\sigma(\theta),0)=s(\theta).$ (19)
The above feature is not a result of using Algorithm 1. Indeed, every gradient feedback algorithm has this feature, as shown next.
Proposition 7.
Let the assumptions of Theorem 5 hold and assume that $S(\theta)$ is $C^{\infty}.$ For any $F_{c}(z,y)$ and $G_{c}(z)$ that achieve asymptotic tracking, there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin such that $\sigma(\theta)$ is injective in $\Theta_{\circ}$$\square$
Proof.
By contradiction, assume that $F_{c}(z,y)$ and $G_{c}(z)$ satisfy (13), but $\sigma(\theta)$ is not injective at the origin; namely, there exists nonzero $\theta^{\prime}\in\Theta_{\circ}$ such that $\sigma(\theta^{\prime})=z_{0}^{\star}.$ From (13b), we have
$0=\nabla_{x}f(G_{c}(\sigma(\theta^{\prime})),\theta^{\prime}).$
Moreover, from Theorem 4,
$\frac{\partial\sigma(\theta^{\prime})}{\partial\theta}s(\theta^{\prime})$ $=F_{c}(\sigma(\theta^{\prime}),0)=0.$ (20)
For $C^{\infty}$ vector fields $h_{1}(x)$ and $h_{2}(x)$, we let
$L_{h_{1}}(h_{2})(x)=\frac{\partial h_{2}(x)}{\partial x}h_{1}(x).$
By application of (20), we have
$0=L_{s}(\nabla_{x}f)(x,\theta^{\prime})=L_{s}(L_{s}(\nabla_{x}f))(x,\theta^{\prime})=\dots.$
Hence, the matrix
$\begin{bmatrix}L_{s}(\nabla_{x}f)(x,\theta)\\ L_{s}(L_{s}(\nabla_{x}f))(x,\theta)\\ \vdots\end{bmatrix},$
is not invertible at $\theta=\theta^{\prime}.$ By [35, Thm. 3.13], the system (2) is not weakly observable, thus violating Assumption 4. ∎
V Fidelity of the internal model and tracking accuracy
In this section, we illustrate through examples how an imprecise knowledge of the internal model impacts the tracking accuracy. For tractability, we will focus on the quadratic problem (12) (see Examples 3 and 4). Suppose that an imprecise knowledge of the internal model is available; namely, there exists $\Delta\in{}^{n_{c}\times p}$ such that (18) is modified to:
$\Sigma S+\Delta$ $=A_{c}\Sigma,$ (21a)
00 0 $=(RG_{c}\Sigma+Q).$ (21b)
To analyze the asymptotic behavior of $y(t),$ it is useful to define the auxiliary variable $\tilde{z}(t)=z(t)-\Sigma\theta(t).$ Using (21) and (17), the dynamics of $\tilde{z}$ follow the model:
$\dot{\tilde{z}}=(A_{c}+B_{c}RG_{c})\tilde{z}+\Delta\theta.$ (22a)
The corresponding gradient signal in terms of $\tilde{z}$ is:
$y=RG_{c}\tilde{z}.$ (22b)
Assuming that $(A_{c}+B_{c}RG_{c})$ is Hurwitz stable (see Theorem 4), the Final Value Theorem gives:
$y_{\infty}$ $:=\lim_{t\rightarrow\infty}y(t)$ (23)
[…]
As a first illustrative scenario, suppose $S=0;$ namely, that the exosystem states are constants at all times. Notice that, since $S=0,$ from (18a), $A_{c}$ incorporates an internal model of $S$ if and only if $A_{c}\Sigma=0;$ precisely, if and only if
$\text{dim}(\ker(A_{c}))\geq p.$
In this case, assuming that the perturbed $A_{c}$ is such that $(A_{c}+B_{c}RG_{c})$ remains Hurwitz stable, from (23), we have:
[…]
As expected, when $\Delta=0,$ $A_{c}$ incorporates an exact internal model, and $y_{\infty}=0$ for any $\theta(0).$ On the other hand, when $\Delta\neq 0,$
[…]
namely, bounded errors in $\Delta$ result in a bounded $y_{\infty}.$
Interestingly, analogous continuity properties may not hold when the Poisson stability assumption is dropped. Consider an instance of (12) with $n=p=2,$ $R=Q=I,$ and
$S=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}.$
Suppose the algorithm (17) is used, with $G_{c}=I,B_{c}=I,$
$A_{c}=\begin{bmatrix}-\epsilon_{1}&1\\ 0&-\epsilon_{2}\end{bmatrix},$
where $0\leq\epsilon_{1}<1,$ $10\leq\epsilon_{2}<$ 1 . To avoid trivial cases, we will assume $\theta_{1}(0)$ and $\theta_{2}(0)$ are nonzero. When $\epsilon_{1}=\epsilon_{2}=0,$ (21) holds with $\Sigma=-I$ and $\Delta=0$, in which case $y_{\infty}=0.$ On the other hand, when $\epsilon_{1},\epsilon_{2}\neq 0,$ from (23):
$y_{\infty}=\lim_{s\rightarrow 0}\begin{bmatrix}\frac{\epsilon_{1}}{s-\alpha_{1}}&\frac{\epsilon_{1}}{s(s-\alpha_{1})}+\frac{\epsilon_{2}}{(s-\alpha_{1})(s-\alpha_{2})}\\ 0&\frac{\epsilon_{2}}{s-\alpha_{2}},\end{bmatrix}\begin{bmatrix}\theta_{1}(0)\\ \theta_{2}(0)\end{bmatrix}$
where $\alpha_{i}=1-\epsilon_{i}$ for $i\in\{1,2\}.$ Since $\epsilon_{1}\neq 0$ and $\theta_{2}(0)\neq 0$, we conclude that $|y(t)|\rightarrow\infty$ as $t\rightarrow\infty$, i.e., asymptotic tracking is not attained.
In conclusion, the tracking accuracy depends on the fidelity of the internal model as well as the asymptotic behavior of the exosystem.
VI Extensions to constrained problems
We can extend our results to study constrained time-varying optimization problems by searching for a stationary point of the Lagrangian.
Consider the equality-constrained problem
minimize $f(x,\theta(t))$
subject to $h_{i}(x,\theta(t))=0,\quad i=1,\ldots,m$
where the constraint functions $h_{i}(x,\theta(t))$ depend on the parameter vector. The associated Lagrangian function is
$L(x,\lambda,\theta(t))=f(x,\theta(t))+\sum_{i=1}^{m}\lambda_{i}h_{i}(x,\theta(t))$
[…]
$L(x,\bar{\lambda},\theta(t))\leq L(x,\lambda,\theta(t))\leq L(\bar{x},\lambda,\theta(t))$
for all pairs $(\bar{x},\bar{\lambda})\in{}^{n}\times{}^{m}$. For any such saddle-point, $x$ is primal optimal, $\lambda$ is dual optimal, and the optimal duality gap is zero. Moreover, the gradient of the Lagrangian (assuming it exists) is zero at any saddle-point. It follows from the derivations in the previous sections that the gradient-feedback and parameter-feedback algorithms can be directly applied to seek a stationary point of the Lagrangian function by replacing (in (1)) the variable $x$ with the extended decision variable $\tilde{x}=(x,\lambda)$ and by letting $f(\tilde{x},\theta)=L(x,\lambda,\theta).$ Notice that, if the critical point computed by (4a) is also a saddle-point, then it is also a solution to the equality-constrained problem; see [36, Ch. 5].
VII Simulation results
In this section, we illustrate our results and optimization design method through a set of numerical simulations.
Figure 3: Simulation results illustrating the performance of an optimization algorithm synthesized using Algorithm 1 for the quadratic instance (12) of (1). See Example 3 and Section VII-A for a discussion. (Top) Illustration of the temporal variability of the parameter $\theta(t)$ and of $z(t).$ (Second from top) $z(t)$ is an estimator for $\theta(t),$ and thus $z(t)\rightarrow\theta(t)$ as $t\rightarrow\infty.$ (Third from top) The proposed control algorithm is successful in regulating the gradient feedback signal $y(t)$ to zero asymptotically. (Bottom) Illustration that $x(t)\rightarrow x^{\star}(t)$ as $t\rightarrow\infty.$
VII-A Optimization design for quadratic costs
[…]
$F_{c}(z,y)$ $=s(z)+L(y-\nabla_{x}f(H_{c}(z),z)),$
$=Sz+L(y-(RG_{c}z+Qz)),$
$=Sz+Ly.$
The simulation results are presented in Figure 3. It follows from our choice of $F_{c}(z,y),$ that the optimization state $z(t)$ converges to the parameter vector $\theta(t)$ (see the top two plots in Fig. 3). Moreover, it follows from our choice of $G_{c}(z)$ that $y(t)\rightarrow 0$ and thus $x(t)\rightarrow x^{\star}(t)$ as $t\rightarrow\infty$ (see the bottom two plots in Fig. 3).
Figure 4: (Left) Areal view of the highway system between the cities of Wavre and Brussels, Belgium. (Right) Graph utilized to model the portion of traffic network of interest.
VII-B Application to solve the dynamic traffic assignment problem in transportation
We next illustrate the applicability of the framework in solving the dynamic traffic assignment problem in roadway transportation [37]; intuitively, the objective is to decide how traffic flows split among the available paths of a network to minimize the drivers’ travel time to destination. We model a roadway transportation network using a static flow model [37], described by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E}),$ with edges $i\in\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ (modeling traffic roads) and nodes $\mathcal{V}$ (modeling traffic junctions). For $i\in\mathcal{E},$ we denote by $i^{+}\in\mathcal{V}$ and $i^{-}\in\mathcal{V}$ its origin and destination nodes, respectively. We assume that an exogenous, time-varying, inflow of traffic $\theta(t)$ enters the network at a certain origin node, denoted by $o\in\mathcal{V},$ and exits at a certain destination node, denoted by $d\in\mathcal{V};$ for simplicity, we assume that there is only one origin-destination pair, but this is without loss of generality [37]. We describe the network state using a vector $x\in\mathbb{R}_{\geq 0}^{|\mathcal{E}|}$ (where $|\mathcal{E}|$ denotes the number of edges) whose entries $x_{i}$ describe the amount of inflow $\theta$ routed through road $i.$ To each link $i,$ we associate a function $\ell_{i}(x_{i})$ describing the latency (or travel time) of road $i$. For our simulations, we consider the network topology in Fig. 4, and choose the latency functions as follows:
$\ell_{1}(x_{1})$ $=x_{1},$ $\ell_{2}(x_{2})$ $=10x_{2},$ $\ell_{3}(x_{3})$ $=x_{3},$
$\ell_{4}(x_{4})$ $=5x_{4},$ $\ell_{5}(x_{5})$ $=x_{5}.$
According to Wardrop’s first principle [37, pp. 31], transportation networks operate at a condition where travelers select their path to minimize their travel time to destination. Mathematically, a Wardrop’s equilibrium is the optimizer of the following optimization problem:
[…]
subject to: $\sum_{j\in\mathcal{E}:j^{-}=v}x_{j}-\sum_{j\in\mathcal{E}:j^{+}=v}x_{j}=\delta_{v}(\theta(t)),\ \forall v\in\mathcal{V},$
$x_{i}\geq 0,\ \forall i\in\mathcal{E},$ (24)
where $\delta_{v}(\theta),$ $v\in\mathcal{V},$ is defined as:
$\delta_{v}(\theta)=\begin{cases}\theta,&\text{if }v=o,\\ -\theta,&\text{if }v=d,\\ 0,&\text{otherwise}.\end{cases}$
The loss function in (VII-B) is used to model travelers who will switch to a different path if it has shorter travel time to destination, while the first constraint in (VII-B) describes the network topology, namely, that traffic flows are conserved at each node. Notice that (VII-B) is a time-varying optimization problem, where the temporal variability originates from the dependence of the constraint on $\theta(t)$, which describes the inflow of vehicles at the origin and outflow at the destination, measured in vehicles per hour. For our simulations, we assume that the network inflow is sinusoidal:
$\theta(t)=\theta_{0}-\theta_{1}\cos(\omega_{1}t+\phi_{1})-\theta_{2}\cos(\omega_{2}t+\phi_{2}),$ (25)
where $\theta_{0},\theta_{1},\theta_{2},\omega_{1},\omega_{2},\phi_{1},\phi_{2}\in\mathbb{R}_{>0},$ satisfy $\theta_{0}>\theta_{1},$ $\theta_{0}>\theta_{2},$ and $\omega_{2}>\omega_{1}.$ The model (25) states that the network inflow is the sum of a constant term, $\theta_{0},$ a slowly-varying sinusoid with angular frequency $\omega_{1}$ and a quickly-varying sinusoid with angular frequency $\omega_{2}.$ The low-frequency sinusoid is used here to describe slowly-varying (e.g., hourly) traffic demands, while the high-frequency sinusoid is used to model sudden (e.g., at the minute-level) variations in traffic demand. For our simulations, we let $\theta_{0}=3$ veh/h, $\theta_{1}=1$ veh/h, $\theta_{2}=0.1$ veh/h, $\omega_{1}=0.1$ rad/hour, $\omega_{2}=\sqrt{50}$ rad/hour, and $\phi_{1}=\phi_{2}=0,$ see Fig. 5 (top).
We applied Algorithm 1 to derive an optimization algorithm to solve the traffic assignment problem (VII-B). For the synthesis, we utilized the internal model $s(z)=Sz,$ where
$S$ $=\begin{bmatrix}0&1&0&0&0\\ -\omega_{1}^{2}&0&0&0&0\\ 0&0&0&1&0\\ 0&0&-\omega_{2}^{2}&0&0\\ 0&0&0&0&0\end{bmatrix}.$
Notice that knowledge of $\theta_{0},\theta_{1},\theta_{2},\phi_{1},\phi_{2}$ is not required to synthesize the optimization algorithm – only the frequencies $\omega_{1},\omega_{2}$ are required to be known. To seek a solution to the constrained problem (VII-B), consider the Lagrangian function:
[…]
$\quad+\sum_{v\in\mathcal{V}}\lambda_{v}\left(\delta_{v}(\theta(t))-\sum_{\begin{subarray}{c}j\in\mathcal{E}:\\ j^{-}=v\end{subarray}}x_{j}+\sum_{\begin{subarray}{c}j\in\mathcal{E}:\\ j^{+}=v\end{subarray}}x_{j}\right),$
where $\lambda:=(\lambda_{1},\dots,\lambda_{|\mathcal{V}|})$ is the vector of Lagrange multipliers. We applied Algorithm 1 to the optimization problem (1) with $f(\tilde{x},\theta)=L(x,\lambda,\theta)$ (see Section VI); the inequality constraints in (VII-B) have been accounted for by projecting $x(t)$ onto the feasible set. Here, matrix $L$ has been chosen so that the eigenvalues of $S-LQ$ are uniformly distributed in the open real interval $(-1,-2)$. Notice that, since the latencies are strictly increasing, the Lagrangian is strongly convex-strongly concave [37], and thus the problem admits a unique critical point that is a saddle point. It follows that our algorithm is guaranteed to converge to a minimizer of (VII-B). Simulation for this problem are presented in Fig. 5. The simulation shows that $z(t)\rightarrow\theta(t)$ as $t\rightarrow\infty$ (see Fig. 5-second figure) and $y(t)\rightarrow 0,$ which implies that the algorithm successfully computes the critical point (see Fig. 5-third figure). Finally, the bottom figure of Fig. 5 illustrates the rate of convergence of the algorithm, which, as expected, is governed by the placement of the eigenvalues of the observer.
Figure 5: Simulation results illustrating the performance of the optimization algorithm synthesized using Algorithm 1 to solve the dynamic traffic assignment problem (VII-B). See Section VII-B for a discussion. The bottom figure show that, from a zero initial condition, the algorithm is capable of computing a Wardrop equilibrium in about $1.5$ hours, and then of tracking this equilibrium. Notice that, a faster rate of convergence could be obtained by shifting the eigenvalues of $S-LQ.$
VIII Conclusions
We showed that the problem of designing optimization algorithms for time-varying optimization problems can be reconducted to the output regulation problem of nonlinear multivariable systems. This connection allowed us to prove the internal model principle of time-varying optimization, which states that asymptotic tracking can be achieved only if the optimization algorithm incorporates a reduplicated model of the temporal variability of the problem. Further, we showed that asymptotic tracking can be achieved under more relaxed assumptions (namely, assumption (A2)) than what is normally imposed in the literature (see assumption (A3)). Moreover, our algorithm structure is novel in the literature, and it relies on the use of an observer for the temporal variability of the problem. This work opens the opportunity for several directions of future work, including an investigation at discrete-time and possibly the application in feedback optimization.
We now summarize relevant facts in center manifold theory from [18]; see also [38]. Consider the nonlinear system:
$\dot{x}=f(x)$ (26)
where $f$ is a $C^{k}$ vector field defined on an open subset $U$ of n, and let $x_{\circ}\in U$ be an equilibrium point for $f$, i.e., $f(x_{\circ})=0$. Without loss of generality, suppose $x^{\circ}=0$. Let $F=\left[\frac{\partial f}{\partial x}\right]_{x=0},$ denote the Jacobian matrix of $f$ at $x=0$. Suppose the matrix $F$ has $n^{\circ}$ eigenvalues with zero real part, $n^{-}$ eigenvalues with negative real part, and $n^{+}$ eigenvalues with positive real part. Let $E^{-},E^{\circ},$ and $E^{+}$ be the (generalized) real eigenspaces of $F$ associated with eigenvalues of $F$ lying on the open left half plane, the imaginary axis, and the open right half plane, respectively. Note that $E^{\circ},E^{-},E^{+}$ have dimension $n^{\circ},n^{-},n^{+}$, respectively and that each of these spaces is invariant under the flow of $\dot{x}=Fx.$ If the linear mapping $F$ is viewed as a representation of the differential (at $x=0$) of the nonlinear mapping $f$, its domain is the tangent space $T_{0}U$ to $U$ at $x=0$, and the three subspaces in question can be viewed as subspaces of $T_{0}U$ satisfying $T_{0}U=E^{\circ}\oplus E^{-}\oplus E^{+}.$ We refer to [39, Sec. A.II] for a precise definition of $C^{k}$ manifolds; loosely speaking, a set $S\subset U$ is a $C^{k}$ manifold it can be locally represented as the graph of a $C^{k}$ function.
Definition 3 (Locally invariant manifold).
A $C^{k}$ manifold $S$ of $U$ is said to be locally invariant for (26) if, for each $x_{\circ}\in S$, there exists $t_{1}<0<t_{2}$ such that the integral curve $x(t)$ of (26) satisfying $x(0)=x_{\circ}$ is such that $x(t)\in S$ for all $t\in\left(t_{1},t_{2}\right)$. $\square$
Intuitively, by letting $x=(y,\theta)$ and expressing (26) as:
$\dot{y}=f_{y}(\theta,y),$ $\dot{\theta}=f_{\theta}(\theta,y),$ (27)
a curve $y=\pi(\theta)$ is an invariant manifold for (27) if the solution of (27) with $\theta(0)=\theta_{\circ}$ and $y(0)=\pi(\theta_{\circ})$ lies on the curve $y=\pi(\theta)$ for $t$ in a neighborhood of $0$ The notion of invariant manifold is useful as, under certain assumptions, it allow us to reduce the analysis of (26) to the study of a reduced system in the variable $\theta$ only. The remainder of this section is devoted to formalizing this fact.
Definition 4 (Center manifold).
Let $x=0$ be an equilibrium of (26). A manifold $S$, passing through $x=0$, is said to be a center manifold for (26) at $x=0$ if it is locally invariant and the tangent space to $S$ at 0 is exactly $E^{\circ}$. $\square$
Returning to the decomposition (27), intuitively, the invariant manifold $y=\pi(\theta)$ is said to be a center manifold when all orbits of $y$ decay to zero and those of $\theta$ neither decay nor grow exponentially.
In what follows, we will assume that all eigenvalues of $F$ have nonpositive real part, i.e.,
$n^{+}=0.$ (28)
When (28) holds, it is always possible choose coordinates in $U$ such that (26) reads as:
$\dot{y}=Ay+g(y,\theta)$ (29a)
$\dot{\theta}=B\theta+h(y,\theta)$ (29b)
where $A$ is an $n^{-}\times n^{-}$ matrix having all eigenvalues with negative real part, $B$ is an $n^{\circ}\times n^{\circ}$ matrix having all eigenvalues with zero real part, and the functions $g$ and $h$ are $C^{k}$ functions vanishing at $(y,\theta)=(0,0),$ together with all their first-order derivatives. Because of their equivalence, any conclusion drawn for (29) will apply also to (26). The following result ensures the existence of a center manifold.
Theorem 8 (Center manifold existence theorem).
Assume that (28) holds. There exist a neighborhood $V\subset{}^{n^{0}}$ of 00 and a class $C^{k-1}$ mapping $\pi:V\rightarrow{}^{n^{-}}$ such that the set
$S=\{(y,\theta)\in{}^{n^{-}}\times V:y=\pi(\theta)\},$
is a center manifold for (29). $\square$
Some important observations are in order. By definition, a center manifold for (29) passes through $(0,0)$ and is tangent to the subset of points whose $y$ coordinate is zero. Namely,
$\pi(0)=0\qquad\text{and}\qquad\frac{\partial\pi}{\partial\theta}(0)=0.$ (30)
Moreover, this manifold is locally invariant for (29): this imposes on the mapping $\pi$ the constraint:
$\frac{\partial\pi}{\partial\theta}(B\theta+h(\pi(\theta),\theta))=A\pi(\theta)+g(\pi(\theta),\theta),$ (31)
as deduced by differentiating with respect to time any solution $(y(t),\theta(t))$ of (29) on the manifold $y(t)=\pi(\theta(t))$. In other words, any center manifold for (29) can equivalently be described as the graph of a mapping $y=\pi(\theta)$ satisfying the partial differential equation (31), with the constraints specified by (30).
Remark 6.
Theorem 8 shows existence but not uniqueness of a center manifold. Moreover, (i) if $g$ and $h$ are $C^{k},k\in\mathbb{N}_{>0},$ then (29) admits a $C^{k-1}$ center manifold; (ii) if $g$ and $h$ are $C^{\infty}$ functions, then (29) has a $C^{k}$ center manifold for any finite $k$, but not necessarily a $C^{\infty}$ center manifold. $\square$
The next result shows that any $y$-trajectory of (29), starting sufficiently close to the origin, converges, as time tends to infinity, to a trajectory that belongs to the center manifold.
Theorem 9.
Assume that (28) holds and suppose $y=\pi(\theta)$ is a center manifold for (29) at $(0,0)$. Let $(y(t),\theta(t))$ be a solution of (29). There exists a neighborhood $U^{\circ}$ of $(0,0)$ and real numbers $M>0$ and $K>0$ such that, if $(y(0),\theta(0))\in U^{\circ}$, then for all $t\geq 0$,
$\|y(t)-\pi(\theta(t))\|\leq Me^{-Kt}\|y(0)-\pi(\theta(0))\|.$square
From the above discussion, any trajectory of (29) starting at a point $y^{\circ}=\pi\left(\theta^{\circ}\right)$ of a center manifold satisfies:
$y(t)=\pi(\zeta(t)),$ $\theta(t)=\zeta(t),$
where $\zeta(t)$ is any solution of
$\dot{\zeta}=B\zeta+h(\pi(\zeta),\zeta),$ (32)
satisfying the initial condition $\zeta(0)=\theta_{\circ}$. This decomposition allows us to predict the asymptotic behavior of (29) by studying the asymptotic behavior of a reduced-order system, namely, (32). This is formalized in the following result.
Theorem 10.
Suppose $\zeta=0$ is a stable (respectively, asymptotically stable, unstable) equilibrium of (32). Then $(y,\theta)=(0,0)$ is a stable (respectively, asymptotically stable, unstable) equilibrium of (29). $\square$

### v2

@@
 Abstract
-Time-varying optimization problems are central to many engineering applications where performance metrics and system constraints evolve dynamically with time.
-A number of algorithms have been proposed in recent years to solve such problems; a common feature of all these methods is that they implicitly require precise knowledge of the temporal variability of the solutions in order to exactly track the optimizers.
-In this paper, we seek to lift these stringent assumptions.
-Our main result is a fundamental characterization showing that an algorithm can track an optimal trajectory if and only if it contains a model of the temporal variability of the problem.
-We refer to this concept to as the internal model principle of time-varying optimization.
-By recasting the optimization objective as a nonlinear regulation problem and using tools from center manifold theory, we provide necessary and sufficient conditions both for an optimization algorithm to achieve exact asymptotic tracking and for such an algorithm to exist.
-We illustrate the applicability of the approach numerically on both synthetic problems as well as practical problems in transportation.
-I Introduction
+Time-varying optimization problems are central to many engineering applications, where performance metrics and system constraints evolve dynamically with time.
+Several algorithms have been proposed to address these problems; a common characteristic among them is their implicit reliance on knowledge of the optimizers’ temporal variability.
+In this paper, we provide a fundamental characterization of this property: we show that an algorithm can track time-varying optimizers if and only if it incorporates a model of the temporal variability of the optimization problem.
+We refer to this concept as the internal model principle of time-varying optimization.
+Our analysis relies on showing that time-varying optimization problems can be recast as output regulation problems and, by using tools from center manifold theory, we establish necessary and sufficient conditions for exact asymptotic tracking.
+As a result, these findings enable the design of new algorithms for time-varying optimization.
+We demonstrate the effectiveness of the approach through numerical experiments on both synthetic problems and the dynamic traffic assignment problem from traffic control.
+{IEEEkeywords}
+Time-varying optimization, Internal model principle, Gradient regulation, traffic control.
+1 Introduction
+[…]
 Time-varying optimization problems play a central role in several scientific domains, as they underpin many important contemporary engineering problems.
 Examples include training in Machine Learning [1, 2], dynamic signal estimation in Signal Processing [3], trajectory tracking in Robotics [4], system optimization in Industrial Control [5], and much more.
-Historically, discrete-time algorithms for time-varying optimization have been proposed and studied first since they emerge as a natural extension of their time-invariant counterparts, allowing for cost functions and constraints that may change over time [6].
+Historically, discrete-time algorithms for time-varying optimization have been proposed and studied first since they emerged as a natural extension of their time-invariant counterparts, allowing for cost functions and constraints that may change over time [6].
 These approaches build on the classical perspective on mathematical optimization, which seeks to construct methods to determine optimizers and consist of iterative procedures implemented on digital devices.
-Recently, there has been a renewed interest in the use and development of continuous-time dynamics for optimization purposes, due to the possibility of utilizing tools from dynamical systems for their analysis [7, 8] and motivated by practice, where optimization is increasingly used to control physical systems [9].
-Motivated by these recent developments, in this paper we study time-varying convex optimization problems and focus on the use of continuous-time dynamics to track exactly optimal solutions.
-Although, by now, several approaches have been developed for this purpose [10, 11, 12, 13], all these techniques implicitly require full knowledge of the temporal variability of the problem [13].
+In recent years, however, numerous optimization algorithms have been formulated and studied as continuous-time processes [7, 8, 9], mainly due to the wide availability of tools for Ordinary Differential Equations (ODEs) that can facilitate their analysis [9].
+Motivated by these recent developments, in this paper, we study time-varying convex optimization problems and focus on the use of continuous-time dynamics to track exactly optimal solutions.
+Several approaches have already been developed for this purpose [10, 11, 12, 13], yet all these techniques implicitly require full knowledge of the temporal variability of the problem [13].
 Unfortunately, in most practical applications, having such knowledge is impractical, either because the temporal variability enters the optimization in the form of exogenous disturbances that are unknown and cannot be measured (see, e.g., [14, 15]), or simply because it is unrealistic to ask for a noiseless model of how the problem changes with time.
@@
 Our approach relies on reinterpreting the optimization algorithm design as a nonlinear, multivariable regulation problem [18], and our analysis uses tools from center manifold theory [19, 17].
-Figure 1 illustrates the architecture of the gradient-feedback design scheme studied in this paper.
+Fig.
+1 illustrates the architecture of the gradient-feedback design scheme studied in this paper.
 Figure 1: Architecture of the gradient-feedback design scheme studied in this work.
@@
 Because the temporal variability of the problem is unknown (or ignored), algorithms in this class can make a decision only after each variation has been observed, thus incurring a certain ‘regret.’ Mathematically, these techniques are capable of reaching only a neighborhood of an optimizer [13], and exact tracking is out of reach in general for these approaches.
-In contrast, the second class of methods uses a model of the temporal evolution of the problem to seek to exactly track optimal trajectories [26, 2, 27, 28].
-Particularly celebrated is the prediction-correction algorithm (see [10] and the recent work [11]), whereby at each time a prediction step is used to anticipate how the optimizer will evolve over time, and a correction step is used to seek a solution to each instantaneous optimization problem.
+Moreover, suitable assumptions need to be made to provide convergence guarantees for these methods; we refer to [13, 26] and references therein for a detailed discussion.
+In contrast, the second class of methods uses a model of the temporal evolution of the problem to exactly track the optimal trajectories [27, 2, 28, 29].
+Particularly celebrated is the prediction-correction algorithm (see [10], the recent work [11], and the discrete implementations [30]), whereby at each time a prediction step is used to anticipate how the optimizer will evolve over time, and a correction step is used to seek a solution to each instantaneous optimization problem.
 We refer to [13] for a recent overview of the topic.
-Recent years have witnessed a growing interest in this problem: [4] uses contraction to study these methods; [29] uses sampling to estimate the temporal variability of the problem; a recent survey on analyzing optimization algorithms using tools from control has appeared in [30]; constrained optimization problems are studied in [31].
-Despite these efforts, the fundamental question of what is the basic algorithm structure that enables exact asymptotic tracking remains unanswered.
+Recent years have witnessed a growing interest in this problem: [4] uses contraction to study these methods; [31] uses sampling to estimate the temporal variability of the problem; a recent survey has appeared in [32]; constrained optimization problems are studied in [33]; recent works [34, 35, 36, 37] have modified these methods to achieve fixed-time convergence.
+With respect to this body of literature, which assumes precise knowledge of the temporal variability of the problem (see detailed discussion in Section 2.1), our focus here is to answer the following fundamental question: what is the least-restrictive information that is needed to track, asymptotically with zero error, the optimizer of a time-varying problem?
+To the best of the authors’ knowledge, a rigorous answer to this question is missing in the literature.
+In this work, we recast the problem of designing optimization algorithms for time-varying optimization as a nonlinear output regulation problem.
+This connects our work with the literature on output regulation, which is a well-established area of research.
+Initial works can be traced back to the 1970s [16, 38, 18] focusing on linear systems, and later extended to nonlinear systems using both local [39, 40] and global [41] approaches for the analysis.
+In recent years, the field has received new attention using modern methods, see, e.g., [42, 43, 44] and the recent tutorial [45].
+To the best of our knowledge, this is the first work in the literature that establishes a connection between output regulation and optimization algorithms design.
 Contributions.
-This paper makes four main contributions.
-First, we recast the problem of designing a time-varying optimization algorithm as a nonlinear multivariable regulation problem, whereby an algorithm is to be designed to regulate the gradient of a certain function to zero.
-We leverage this formulation to characterize the class of optimization algorithms that can achieve tracking.
-In a net departure from existing approaches (e.g., [23, 13, 11, 12]), our characterization is general and allows us to study not only a single optimization method, but an entire class, which enables us to derive fundamental results for all algorithms in the class.
+This paper features four main contributions.
+First, we recast the problem of designing a time-varying optimization algorithm as a nonlinear multivariable regulation problem, whereby the signal to be regulated is the gradient of the loss function.
+We leverage this formulation to derive necessary and sufficient conditions to achieve exact asymptotic tracking for a large class of optimization methods (described by sufficiently-smooth functions).
+In a net departure from existing approaches (e.g., [24, 13, 11, 12]), our characterization is general and allows us to study not only a single optimization method, but an entire class, which enables us to derive fundamental results for all algorithms in this class.
 Second, by harnessing tools from center manifold theory [19, 17], we provide necessary and sufficient conditions for an optimization algorithm to ensure tracking.
@@
 This property allows us to prove the internal model principle of time-varying optimization, which states that for an optimization algorithm to achieve asymptotic tracking, it must incorporate a reduplicated model of the temporal variability of the problem.
-This feature is implicit in all existing approaches for time-varying optimization [13] but, to the best of the authors’ knowledge, lacked a formal proof until now.
+This feature is implicit in all existing approaches for time-varying optimization [10, 11, 34, 36, 37] but, to the best of the authors’ knowledge, lacked a rigorous understanding until now.
 Third, we derive necessary and sufficient conditions for the existence of a tracking algorithm.
 Fourth, we use our characterizations to design algorithms for time-varying optimization.
-With respect to the existing literature, our algorithm does not require one to know or measure exactly the temporal variability of the problem, and thus uses less stringent assumptions.
-Finally, we illustrate the applicability of the approach numerically on both synthetic problems as well as practical problems in transportation.
+With respect to the existing literature, our algorithm does not require one to know or measure exactly the temporal variability of the problem and thus uses less stringent assumptions.
+Finally, we illustrate the applicability of the approach numerically on both synthetic problems and the traffic assignment problem in transportation.
 Organization.
-Section II presents the problem of interest, Section III studies the simpler problem of parameter-feedback optimization, which allows us to derive the necessary framework, and Section IV tackles the time-varying optimization design problem in its full generality and contains our main results.
-Section V discusses the tracking accuracy in relationship to the fidelity of the internal model, Section VI presents some extensions to constrained optimization, Section VII validates numerically the results, and Section VIII illustrates our conclusions.
-Finally, in the Appendix we summarize basic facts on center manifold theory that are used extensively throughout the paper.
+Section 2 presents the problem of interest, Section 3 the parameter-feedback problem (where the temporal variability of the problem can be explicitly measured), and Section 4 the gradient-feedback problem (where the algorithm has access only to first-order functional evaluations).
+Section 5 discusses the tracking accuracy in relationship to the fidelity of the internal model, Section 6 presents extensions to constrained optimization and discrete-time algorithms, Section 7 validates numerically the results, and Section 8 illustrates our conclusions.
+Finally, in the Appendix, we summarize basic facts on center manifold theory used in the paper.
 […]
-II Problem setting
-II-A Problem statement
+2 Problem setting
+2.1 Optimization objectives
 We consider the time-varying optimization problem:
 $\min_{x\in{}^{n}}f(x,\theta(t)),$ (1)
-where $t\in\mathbb{R}_{\geq 0}$ denotes time and $f:{}^{n}\times\Theta\rightarrow\real,$ $\Theta\subseteq{}^{p},$ is a loss function that is parametrized by the time-varying parameter vector $\theta:\mathbb{R}_{\geq 0}\rightarrow\Theta.$ We make the following assumptions throughout.
+where $t\in\mathbb{R}_{\geq 0}$ denotes time and $f:{}^{n}\times\Theta\rightarrow\real,$ $\Theta\subseteq{}^{p},$ is a loss function that is parametrized by the time-varying parameter vector $\theta:\mathbb{R}_{\geq 0}\rightarrow\Theta.$ We make the following standard assumptions on the loss.
 Assumption 1 (Properties of the objective function).
 The map $x\mapsto f(x,\theta)$ is convex and $x\mapsto\nabla_{x}f(x,\theta)$ is Lipschitz continuous in n, for each $\theta\in\Theta.$square
-Assumption 2 (Existence of an exosystem).
+Convexity and smoothness are standard assumptions in optimization [26], which have been widely used in works on related problems [8, 28, 14, 15, 10, 11, 12, 13, 24, 32, 33, 29, 31].
+In (1), the parameter $\theta(t)$ is used to model the temporal variability of the problem. We will require that $\theta(t)$ belongs to a certain class of temporal variabilities, as specified next.
+Assumption 2 (Class of temporal variabilities).
 There exists a smooth (i.e., $C^{\infty}$) vector field $s:\Theta\rightarrow{}^{p}$ and $\theta(0)\in\Theta$ such that the parameter vector $\theta(t)$ satisfies
-$\dot{\theta}(t)=s(\theta(t))$ (2)
-for all $t\in\mathbb{R}_{\geq 0}.$square
-Convexity and smoothness are standard assumptions in optimization [25].
-Assumption 2 ensures the existence of an autonomous system, called exosystem, describing the class of temporal variabilities of the cost taken into consideration.
-In what follows, we will say that $x^{\star}:\mathbb{R}_{\geq 0}\rightarrow{}^{n}$ is a critical trajectory of (1) if it satisfies:
-$0=\nabla_{x}f(x^{\star}(t),\theta(t)),\quad\forall t\in\mathbb{R}_{\geq 0}.$
-We will assume existence of a critical trajectory and that any critical trajectory is continuous.
-Remark 1.
-The time variation in (1) is captured implicitly through the parameter vector $\theta(t)$. A related, yet slightly more general, problem is as follows:
+$\dot{\theta}(t)=s(\theta(t)),$ (2)
+for $t\in\mathbb{R}_{\geq 0}.$ Moreover, $\theta=0$ is an equilibrium of (2) and the trajectories of (2) are bounded.  $\square$
+We stress that, a priori, we do not assume that $s(\theta)$ nor $\theta(0)$ are known (see Problem 2.2 shortly below for additional details).
+Assumption 2 characterizes the class of temporal variabilities taken into consideration.
+This assumption is mild, as it only requires that $\theta(t)$ is deterministic, sufficiently smooth (so that its derivative is some $C^{\infty}$ function of $\theta(t)$ as in (2)), and its trajectories remain bounded.
+In line with [39, 38], we call the autonomous system (2) the exosystem.
+For simplicity of the presentation, we assume that $\Theta$ is some neighborhood of the origin of p.
+We put no restrictions on the size of this neighborhood (which is, e.g., allowed to be the entire space $\Theta={}^{p}$), and thus on the size of $\theta(t)$ nor on its temporal variation.
+Moreover, there is no restriction with asking that $\Theta$ contains the origin because, if $\theta(t)$ takes values in the neighborhood of any other point, such a point can be shifted to the origin through a time-invariant change of variables without altering the solutions of (1).
+Remark 1 (Parametrization in time-varying optimization).
+In (1), the time variation is captured implicitly by the parameter $\theta(t)$. A related, yet slightly more general, problem is:
 $\min_{x\in{}^{n}}f_{0}(x,t),$ (3)
@@
 On the other hand, in general, there exists an infinite number of ways to parametrize (3) as in (1), thus leading to possible ambiguities.
-For instance, any function $f_{0}(x,t)$ may be parametrized by $\theta(t)=t$ (so that $f_{0}\equiv f$), although this is not always compatible with our assumptions (see Assumption 3 and Remark 2 below).
+For instance, any $f_{0}(x,t)$ may be parametrized by $\theta(t)=t$ (so that $f_{0}\equiv f$), although this is not compatible with our boundedness trajectory assumption (Assumption 2).
 $\square$
-In this work, we are interested in the problem of designing an optimization algorithm that is capable of determining and tracking a critical trajectory of (1).
-We are interested in doing so with as little knowledge as possible on $\theta(t)$ and its temporal variability.
-From an optimization perspective, three types of assumptions may be considered (presented in increasing order of restrictiveness):
-  1. (A1)
-    The parameter signal $\theta(t)$ is known or measurable at each $t\in\mathbb{R}_{\geq 0},$ but no knowledge of $s(\theta)$ is available.
-  2. (A2)
-    The vector field $s(\theta)$ is known, but $\theta(0)$ (and hence also $\theta(t)$) is unknown.
-  3. (A3)
-    The vector field $s(\theta)$ as well as $\theta(0)$ are known.
-Under these assumptions, different methods have been developed in the literature, each with corresponding guarantees:
-  1. (M1)
-    Under (A1), all methods from static optimization (e.g., [23, 24, 19, 20, 21, 22, 13]) can be applied directly.
-All these methods can achieve convergence only to a neighborhood of a critical trajectory.
-  2. (M2)
-    The approach of [12] is applicable in this case, which is however limited to strongly convex quadratic optimization problems with linear temporal variablities.
-  3. (M3)
-    The class of prediction correction methods [10] and their variations [11] take advantage of this knowledge and ensures asymptotic tracking of $x^{\star}(t)$.
-A natural question emerges directly from the above discussion: is (A3) necessary to ensure exact tracking and, if not, can it be relaxed (e.g., to (A1) or (A2))?
-With this motivation, in this paper we are driven by the following question.
-Problem 0 (Minimal knowledge for exact tracking).
-What is the least-restrictive set of assumptions (cf. (A1)–(A3)) under which there exists an optimization algorithm that achieves exact asymptotic tracking of the critical points of (1)?  $\square$
-Our objective is to seek an optimization algorithm that assumes no access to $\theta(t)$; instead, as common in first-order optimization approaches [25], we will assume only the availability of functional evaluations of the gradient function $\nabla_{x}f(x,\theta(t))$ at points $x\in{}^{n},$ selected by the algorithm.
-Formally, the optimization algorithm will be described by an internal state $z(t),$ which takes values on an open subset $\mathcal{Z}\subseteq{}^{n_{c}},n_{c}\in\mathbb{N}_{>0}$; the optimization algorithm generates a sequence of points $x(t)\in{}^{n}$ (called exploration signal) at which the gradient shall be evaluated, and processes functional evaluations of the gradient at these points, $y(t)=\nabla_{x}f(x(t),\theta(t))$ (called gradient feedback signal).
-Mathematically, the optimization algorithm is:
-$\dot{z}(t)$ $=F_{c}(z(t),y(t)),$ $x(t)$ $=G_{c}(z(t)),$ (4a)
-together with the gradient-feedback signal:
-$y(t)=\nabla_{x}f(x(t),\theta(t)),$ (4b)
+Remark 2 (Discrete-time implementations).
+In this work, we pursue the continuous-time time-varying optimization problem (1), together with a continuous-time optimization algorithm.
+This is inspired by control applications (see [14] and references therein), where (1) models performance objectives associated with a physical plant to be controlled, and $\theta(t)$ models exogenous disturbances affecting the plant.
+If instead the parameter vector were to evolve in discrete time, we would have the alternative optimization problem
+$\min_{x\in\mathbb{R}^{n}}f\left(x,\theta_{k}\right),$ (4)
+where $k\in\mathbb{N}_{\geq 0}$ denotes time or iteration, and $f:\mathbb{R}^{n}\times\Theta\rightarrow\mathbb{R}$.
+While we could discretize $\theta(t)$ and then study discrete-time algorithms to solve this problem, physical control applications motivate studying the continuous-time problem directly.
+We provide additional details on discrete-time implementations of our method in Section 6.2.
+$\square$
+In what follows, we say that $x^{\star}(t),$ with $x^{\star}:\mathbb{R}_{\geq 0}\rightarrow{}^{n}$, is a critical trajectory of (1) if it satisfies:
+$0=\nabla_{x}f(x^{\star}(t),\theta(t)),\quad\forall t\in\mathbb{R}_{\geq 0}.$ (5)
+We assume the existence of a critical trajectory and that any critical trajectory is continuous.
+The existence of a critical trajectory can be guaranteed under standard assumptions on the optimization problem; for example, coercivity of the cost [46] (i.e., $f(x,\theta)\to\infty$ when $\|{x}\|\to\infty$), or (by Weierstrass’ theorem [46]) when the search domain can be restricted to a compact set without altering the optimizers.
+Continuity of the critical trajectories can also be ensured under standard assumptions: for example, by Berge’s theorem [46], by requiring that $f(x,\theta)$ is continuous in $\theta$.
+In our analysis, we will use a critical point with $\theta$ at rest, which is a constant vector $x^{\star}_{\circ}\in{}^{n}$, defined implicitly as:
+$0=\nabla_{x}f(x^{\star}_{\circ},0).$ (6)
+We assume that $x^{\star}_{\circ}$ exists and is locally unique [^1] Existence and local uniqueness of $x^{\star}_{\circ}$ can be guaranteed, for instance, under the assumptions of the Implicit Function Theorem [47].
+Namely, let $X_{\circ}\times\Theta_{\circ}$ be some neighborhood of $\left(x_{\mathrm{o}}^{\star},0\right)$, $x^{\star}_{\circ}$ exists and is locally unique when: (i) $f$ is $C^{1}$ on $X_{\circ}\times\Theta_{\circ}$, (ii) $x\mapsto f(x,\theta)$ is $C^{2}$ on $X_{\circ}$ for each $\theta\in\Theta_{\circ}$, and (iii) the $\left.\nabla_{xx}^{2}f(x,\theta)\right|_{x=x_{\circ}^{\star},\theta=0}$ is locally positive definite..
+Remark 3 (Critical trajectory vs. critical point at rest).
+We stress that $x^{\star}(t)$ and $x^{\star}_{\circ}$ are distinct quantities.
+Unlike the critical trajectory (defined in (5)), which denotes the optimal input to be tracked by the optimization algorithm, the critical point with $\theta$ at rest (defined in (6)) is a constant vector that need not be tracked and may never coincide with $x^{\star}(t)$$\square$
+We conclude this discussion by reviewing two important classes of optimization methods commonly used to solve (1).
+Remark 4 (Basic gradient flow algorithms).
+The basic gradient flow algorithm [48, 26] for (1) reads as:
+$\dot{x}(t)=-\eta\nabla_{x}f(x(t),\theta(t)),$ (7)
+where $\eta>0$ models the algorithm’s step size.
+It is known [48] that this class of algorithms is capable of computing (asymptotically) critical points of (1) only when $\theta(t)$ is a constant signal; in all other cases, the algorithm converges to a neighborhood of a critical point and convergence is inexact.
+We refer the reader to Remark 11 for an insightful interpretation of this fundamental limitation.
+$\square$
+Remark 5 (Prediction-correction algorithms).
+An established family of methods to solve (1) is that of prediction-correction algorithms [10, 11, 34, 36, 37]. Tailored to our setting, the basic method of this class reads as:
+[…]
+where $\nabla_{xx}f(x,\theta)$ is the Hessian of $f(x,\theta)$, and $\nabla_{x\theta}f(x,\theta)$ is the partial derivative of $\nabla_{x}f(x,\theta)$ with respect to $\theta$.
+Under strong convexity assumptions, this algorithm is known to converge exponentially, with zero asymptotic error, to a critical trajectory [10, 11].
+Variations of this algorithm have also been proposed to achieve fixed-time or finite-time convergence [34].
+We conclude by noting that implementing (8) requires (i) the cost to be twice continuously differentiable and strongly convex, (ii) knowledge of the maps $s(\theta)$, $\nabla_{x}f(x,\theta)$, $\nabla_{xx}f(x,\theta)$, and $\nabla_{x\theta}f(x,\theta)$, and (iii) knowledge of $\theta(t).$square
+2.2 Problem statement
+In line with the literature [26, 10, 11, 34, 12], we focus on gradient-type algorithms for solving (1), which are algorithms that have access to first-order oracles of the cost; that is, functional evaluations of the map:
+$(t,x)\mapsto\nabla_{x}f(x,\theta(t)).$ (9)
+Remark 6 (Evaluating the gradient).
+In the optimization literature [26], functional evaluations of (9) are generally obtained through two main approaches: (i) the algorithm has access to both the analytical form of $\nabla_{x}f(x,\theta)$ and $\theta(t)$ at each time $t$ (either because this signal is known or measurable), or (ii) an oracle returns evaluations of (9) (computed, e.g., through numerical differentiation).
+These alternatives will be discussed more in detail in (O1)–(O4), shortly below.
+$\square$
+We let the optimization algorithm be described by a dynamic internal state $z(t),$ taking values on an open subset $\mathcal{Z}\subseteq{}^{n_{c}},n_{c}\in\mathbb{N}_{>0}$.
+The optimization algorithm generates a sequence of points $x(t)\in{}^{n}$ (called exploration signal), and has access to functional evaluations of (9) at these points (called gradient feedback signal):
+$y(t)=\nabla_{x}f(x(t),\theta(t)).$ (10a)
+Together with this gradient feedback signal, the optimization algorithm is described by [^2] While one could allow $G_{c}$ to also depend on $y$, so that $x(t)=G_{c}(z(t),y(t))$, we will show in Section 4 that this is unnecessary.
+$\dot{z}(t)$ $=F_{c}(z(t),y(t)),$ $x(t)$ $=G_{c}(z(t)),$ (10b)
 where $F_{c}:\mathcal{Z}\times{}^{n}\rightarrow{}^{n_{c}}$ and $G_{c}:\mathcal{Z}\rightarrow{}^{n}$ are functions to be designed.
 In the remainder, we refer to (10) as a dynamic gradient-feedback optimization algorithm.
-The architecture of the studied gradient-feedback algorithm is illustrated in Figure 1.
-Notice that the dynamics of the optimization algorithm (4), coupled with the time-variability generator (2), have the form of a nonlinear autonomous system:
-$\dot{z}(t)$ $=F_{c}(z(t),y(t)),$
-$y(t)$ $=\nabla_{x}f(G_{c}(z(t)),\theta(t)),$
-$\dot{\theta}(t)$ $=s(\theta(t)).$ (5)
-Our objective is to design $F_{c}(z,y),$ $G_{c}(z),$ and $n_{c},$ so that $y(t)\rightarrow 0$ as $t\rightarrow\infty,$ which ensures that $x(t)$ tracks, with zero asymptotic error, a critical trajectory $x^{\star}(t)$ of (1).
-We will assume that $\theta(t)$ takes values in a neighborhood of the origin, and thus let $\Theta$ be some neighborhood of the origin of ${}^{p}.$ Note that there is no loss of generality in doing so, because if $\theta(t)$ takes values in the neighborhood of any other point, the former can be shifted to the origin via a change of variables without altering the critical points of (1).
-In what follows, we will denote by $x^{\star}_{\circ}\in{}^{n}$ a point such that
-$0=\nabla_{x}f(x^{\star}_{\circ},0),$ (6)
-and assume that $x^{\star}_{\circ}\in{}^{n}$ is locally unique.
-Moreover, we will let $z^{*}_{\circ}$ be such that $x_{\circ}^{\star}=G_{c}(z^{\star}_{\circ}),$ and assume that $z^{*}_{\circ}$ exists and is locally unique.
-Definition 1.
-We say that (4) asymptotically tracks a critical trajectory of (1) if, for each initial condition $(z(0),\theta(0))$ in some neighborhood of $(z_{\circ}^{*},0)\in\mathcal{Z}\times\Theta$, the solution of (II-A) satisfies $y(t)\to 0$ as $t\to\infty$.
+The architecture of the presented gradient-feedback algorithm is illustrated in Fig.
+1.
+For simplicity of the presentation, we will require that $F_{c}(z,y)$ and $G_{c}(z)$ are such that [^3] Notice that this is without loss of generality since $F_{c}(z,y)$ and $G_{c}(z)$ are to be designed and $x_{\circ}^{\star}$ is known through (6).:
+00 0 $=F_{c}(z_{\circ}^{\star},0)\qquad\text{and}\qquad x_{\circ}^{\star}=G_{c}(z_{\circ}^{\star}),$ (11)
+for some locally unique $z_{\circ}^{\star}\in{}^{n_{c}}$. Together with (6), this ensures that $z_{\circ}^{\star}$ is an equilibrium of (10) when the gradient feedback signal $y(t)$ is equal to zero.
+The dynamics of algorithm (10), coupled with the exosystem (2), have the form of a nonlinear autonomous system:
+$\dot{z}(t)$ $=F_{c}(z(t),y(t)),$ (12a)
+$y(t)$ $=\nabla_{x}f(G_{c}(z(t)),\theta(t)),$ (12b)
+$\dot{\theta}(t)$ $=s(\theta(t)).$ (12c)
+Definition 1 (Exact asymptotic tracking).
+We say that (10) exactly asymptotically tracks a critical trajectory of (1) if there exists a neighborhood of $(z_{\circ}^{*},0)$ in $\mathcal{Z}\times\Theta$ such that, for each initial condition $(z(0),\theta(0))$ in the neighborhood, the solution of (12) satisfies $y(t)\to 0$ as $t\to\infty$.
 $\square$
-Notice that, when an algorithm asymptotically tracks a critical trajectory, $x(t)\rightarrow x^{\star}(t)$ for some critical trajectory $x^{\star}(t);$ namely, critical point is readily given by the exploration signal.
-We make the objective of this work formal as follows.
-Problem 1 (Dynamic gradient-feedback problem).
-Find necessary and sufficient conditions (in terms of the loss function $f$) for the existence of a gradient-feedback optimization algorithm that asymptotically tracks a critical trajectory of (1).
-Derive a method to design such an algorithm when these conditions hold.
+By considering the entire class of optimization algorithms as in (10), and without requiring any assumption in addition to (9), we begin by posing the following fundamental question.
+Problem 0 (Minimal knowledge for exact asymptotic tracking).
+Consider the class of optimization algorithms (10).
+Determine the minimal necessary knowledge (beyond (9)), concerning the exosystem (2) and the optimization (1), needed to design an algorithm from this class that tracks, with zero asymptotic error, the critical trajectories of (1).
 $\square$
-It is important to note that existence conditions for $F_{c}(z,y)$ and $G_{c}(z)$ shall not depend on the temporal variability of the optimization (namely, $s(\theta)$), but only on the properties of the optimization problem.
-Conversely, if such conditions were to depend on $s(\theta)$, a more general class of optimization algorithms than (4) could be constructed, implying the formulation (4) would not be sufficiently general.
-II-B Standing assumptions
-We list hereafter the basic assumptions on which our approach to gradient-feedback theory is based.
-Assumption 3 (Stability of the exosystem).
-The point $\theta=0$ is a stable equilibrium of (2), and there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of 00 with the property that each initial condition $\theta(0)\in\Theta_{\circ}$ is Poisson stable [^1] Recall that, for a nonlinear system of the form (2), an initial condition $\theta^{\circ}$ is said to be Poisson stable if the solution starting from $\theta^{\circ}$ at time $t$ of (2) (here denoted by $\theta_{t}\left(\theta^{\circ}\right)$)) is defined for all $t\in\mathbb{R}_{\geq 0}$ and, for every neighborhood $U$ of $\theta^{\circ}$ and each $T\in\mathbb{R}_{>0}$, there exists a time $t_{1}>T$ such that $\theta_{t_{1}}(\theta^{\circ})\in U$, and a time $t_{2}<-T$ such that $\theta_{t_{2}}(\theta^{\circ})\in U$..
+Before addressing this question, we first outline various pieces of information on the exosystem and objective that the algorithm may (or may not) be able to access.
+Two (non mutually-exclusive [^4] Notice that requiring both (E1) and (E2) is equivalent to asking that both $s(\theta)$ as well as $\theta(0)$ are known.) assumptions on the exosystem from the literature (see, e.g., [48, 26, 10, 11, 34, 12]) are:
+  1. (E1)
+    The parameter signal $\theta(t)$ is known or measurable at each $t\in\mathbb{R}_{\geq 0},$ but no knowledge of $s(\theta)$ is available.
+  2. (E2)
+    The vector field $s(\theta)$ is known, but $\theta(0)$ (and hence also $\theta(t)$) is unknown.
+A second distinction can be made in terms of what knowledge concerning the objective function is available.
+In this case, several (non mutually-exclusive) options can be found in the literature (see, e.g., [48, 26, 10, 11, 34, 12]):
+  1. (O1)
+    The algorithm has access to oracle evaluations of (9).
+  2. (O2)
+    The analytical form of the gradient $\nabla_{x}f(x,\theta)$ is known.
+  3. (O3)
+    The sensitivity $\nabla_{x\theta}f(x,\theta)$ of the cost with respect to the parameter is known.
+  4. (O4)
+    The Hessian $\nabla_{xx}^{2}f(x,\theta)$ of the cost is known.
+  5. (O5)
+    The objective $f(x,\theta)$ is quadratic in both $x$ and $\theta$.
+Table 1: Summary of required knowledge and known guarantees by various established approaches.
+The symbol ()($\checkmark$)( ✓ ) is used when (O1) follows from (E1) and (O2).
+See Remarks 4, 5, and 7 for discussions.
+Approximate (E1) (E2) (O1) (O2) (O3) (O4) (O5)
+solutions $\theta(t)$ $s(\theta)$ $(t,x)\mapsto\nabla_{x}f(x,\theta(t))$ $\nabla_{x}f(x,\theta)$ $\nabla_{x\theta}f(x,\theta)$ $\nabla_{xx}^{2}f(x,\theta)$ $f$ quadratic
+Basic gradient flow [48, 26]
+Control-based methods [12]
+Prediction-correction [10, 11, 34] (✓)
+This work (Section 3) (✓)
+This work (Section 4)
+These are not assumptions we impose a priori in this work, but rather widely used requirements in the literature [48, 26, 10, 11, 34, 12].
+Various methods have been developed in the literature based on different combinations of these assumptions: for example, under (E1) and (O1), the basic gradient flow algorithm (7) converges (asymptotically) to a neighborhood of the critical points (see Remark 4).
+Under assumptions (E1), (E2), (O2), (O3), and (O4), the prediction-correction method (8) converges (asymptotically or in fixed-time) exactly to the optimizers (see Remark 5).
+Table 1 provides a comparison of the assumptions required by established methods and their positioning with respect to this work.
+Remark 7 (Generality of the algorithm class).
+The gradient flow algorithm (7) can be viewed as a special case of (10) with $F_{c}(z,y)=-\eta y$ and $G_{c}(z,y)=z$.
+Similarly, the prediction-correction method (8) can be viewed as a special case of (10) with $z=(z_{1},z_{2})$ and
+[…]
+$G_{c}(z)$ $=z_{1},$ (13)
+with $z_{2}(0)=\theta(0).$ It follows that (10) is general enough to encompass these two established methods as special cases, and thus all conclusions concerning the class of algorithms (10) made here will also apply to these methods.
 $\square$
-[…]
-For convenience, we will require that $F_{c}(z^{\star}_{\circ},0)=0.$ This ensures that the optimization algorithm (4) has an equilibrium at $z=z^{\star}_{\circ},$ and the corresponding gradient feedback signal is identically zero: $y(t)=\nabla_{x}f(x_{\circ}^{\star},0)=0.$ Since $\theta(t)$ is unmeasurable, the temporal variability of the cost can only be evaluated through measurements of $y(t);$ hence, we make the following assumption.
-Assumption 4 (Detectability of the exosystem).
-Let
-$Q:=\left[\frac{\partial\nabla_{x}f}{\partial\theta}\right]_{(x,\theta)=(x_{\circ}^{\star},0)},$ $S:=\left[\frac{\partial s}{\partial\theta}\right]_{\theta=0}.$ (7)
-The pair $(Q,S)$ is detectable. $\square$
-Detectability ensures that all the modes of the exosystem (2) can be reconstructed from the available measurements $y(t).$ Undetectability of $(Q,S)$ corresponds to a redundant description of the exogenous signal: indeed, if some modes of the exosystem do not influence the gradient, then $y(t)$ would be independent of those modes and they can thus be removed without altering the problem.
-Example 1 (Detectability).
-Suppose the objective function admits the decomposition $f(x,\theta(t))=\hat{f}(x)+x^{\mathsf{T}}\theta_{1}(t),$ where $\theta_{1}(t)=\alpha\cos(\omega t)$ for some $\alpha$ and $\omega$.
-The parameter signal can be generated by (2) with initial condition $\theta(0)=(\alpha,0)$ and vector field:
-$s(\theta)=\begin{bmatrix}0&1\\ -\omega^{2}&0\end{bmatrix}\theta.$
-Notice that, although the gradient signal $y(t)=\nabla_{x}\hat{f}(x)+\theta_{1}(t)$ is independent of $\theta_{2}(t),$ both modes of the exosystem affect the gradient (indeed, $\theta_{1}(t)=\alpha\cos(\omega t)$ cannot be generated by a linear system with a single mode).
-Notice also that the pair $(Q,S)$ is detectable.
+Our second objective is to design optimization algorithms from the class (10) that track, with zero asymptotic error, the critical trajectories of (1). We formalize this notion next.
+Problem 1 (Design of dynamic gradient-feedback optimization algorithms).
+Design $F_{c}(z,y),$ $G_{c}(z),$ and $n_{c},$ requiring the minimal necessary knowledge for exact asymptotic tracking (as in Problem 2.2), so that (10) exactly asymptotically tracks a critical trajectory of (1).
 $\square$
-III The parameter-feedback problem
-To address Problem 1, we first study a simpler problem that allows us to derive the necessary framework to tackle our objectives in their generality.
-To this end, in place of the dynamic optimization algorithm (4), we begin by considering an algebraic optimization algorithm of the form:
+3 The parameter-feedback problem
+In many cases of interest [26], having access to $y(t)$ is a byproduct of having access to both the function $\nabla_{x}f(x,\theta)$ and knowledge or measurements of $\theta(t)$; see Remark 6.
+In these cases, assumptions (E1) and (O2) are automatically satisfied.
+In this section, we analyze this scenario.
+We anticipate that the results derived in this section will also serve as an intermediate step to tackle the more challenging problem where assumption (E1) is relaxed, which is the focus of Section 4.
+When the algorithm has access to both $\theta(t)$ at each time $t\in\mathbb{R}_{\geq 0},$ and the function $\nabla_{x}f(x,\theta)$, the measurements $y(t)$ do not provide additional information.
+Hence, under the assumptions of this section, the optimization algorithm (10) can be replaced by the algebraic relationship [^5] While one could consider a dynamic optimization algorithm of the form $\dot{z}(t)=F_{c}(z(t),\theta(t))$ and $x(t)=G_{c}(z(t)),$ we will prove in Theorem 1 that a dynamic structure is unnecessary.
 $x(t)=H_{c}(\theta(t)),$ (14)
-where $H_{c}:\Theta\rightarrow{}^{n}$ is a mapping to be designed; we will require that $H_{c}$ is of class $C^{0}$ and satisfies the fixed-point condition $x^{\star}_{\circ}=H_{c}(0).$ Because of the explicit dependence on $\theta(t),$ we will refer to (8) to as a parameter-feedback optimization algorithm.
-Our objective is to design the map $H_{c}$ so that the composition of (2), (4b), and (8):
-$y(t)$ $=\nabla_{x}f(H_{c}(\theta(t)),\theta(t)),$
-$\dot{\theta}(t)$ $=s(\theta(t)),$ (9)
-tracks, with zero asymptotic error, the critical trajectories of (1). For the framework considered in this section, Problem 1 is reformulated as follows.
-Problem 2 (Static parameter-feedback problem).
-Find necessary and sufficient conditions (in terms of the loss function $f$) for the existence of a parameter-feedback optimization algorithm that asymptotically tracks a critical trajectory of (1).
-Derive a method to design such an algorithm when these conditions hold.
+where $H_{c}:\Theta\rightarrow{}^{n}$ is a mapping to be designed.
+Because of the explicit dependence on $\theta(t),$ we will refer to (14) to as a parameter-feedback optimization algorithm.
+In analogy with (11), for simplicity of the presentation, in this section, we will impose that $H_{c}(\theta)$ satisfies:
+$x^{\star}_{\circ}=H_{c}(0),$
+and require that $H_{c}$ is of class $C^{0}.$
+The composition of (2), (10a), and (14) is given by:
+$y(t)$ $=\nabla_{x}f(H_{c}(\theta(t)),\theta(t)),$ (15a)
+$\dot{\theta}(t)$ $=s(\theta(t)).$ (15b)
+In analogy with Definition 1, we will say that the parameter-feedback algorithm (14) exactly asymptotically tracks a critical trajectory of (1) if there exists a neighborhood $\Theta_{s}\subset\Theta$ of the origin such that, for each initial condition $\theta(0)\in\Theta_{s}$, the solution of (15) satisfies $y(t)\to 0$ as $t\to\infty$.
+For the algorithm (14) considered in this section, Problems 2.2 and 1 are reformulated as follows.
+Problem 2 (Minimal knowledge for parameter feedback).
+Consider the class of optimization algorithms (14).
+Determine the minimal necessary knowledge concerning (1) and (2), needed to design an algorithm from this class that exactly asymptotically tracks a critical trajectory of (1).
 $\square$
-Solvability of the static parameter-feedback problem will depend on the existence of a function that zeros the gradient.
+Problem 3 (Design of parameter-feedback optimization algorithms).
+Design $H_{c}(\theta),$ requiring minimal knowledge for parameter-feedback (as in Problem 2), so that (14) exactly asymptotically tracks a critical trajectory of (1).  $\square$
+3.1 Fundamental results
+Solvability of the parameter-feedback problem will depend on the existence of a function that zeros the gradient on the set of limit points of the exosystem; we now define these notions.
 Definition 2 (Mapping zeroing the gradient).
@@
 00 0 $=\nabla_{x}f(H_{c}(\theta),\theta).$ (16)
-Moreover, we say that $H_{c}$ zeros the gradient globally if (10) holds for all $\theta\in\Theta$, and locally if (10) holds for all $\theta\in\Theta_{\circ},$ where $\Theta_{\circ}\subseteq{}^{p}$ is some neighborhood of the origin [^2] The neighborhoods in Defn.
-2 and Assumption 3 need not be the same..
-$\square$
-The following result characterizes all parameter-feedback optimization algorithms that achieve asymptotic tracking of a critical trajectory.
-Theorem 1 (Parameter-feedback algorithm characterization).
-Let Assumptions 1, 2, and 3 hold. The parameter-feedback algorithm (III) asymptotically tracks a critical trajectory of (1) if and only if the mapping $H_{c}$ zeros the gradient locally.
+Moreover, we say that $H_{c}$ zeros the gradient on a set $\Theta_{\circ}\subseteq\Theta$ if (16) holds for all $\theta\in\Theta_{\circ}$. $\square$
+Definition 3 (Limit point and limit set).
+A point $\theta_{\omega}\in\Theta$ is a limit point with respect to the initialization $\theta_{\circ}\in\Theta$ if there exists a sequence $\{t_{i}\}_{i\in\mathbb{N}_{\geq 0}}$ with $t_{i}\to\infty$ as ${i\to\infty}$ such that the exosystem (2) with $\theta(0)=\theta_{\circ}$ satisfies $\theta({t_{i}})\to\theta_{\omega}$ as ${i\to\infty}$.
+Let $\Omega(\theta_{\circ})$ denote the set of all limit points (i.e., for all sequences $t_{i}$) of (2) with respect to the initialization $\theta_{\circ}\in\Theta.$ Given $\Theta_{\circ}\subseteq\Theta,$ the set $\Omega(\Theta_{\circ}):=\cup_{\theta_{\circ}\in\Theta_{\circ}}\Omega(\theta_{\circ})$ is called the limit set with respect to initializations in $\Theta_{\circ}$ [49$\square$
+Intuitively, $\Omega(\Theta_{\circ})$ denotes the set of all limit points (equilibria, limit cycles, etc.) that can be reached by the exosystem when initialized at points in $\Theta_{\circ}$.
+By the boundedness trajectories assumption (Assumption 2), $\Omega(\Theta_{\circ})$ is a bounded set.
+The following result characterizes all parameter-feedback optimization algorithms that achieve asymptotic tracking.
+Theorem 1 (Existence and characterization of parameter-feedback algorithms).
+Let Assumptions 12 hold.
+The parameter-feedback algorithm (14) asymptotically tracks a critical trajectory of (1) if and only if there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin such that the mapping $H_{c}$ zeros the gradient on the limit set $\Omega(\Theta_{\circ})$.
 Proof.
-(Only if) Suppose $y(t)\to 0$ as $t\to\infty$.
-We will show that $H_{c}$ zeros the gradient locally.
-By Poisson stability of the exosystem (cf.
-Assumption 3), there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin such that, for every $\theta_{\circ}\in\Theta_{\circ},$ every $\varepsilon>0,$ and every $T>0,$ the trajectory of (III) satisfies
-$\|\theta(t)-\theta_{\circ}\|<\varepsilon,$
-at some $t>T$. Then, $\lim_{t\rightarrow\infty}y(t)=0$ can hold only if (10) holds for all $\theta\in\Theta_{\circ}$.
-(If) Suppose $H_{c}$ zeros the gradient on some neighborhood $\Theta_{0}\subset\Theta$ of the origin, i.e., (10) holds for all $\theta\in\Theta_{0}$.
-By Assumption 3, the point $\theta=0$ is a stable equilibrium of the exosystem, so there exists some other neighborhood $\Theta_{1}\subset\Theta$ of the origin and time $T>0$ such that $\theta(t)\in\Theta_{0}$ for all $t>T$ when $\theta(0)\in\Theta_{1}$.
-Since $H_{c}$ zeros the gradient on $\Theta_{0}$, this implies that $y(t)\to\infty$ as $t\to\infty$.
+(Only if) Suppose $y(t)\to 0$ as $t\to\infty$ for initializations $\theta(0)\in\Theta_{\circ}$; we will show that $H_{c}$ zeros the gradient on $\Omega(\Theta_{\circ})$.
+By Assumption 2, the trajectories of the exosystem are bounded, and thus, by the Bolzano–Weierstrass theorem, there exists an increasing subsequence $\{t_{i}\}_{i\in\mathbb{N}_{\geq 0}}$ such that $\theta(t_{i})$ converges to some limit point $\theta_{\omega}\in\Omega\left(\theta(0)\right)$.
+We then have:
+$\lim_{i\to\infty}y({t_{i}})=\lim_{i\to\infty}\nabla_{x}f(H_{c}(\theta({t_{i}})),\theta({t_{i}}))=\nabla_{x}f(H_{c}(\theta_{\omega}),\theta_{\omega}),$ (17)
+where the second identity follows by the continuity of the gradient (see Assumption 1) and that of $H_{c}$.
+Because, $y(t)\to 0$, the left-hand side of (17) is equal to zero, and thus (17) implies that $\nabla_{x}f(H_{c}(\theta_{\omega}),\theta_{\omega})=0.$ Since this holds for any limit point $\theta_{\omega}\in\Omega(\Theta_{\circ})$, the statement follows.
+(If) Suppose $\theta(0)\in\Theta_{\circ}$ and that $H_{c}$ zeros the gradient on $\Omega(\Theta_{\circ})$.
+Then, the right-hand side of (17) is equal to zero, which implies the existence of a sequence $t_{i}$ such that $y({t_{i}})\to 0$ as ${i\to\infty}$.
+Since this holds for any limit point $\theta_{\omega}\in\Omega(\theta(0))$, any convergent subsequence of $y(t)$ converges to zero.
+Moreover, $y(t)$ is bounded due to the Lipschitz continuity of the gradient, so $y(t)$ also converges to zero as $t\to\infty$.
+By iterating the reasoning for all $\theta(0)\in\Theta_{\circ},$ it follows that $y(t)\to 0$ for all $\theta(0)\in\Theta_{\circ}$, and the claim follows.
 ∎
-Remark 2.
-Poisson stability of the exosystem is needed for the gradient zeroing condition to be necessary for asymptotic tracking of a critical trajectory (but is not needed for sufficiency).
-This is because Poisson stability implies that all points sufficiently close to the origin are revisited infinitely often, so $H_{c}$ must zero the gradient on a neighborhood of the origin.
-More generally, $H_{c}$ must zero the gradient on all limit points of $\theta(t)$$\square$
-The result provides a full characterization of all parameter-feedback algorithms that ensure exact asymptotic tracking.
-In words, $x=H_{c}(\theta)$ is a parameter-feedback algorithm if and only if it zeros the gradient everywhere in a neighborhood of the critical point.
-As a byproduct, the result also provides a necessary and sufficient condition for the solvability of the parameter-feedback problem: the problem is solvable if and only if the set of solutions to the system of equations $0=\nabla_{x}f(x,\theta)$ can be expressed, everywhere in a neighborhood of the origin of $\Theta$, as the graph of a function $x=H_{c}(\theta).$ Finally, the theorem provides an explicit form for the desired parameter-feedback algorithm (8): this is given by $x(t)=H_{c}(\theta(t));$ this algorithm ensures that, for all initial states $\theta(0)$ sufficiently close to the origin, $y(t)\rightarrow 0$ as $t\rightarrow\infty.$ In this sense, Theorem 1 provides a complete answer to Problem 2.
+The theorem shows that the existence of a parameter-feedback algorithm is dependent upon the existence of a mapping that zeros the gradient.
+An application of this theorem is illustrated in Example 1, which shows that a parameter-feedback algorithm may fail to exist in some circumstances.
+The theorem also provides a full characterization of all parameter-feedback algorithms that achieve asymptotic tracking.
+Loosely speaking, $x=H_{c}(\theta)$ is a parameter-feedback algorithm if and only if it zeros the gradient on the limit set of the exosystem.
+As a byproduct, the result also provides a necessary and sufficient condition for the solvability of the parameter-feedback problem: the problem is solvable if and only if the set of solutions to the system of equations $0=\nabla_{x}f(x,\theta)$ can be expressed, at all limit points of the exosystem, as the graph of a function $x=H_{c}(\theta).$ We present a set of sufficient conditions for this to hold in Section 3.2.
+By Theorem 1, the problem of designing a parameter-feedback algorithm for exact asymptotic tracking can be reduced to that of finding a mapping $x=H_{c}(\theta)$ that zeros the gradient.
+Therefore, we have the following.
+Answer to Problem 2.
+When a parameter-feedback algorithm that achieves exact tracking exists, the minimal knowledge needed to design such an algorithm is (O2).
+Notice that assumption (E1) is implicitly required to execute the algorithm (14).
 We illustrate the applicability of the result and the necessity of the provided condition in the following example.
 Figure 2: Investigation of the condition (16).
-(Left) Loss function $f(x,\theta)$ studied in Example 2, plotted for $\theta=0.$ (Right) Gradient $\nabla f(x,\theta).$ The function $f(x,0)$ admits two critical points: $x^{\star}_{\circ,1}=\frac{1}{\sqrt{3}}$ and $x^{\star}_{\circ,2}=-\frac{2}{\sqrt{3}}.$ At $x^{\star}_{\circ,1},$ condition (10) is not satisfied, since with an an upward shift of the graph of $\nabla_{x}f(x,0),$ $x^{\star}_{\circ,1}$ is no longer a critical point of $f(x,0).$ On the other hand, (10) holds for $x^{\star}_{\circ,2},$ since $x^{\star}_{\circ,2}$ varies continuously as $\theta$ is perturbed.
+(Left) Loss function $f(x,\theta)$ studied in Example 1, plotted for $\theta=0.$ (Right) Gradient $\nabla f(x,\theta).$ The function $f(x,0)$ admits two critical points: $x^{\star}_{\circ,1}=\frac{1}{\sqrt{3}}$ and $x^{\star}_{\circ,2}=-\frac{2}{\sqrt{3}}.$ At $x^{\star}_{\circ,1},$ condition (16) is not satisfied, since with an upward shift of the graph of $\nabla_{x}f(x,0),$ $x^{\star}_{\circ,1}$ is no longer a critical point of $f(x,0).$ On the other hand, (16) holds for $x^{\star}_{\circ,2},$ since $x^{\star}_{\circ,2}$ varies continuously as $\theta$ is perturbed.
 See Example 1 for a discussion.
@@
 2(b)) shifts upward and the equation $\nabla_{x}f(x,\epsilon)=0$ no longer admits a solution in a neighborhood of $x^{\star}_{\circ,1}.$
-On the other hand, for the critical point $x^{\star}_{\circ,2},$ any arbitrarily-small upward or downward shift of the graph of $\nabla_{x}f(x,0)$ results in a continuous perturbation of $x^{\star}_{\circ,2}$ (see Fig.
+On the other hand, for the critical point $x^{\star}_{\circ,2},$ any arbitrarily small upward or downward shift of the graph of $\nabla_{x}f(x,0)$ results in a continuous perturbation of $x^{\star}_{\circ,2}$ (see Fig.
 2(b)), thus suggesting existence of $x=H_{c}(\theta)$ as in (16).
@@
 $\square$
-Existence of a map $x=H_{c}(\theta)$ as in (10) can be ensured for general problems with the aid of the implicit function theorem [32], as illustrated next.
+We conclude by observing that the proof of Theorem 1 is constructive in that it provides an explicit procedure to construct parameter-feedback algorithms.
+Answer to Problem 3.
+When a parameter-feedback algorithm exists, a procedure to design it requiring minimal knowledge (O2) is to first determine the limit set $\Omega(\Theta_{\circ})$ for the given set of initial conditions $\Theta_{\circ}$ and then find a mapping $H_{c}(\theta)$ that zeros the gradient on the limit set, in which case the parameter-feedback algorithm is given by (14).
+We conclude by illustrating the parameter-feedback design procedure on a quadratic problem.
+Example 2.
+Consider an instance of (1) with quadratic cost and time-variability that depends linearly on $\theta(t)$ (which, e.g., has been investigated in [12]):
+$f(x(t),\theta(t))=\tfrac{1}{2}x(t)^{\mathsf{T}}Rx(t)+x(t)^{\mathsf{T}}Q\theta(t),$ (19)
+with matrices $R\in\mathbb{S}^{n}$ and $Q\in{}^{n\times p}$. This loss is convex and Lipschitz smooth and thus satisfies Assumption 1. In this case, the signal we wish to regulate to zero is:
+$y(t)=\nabla_{x}f(x(t),\theta(t))=Rx(t)+Q\theta(t).$
+For arbitrary $\theta,$ this problem admits a critical point if and only if $\operatorname{Im}Q\subseteq\operatorname{Im}R$, in which case $x_{\circ}^{\star}$ is unique.
+Applying Theorem 1 amounts to finding a linear transformation $H_{c}\in{}^{n\times p}$ such that $0=(RH_{c}+Q)\theta$ for all $\theta$ in a neighborhood of the origin.
+Assuming $\operatorname{Im}Q\subseteq\operatorname{Im}R$, we can choose $H_{c}=-R^{\dagger}Q$, where $R^{\dagger}$ is the pseudo-inverse of $R$.
+Note that, by substituting into (15), we have ${y(t)=RH_{c}\theta(t)+Q\theta(t)=0}$ for all times $t\in\mathbb{R}_{\geq 0}.$ Namely, the gradient is identically zero.
+This property holds because the mapping $H_{c}(\theta)$ derived here zeros the gradient globally (and not only in some neighborhood of the origin).
+In fact, this feature holds for the any optimization problem for which the gradient of the cost admits a mapping that zeros the gradient globally.
+A set of sufficient conditions to check for this property are given in Theorem 3.
+$\square$
+Remark 8 (Prediction-correction algorithms asymptotically compute maps that zero the gradient).
+As an illustration, consider the quadratic problem studied in Example 2.
+Let $x_{\text{PC}}(t)$ denote the iterates of the prediction-correction algorithm (8) applied to the quadratic objective (19), and let $x_{\text{PF}}(t)=H_{c}\theta(t)$ with $H_{c}=-R^{\dagger}Q$ be the parameter-feedback algorithm.
+Then, the error $e(t)=x_{\text{PC}}(t)-x_{\text{PF}}(t)$ satisfies the dynamics $\dot{e}(t)=-e(t)$, indicating that the iterates of prediction-correction converge exponentially to those of the parameter-feedback algorithm, which zeros the gradient.
+$\square$
+3.2 Sufficient conditions for the existence of a parameter-feedback algorithm
+Existence of a map $x=H_{c}(\theta)$ as in (16) can be ensured for general problems with the aid of the implicit function theorem [47], as illustrated by the following result.
 Proposition 2 (Existence of local parameter-feedback algorithms).
-Let Assumptions 1, 2, and 3 hold, and let $X_{\circ}\times\Theta_{\circ}$ be some neighborhood of $(x_{\circ}^{\star},0).$ Further, assume that:
+Let Assumptions 12 hold, and let $X_{\circ}\times\Theta_{\circ}$ be some neighborhood of $(x_{\circ}^{\star},0).$ Further, assume that:
   1. (i)
@@
     the Hessian $\left.\nabla^{2}_{xx}f(x,\theta)\right|_{x=x^{\star}_{\circ},\theta=0}$ is positive definite.
-Then, there exits a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient locally. $\square$
+Then, there exists a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient on some neighborhood of the origin of $\Theta$. $\square$
 Proof.
@@
 $\det\left[\frac{\partial F(x,\theta)}{\partial x}\right]_{(x,\theta)=(x_{\circ}^{\star},0)}\neq 0.$
-Then, the result follows from the implicit function theorem (see, e.g., [33, Thm.
+Then, the result follows from the implicit function theorem (see, e.g., [50, Thm.
 1]) applied to the first-order optimality conditions $0=F(x,\theta)$ at the point $(x,\theta)=(x^{\star}_{\circ},0)$.
 ∎
-Intuitively, when the loss function is twice continuously differentiable in the decision variable with a positive definite Hessian, every critical point is a local isolated minimum and existence of $H_{c}(\theta(t))$ is guaranteed by the implicit function theorem.
-Notice that, differently from the characterization in Theorem 1, the conditions offered by Proposition 2 are sufficient but not necessary.
-Although the conditions are immediate to verify, the convergence claims of Theorem 1 and Proposition 2 are of local nature, namely, $y(t)\rightarrow 0$ is ensured provided that $\theta(0)$ is sufficiently close to the origin.
+In other words, Proposition 2 states that when the loss function is sufficiently smooth and the Hessian is locally positive definite, the existence of $H_{c}(\theta(t))$ is guaranteed by the implicit function theorem.
+Notice that these conditions are sufficient but not necessary.
+Remark 9 (Computing mappings that zero the gradient).
+Under the assumptions of Proposition 2, the implicit function theorem gives that the linear function
+[…]
+is a first-order approximation of a mapping that zeros the gradient on a neighborhood of the origin. $\square$
+Although the conditions in Proposition 2 are immediate to verify, the convergence claims of Theorem 1 and Proposition 2 are of local nature, namely, $y(t)\rightarrow 0$ is ensured provided that $\theta(0)$ is sufficiently close to the origin.
 The following result provides a sufficient condition for global convergence.
 Theorem 3 (Existence of global parameter-feedback algorithms).
-Let Assumptions 1, 2, and 3 hold. Further, assume:
+Let Assumptions 12 hold. Further, assume:
   1. (i)
@@
     the mappings $\nabla_{xx}^{2}f$ and $\frac{\partial\nabla_{xx}^{2}f}{\partial x}$ are $C^{0}$ on ×${}^{n}\times\Theta.$
-Then, there exits a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient globally. $\square$
+Then, there exits a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient everywhere in $\Theta.$square
 Proof.
@@
 and the mappings $\frac{\partial F}{\partial x}$ and $\frac{\partial^{2}F}{\partial x^{2}}$ are $C^{0}$ in ×${}^{n}\times\Theta$.
-Hence, assumptions (B1)–(B5) of a global version of the implicit function theorem [33, Thm.
-2] (see also [34, Thm.
+Hence, assumptions (B1)–(B5) of a global version of the implicit function theorem [50, Thm.
+2] (see also [51, Thm.
 6]) are satisfied for the first-order optimality conditions $0=F(x,\theta)$, and the claim follows.
 ∎
-Theorem 3 shows that, under additional continuity assumptions on the loss function (cf.
-conditions (i), (ii), and (iv)) and when the Hessian of the loss is positive definite everywhere, convergence of (III) is ensured globally.
-More formally, under the assumptions of Theorem 3, the parameter-feedback algorithm $x(t)=H_{c}(\theta(t))$ ensures that $y(t)\rightarrow 0$ as $t\rightarrow\infty$ for all initial conditions $\theta(0)\in\Theta$.
-This result is particularly relevant, as it ensures the existence of parameter-feedback algorithms for the (very important in practice) class of strictly convex loss functions (studied in, e.g., [11, 12]).
-We conclude this section by illustrating the design procedure for parameter-feedback algorithms on a quadratic problem.
-Example 3.
-Consider an instance of (1) with quadratic cost and time-variability that depends linearly on $\theta(t)$ (which has been investigated in [12]):
-$f(x(t),\theta(t))=\tfrac{1}{2}x(t)^{\mathsf{T}}Rx(t)+x(t)^{\mathsf{T}}Q\theta(t),$ (12)
-with matrices $R\in\mathbb{S}^{n}$ and $Q\in{}^{n\times p}$.
-In this case, the signal we wish to regulate to zero is: $y(t)=\nabla_{x}f(x(t),\theta(t))=Rx(t)+Q\theta(t).$ For arbitrary $\theta,$ this problem admits a critical point if and only if $\operatorname{Im}Q\subseteq\operatorname{Im}R$, in which case $x_{\circ}^{\star}$ is unique.
-Applying Theorem 1 amounts to finding a linear transformation $H_{c}\in{}^{n\times p}$ such that $0=(RH_{c}+Q)\theta$ for all $\theta$ in a neighborhood of the origin.
-Assuming $\operatorname{Im}Q\subseteq\operatorname{Im}R$, we can choose $H_{c}=-R^{\dagger}Q$, where $R^{\dagger}$ is the pseudo-inverse of $R$.
-Note that, by substituting into (III), we have
-$y(t)=RH_{c}\theta(t)+Q\theta(t)=0,\quad\forall t\in\mathbb{R}_{\geq 0}.$
-Namely, the gradient is identically zero at all times.
-We also note that this is a particular feature of the parameter feedback problem, which can be achieved because $\theta(t)$ is assumed to be known when implementing (8).
-As we will see shortly below (cf.
-Section IV), $y(t)=0$ can be achieved only as $t\rightarrow\infty$ for the gradient-feedback problem.
-We conclude by noting that $y(t)=0$ for all $t\in\mathbb{R}_{\geq 0}$ holds for all initial conditions $\theta(0)$ because the mapping $H_{c}(\theta)$ derived here zeros the gradient globally (and not only in some neighborhood of the origin).
-See also Theorem 3.
-$\square$
-IV The dynamic gradient-feedback problem
-In this section, we will build on the framework derived for the parameter-feedback problem to address the main objectives of this work: Problem 1.
-IV-A Fundamental results
-We begin with the following instrumental characterization.
+Theorem 3 shows that, under additional continuity assumptions on the loss function and when the Hessian of the loss is positive definite everywhere, there exists a mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient everywhere in $\Theta.$
+By combination of Theorem 3 and Theorem 1, it follows that under the assumptions of Theorem 3, the parameter-feedback algorithm $x(t)=H_{c}(\theta(t))$ ensures that $y(t)\rightarrow 0$ as $t\rightarrow\infty$ for all initial conditions $\theta(0)\in\Theta$.
+This result ensures the existence of parameter-feedback algorithms for the particular class of strictly convex loss functions [11, 12].
+4 The dynamic gradient-feedback problem
+In the previous section, we analyzed algorithms assuming knowledge of the function $\nabla_{x}f(x,\theta)$ and of the signal $\theta(t),t\in\mathbb{R}_{\geq 0}$ (cf.
+assumptions (E1) and (O2)).
+In this section, we relax these assumptions and require only that $y(t)$ is available through oracle evaluations of (9).
+For this reason, we will shift our attention back to the general class of dynamic gradient-feedback algorithms (10).
+Because $\theta(t)$ is unmeasurable by the algorithm, the temporal variability of the cost can only be evaluated through measurements of $y(t);$ for this reason, we impose the following.
+Assumption 3 (Detectability of the exosystem).
+Let
+$Q:=\left[\frac{\partial\nabla_{x}f}{\partial\theta}\right]_{(x,\theta)=(x_{\circ}^{\star},0)},$ $S:=\left[\frac{\partial s}{\partial\theta}\right]_{\theta=0}.$ (21)
+The pair $(Q,S)$ is detectable. $\square$
+The detectability property of Assumption 3 is equivalent to the existence of a local exponential observer for all (non-asymptotically stable) modes of $\theta(t)$ based on measurements $y(t)$ [52, Cor.
+3.4].
+Undetectability corresponds to a redundant description of the exogenous signal: if some modes of $\theta(t)$ are undetectable, these modes do not influence the gradient (and, as a consequence, the critical points of (1)), and thus they can be removed from the optimization problem (1) without altering its critical points.
+In other words, Assumption 3 ensures that (2) does not contain redundancies.
+4.1 Existence and characterization of dynamic gradient feedback algorithms
+We begin by considering the class of algorithms (10), restricted to ensure local exponential stability, and we characterize a set of necessary and sufficient conditions for this class to achieve exact asymptotic tracking.
 Theorem 4 (Gradient-feedback algorithm characterization).
-Suppose Assumptions 1, 2, and 3 hold, and assume that $F_{c}(z,y)$ and $G_{c}(z)$ are such that the equilibrium $z=z^{\star}_{\circ}$ of
-$\dot{z}(t)$ $=F_{c}(z(t),\nabla_{x}f(G_{c}(z(t)),\theta(t))),$
-is exponentially stable.
-Then, the gradient-feedback optimization algorithm (II-A) asymptotically tracks a critical trajectory of (1) if and only if there exists a $C^{2}$ mapping $z=\sigma(\theta)$ with $\sigma(0)=z_{\circ}^{\star}$, defined in a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin, that satisfies the following for all $\theta\in\Theta_{\circ}$:
-$\frac{\partial\sigma(\theta)}{\partial\theta}s(\theta)$ $=F_{c}(\sigma(\theta),0),$ (13a)
-00 0 $=\nabla_{x}f(G_{c}(\sigma(\theta)),\theta).$ (13b)
+Suppose Assumptions 13 hold, and assume that $F_{c}(z,y)$ and $G_{c}(z)$ are such that the equilibrium $z=z^{\star}_{\circ}$ of
+$\dot{z}(t)$ $=F_{c}(z(t),\nabla_{x}f(G_{c}(z(t)),0)),$ (22)
+is locally exponentially stable.
+The optimization algorithm (12) exactly asymptotically tracks a critical trajectory of (1) if and only if there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin and a $C^{2}$ mapping $z=\sigma(\theta)$ with $\sigma(0)=z_{\circ}^{\star}$, such that:
+$\left.\frac{\partial\sigma(\theta)}{\partial\theta}\right|_{\theta=\theta_{\omega}}s(\theta_{\omega})$ $=F_{c}(\sigma(\theta_{\omega}),0),$ (23a)
+00 0 $=\nabla_{x}f(G_{c}(\sigma(\theta_{\omega})),\theta_{\omega}),$ (23b)
+at all limit points $\theta_{\omega}\in\Omega(\Theta_{\circ})$.
 Proof.
-(Only if). We first prove that $\lim_{t\rightarrow\infty}y(t)=0$ implies (13). The coupled dynamics (II-A) have the form:
-$\dot{z}$ $=(A_{c}+B_{c}RM)z+B_{c}Q\theta+\chi(x,\theta),$
-$\dot{\theta}$ $=S\theta+\psi(\theta).$ (14)
-for some mappings $\chi(x,\theta)$ and $\psi(\theta)$ that vanish at the origin along with their first-order derivatives, and where $Q$ and $S$ are defined in (7) and
+(Only if). We first prove that $\lim_{t\rightarrow\infty}y(t)=0$ implies (23). The coupled dynamics (12) have the form:
+$\dot{z}$ $=(A_{c}+B_{c}RM)z+B_{c}Q\theta+\chi(z,\theta),$
+$\dot{\theta}$ $=S\theta+\psi(\theta),$ (24)
+for some mappings $\chi(x,\theta)$ and $\psi(\theta)$ that vanish at the origin along with their first-order derivatives, with $Q$ and $S$ are defined in (21), and
 $A_{c}$ $=\left[\frac{\partial F_{c}}{\partial z}\right]_{(z,y)=(z_{\circ}^{\star},0)},$ $B_{c}$ $=\left[\frac{\partial F_{c}}{\partial y}\right]_{(z,y)=(z_{\circ}^{\star},0)},$
 $R$ $=\left[\frac{\partial\nabla_{x}f}{\partial x}\right]_{(x,\theta)=(x_{\circ}^{\star},0)},$ $M$ $=\left[\frac{\partial G_{c}}{\partial z}\right]_{z=z_{\circ}^{\star}}.$
-By assumption, the eigenvalues of the matrix $A_{c}+B_{c}RM$ are in $\mathbb{C}^{-}$ and those of $S$ on the imaginary axis.
-By Theorem 8, the system (IV-A) has a center manifold at (0,0): the graph of a mapping $z=\sigma(\theta),$ with $\sigma(\theta)$ satisfying (see (31))
-$\frac{\partial\sigma(\theta)}{\partial\theta}s(\theta)=F_{c}(\sigma(\theta),\nabla_{x}f(G_{c}(\sigma(\theta)),\theta)).$
-By Assumption 3, no trajectory on this manifold converges to zero, so $\lim_{t\rightarrow\infty}y(t)=0$ can hold only if (13b) holds, in which case the above equation reduces to (13a).
+By assumption, the eigenvalues of the matrix $A_{c}+B_{c}RM$ are in $\mathbb{C}^{-}$.
+By Theorem 9, the system (4.1) has a center manifold at $(z^{\star}_{\circ},0)$: the graph of a mapping $z=\sigma(\theta),$ with $\sigma(\theta)$ satisfying (see (41))
+$\sigma(s(\theta))=F_{c}(\sigma(\theta),\nabla_{x}f(G_{c}(\sigma(\theta)),\theta)).$
+Next, similarly to the parameter-feedback case (Theorem 1), by Assumption 2, $\theta(t_{i})$ converges to some limit point $\theta_{\omega}\in\Omega\left(\theta(0)\right)$ for some subsequence $\{t_{i}\}_{i\in\mathbb{N}_{\geq 0}}$.
+By continuity of the gradient (Assumption 1) and that of $G_{c}$,
+$\lim_{i\to\infty}y({k_{i}})=\lim_{i\to\infty}\nabla_{x}f(G_{c}(\theta_{t_{i}}),\theta_{t_{i}})=\nabla_{x}f(G_{c}(\theta_{\omega}),\theta_{\omega}).$ (25)
+Because, $y(t)\to 0$, the left-hand side of (25) is equal to zero, and thus $\nabla_{x}f(G_{c}(\theta_{\omega}),\theta_{\omega})=0.$ Since this holds for any limit point $\theta_{\omega}\in\Omega(\theta_{0})$, the claim follows.
 (If).
 We now prove that (23) implies ${\lim_{t\rightarrow\infty}y(t)=0}$.
-By Theorem 9, the center manifold $z=\sigma(\theta)$ is locally attractive; namely, $z(t)\rightarrow\sigma(\theta(t))$ as $t\rightarrow\infty.$ Then, the fulfillment of (13b) guarantees that $y(t)\rightarrow 0.$
-The two conditions in (13) fully characterize the class of optimization algorithms that achieve asymptotic tracking of a critical trajectory.
-In words, (4) tracks a critical trajectory if and only if, for some mapping $\sigma,$ the composite function $G_{c}\circ\sigma$ zeros the gradient locally (see (13b)), and the controller $F_{c}(z,y)$ is algebraically related to the exosystem $s(\theta)$ as given by (13a).
+It follows from (23) and Theorem 10 that the graph of the mapping $z=\sigma(\theta)$ (i.e., $(\sigma(\theta),\theta)$) is a center manifold for the coupled dynamics (12).
+Moreover, such manifold is locally attractive; namely, $z(t)\rightarrow\sigma(\theta(t))$ as $t\rightarrow\infty.$ Then, the fulfillment of (23b) guarantees that the right-hand side of (25) $y(t)\rightarrow 0.$ The conclusion then follows by iterating the (If) part of the proof of Theorem 1.
+∎
+The two conditions in (23) fully characterize the class of optimization algorithms that achieve exact asymptotic tracking: (10) tracks a critical trajectory if and only if, for some mapping $\sigma,$ the composite function $G_{c}\circ\sigma$ zeros the gradient locally (see (23b)), and the controller $F_{c}(z,y)$ is algebraically related to the exosystem $s(\theta)$ as given by (23a).
 Notice that, by Theorem 1, the former condition implies that
 $x(t)=G_{c}(\sigma(\theta)),$ (26)
-is a parameter-feedback optimization algorithm for (1).
+is a parameter-feedback optimization algorithm for (1). Finally, we note that the existence of an exponentially stable system as in (22) is always guaranteed under Assumption 3 [52, Sec. 3].
+The following interpretation follows from our findings.
 Remark 10 (The internal model principle).
-We interpret condition (13a) as the internal model principle of time-varying optimization, as it expresses the requirement that any optimization algorithm that achieves asymptotic tracking must include an internal model of the exosystem.
-Note that the use of a copy of the temporal variability of the optimization problem is explicit in the prediction-correction algorithm [10] (precisely, through the term $\nabla_{xt}f_{0}(x(t),t)$ – see Remark 1 for notation).
+We refer to condition (23a) as the internal model principle of time-varying optimization, as it encapsulates the requirement that an optimization algorithm must incorporate an internal model of the exosystem to achieve exact asymptotic tracking.
+Note that the use of a copy of the temporal variability of the optimization problem is explicit in the prediction-correction algorithm (see (7)).
 $\square$
-It is important to note that, by Theorem 4, the exosystem state $\theta$ and that of the optimization $z$ must be related, everywhere in $\Theta_{\circ},$ by the relationship:
+Theorem 4 allows us to re-interpret the basic gradient-flow algorithm as follows.
+Remark 11 (Internal model-based interpretation of basic gradient flow methods).
+Recall that the gradient-flow algorithm (4) can be viewed as an instance of (10a) with $F_{c}(z,y)$ and $G_{c}(z)$ as in Remark 7.
+By direct substitution into (23), it is immediate to see that this algorithm satisfies (23a) with $s(\theta)=0$ and $\sigma(\theta)$ arbitrary.
+Since $s(\theta)=0$ describes the internal model of a constant signal, it follows from Theorem 4 that these algorithms achieve exact asymptotic tracking if and only if $\theta(t)$ is a constant signal.
+This observation is in line with the inexact convergence properties of these algorithms typically given for these algorithms [48, 26, 14$\square$
+Since by, Theorem 4, the problem of exact asymptotic tracking can be reduced to that of finding an optimization algorithm that is algebraically equivalent to the exosystem and a mapping that zeros the gradient, we have the following.
+Answer to Problem 2.2.
+When an algorithm as in (10) that achieves exact asymptotic tracking exists, the minimal knowledge needed to design such an algorithm is given by assumptions (E2) and (O2).
+Notice that assumption (O1) is implicitly required to execute the algorithm (10).
+By comparison with Answer to Problem 2, it follows that relaxing the knowledge of $\theta(t)$ (i.e., relaxing (E1)) comes at the cost of requiring knowledge of the exosystem (i.e., requiring (E2)).
+By Theorem 4, the exosystem state $\theta$ and that of the optimization $z$ must be related, everywhere in $\Theta_{\circ},$ by
 $z(t)=\sigma(\theta(t)).$ (27)
-Intuitively, (16) is interpreted as the existence of a change of coordinates between the state of the exosystem and that of the optimization (we discuss the invertibility properties of $\sigma(\theta)$ shortly below; see Section IV-B and Remark 4).
-Remark 4.
-An important special case is obtained when $\sigma$ is the identity operator on $\Theta$; in this case, (13) simplifies to:
-$s(\theta)$ $=F_{c}(\theta,0),$
-which states that the controller vector field $F_{c}(z,y)$ must coincide with that of the exosystem $s(\theta)$ in $\Theta_{\circ}.$ In this case, (16) gives $z(t)=\theta(t);$ namely, the controller state $z(t)$ and that of the exosystem $\theta(t)$ coincide in $\Theta_{\circ}.$square
-While Theorem 4 provides a full characterization of all gradient-feedback algorithms that achieve tracking, it remains to address under what conditions on the loss $f(x,\theta)$ such an algorithm is guaranteed to exist.
-This question is addressed by the following result.
+Intuitively, (27) is interpreted as the existence of a change of coordinates between the state of the exosystem and that of the optimization; see Section 4.3 for a discussion on the invertibility properties of the function $\sigma(\theta)$.
+Remark 12 (Special case with $\sigma$ being the identity operator).
+An important special case is obtained when $\sigma$ is the identity operator on $\Theta$; in this case, (23) simplifies to $s(\theta)=F_{c}(\theta,0)$, which states that the controller vector field $F_{c}(z,y)$ must coincide with that of the exosystem $s(\theta)$ in $\Omega(\Theta_{\circ}).$ In this case, (27) gives $z(t)=\theta(t);$ namely, the controller state $z(t)$ and that of the exosystem $\theta(t)$ coincide in $\Omega(\Theta_{\circ}).$square
+While Theorem 4 provides a full characterization of all gradient-feedback algorithms that achieve tracking, it remains to address under what conditions such an algorithm is guaranteed to exist.
+This question is addressed next.
 Theorem 5 (Existence of gradient-feedback algorithms).
-Suppose Assumptions 1, 2, 3, and 4 hold.
-There exists a gradient-feedback optimization algorithm that solves Problem 1 if and only if there exists a mapping $H_{c}:\Theta\to{}^{n}$ that zeros the gradient locally.
+Suppose Assumptions 13 hold.
+Then, there exists a gradient-feedback algorithm, such that $z=z^{\star}_{\circ}$ is exponentially stable for (22), that solves Problem 1 if and only if there exists a mapping $H_{c}:\Theta\to{}^{n}$ that zeros the gradient on the limit set of (2) with respect to its initializations.
 $\square$
@@
 (If) We will prove this claim by constructing a gradient-feedback algorithm that achieves $y(t)\rightarrow 0$ as $t\to\infty$.
-First, notice that by Assumption 4, there exists a matrix $L$ such that $S-LQ$ has eigenvalues in $\mathbb{C}^{-}.$ Consider the algorithm (4) with $n_{c}=p$ and [^3] Note that the second argument of the gradient is evaluated at the algorithm state $z$ instead of the parameter vector $\theta$ as the latter is unknown.
-$F_{c}(z,y)$ $=s(z)+L(y-\nabla_{x}f(H_{c}(z),z)),$
-$G_{c}(z)$ $=H_{c}(z),$
+First, notice that by Assumption 3, there exists a matrix $L$ such that $S-LQ$ has eigenvalues in $\mathbb{C}^{-}$.
+Consider the algorithm (10) with $n_{c}=p$ and [^6] Note that the second argument of the gradient is evaluated at the algorithm state $z$ instead of the parameter vector $\theta$ as the latter is unknown.
+$F_{c}(z,y)$ $=s(z)+L(y-\nabla_{x}f(H_{c}(z),z)),$ (28a)
+$G_{c}(z)$ $=H_{c}(z),$ (28b)
 where $H_{c}(z)$ is as in (16).
-The Jacobian of $\dot{z}(t)=F_{c}(z(t),y(t))$ with respect to $z$ is given by $S-LQ$ (notice that two terms of the form $LRM$ with opposite sign cancel out in forming the Jacobian: one from the gradient term $\nabla_{x}f(H_{c}(z),z)$ and the other one from $y$).
-The claim thus follows by application of Theorem 4 with $\sigma$ the identity operator on $\Theta$.
+The Jacobian of $\dot{z}(t)=F_{c}(z(t),y(t))$ with respect to $z$ is given by $S-LQ$ (two terms of the form $LRM$ with opposite sign cancel out in forming the Jacobian: one from the gradient term $\nabla_{x}f(H_{c}(z),z)$ and the other one from $y$); since $S-LQ$ has eigenvalues in $\mathbb{C}^{-}$, $z=z^{\star}_{\circ}$ is exponentially stable for (22).
+The claim thus follows by Theorem 4 with $\sigma$ the identity operator on $\Theta$.
 ∎
-Interestingly, the conditions for existence of a gradient-feedback algorithm and those for existence of a parameter-feedback algorithm are identical.
-This is not surprising, since a parameter-feedback algorithm is assumed to have access to $\theta(t)$ while a gradient-feedback algorithm needs to measure $\theta(t)$ indirectly through $y(t).$ More precisely, as stated by (16), the dynamic state of the controller $z(t)$ acts as alternative representation (i.e., in different coordinates) of the exosystem state $\theta(t),$ while the exploration signal $x(t)$ acts as a parameter-feedback algorithm (see (15)).
-The proof of Theorem 5 is constructive, as it provides a design procedure to construct $F_{c}(z,y)$ and $G_{c}(z)$ that constitute a gradient-feedback optimization algorithm.
-Such a procedure is presented in Algorithm 1, where a Luenberger observer is used to estimate the exosystem state $\theta(t)$ (see line $4$), and a parameter feedback algorithm is then applied to the estimated exosystem state to regulate the gradient to zero (precisely, $G_{c}(z)$ is designed following the approach of Theorem 1 – see line $3$).
-Data: $s(\theta),$ $\nabla_{x}f(x,\theta),$ $H_{c}(\theta)$ satisfying (13), Jacobian matrices $Q$ and $S$ in (7)
-1 $n_{c}\leftarrow n$;
+Interestingly, the conditions for the existence of a gradient-feedback algorithm and those for the existence of a parameter-feedback algorithm (cf.
+Section 3) are identical.
+This is not surprising, since a parameter-feedback algorithm is assumed to have access to $\theta(t)$ while a gradient-feedback algorithm needs to measure $\theta(t)$ indirectly through $y(t).$ More precisely, as stated by (27), the dynamic state of the controller $z(t)$ acts as an alternative representation (i.e., in different coordinates) of the exosystem state $\theta(t),$ while the exploration signal $x(t)$ acts as a parameter-feedback algorithm (see (26)).
+4.2 Design of dynamic gradient feedback algorithms
+We now focus on addressing Problem 1. A design procedure to construct $F_{c}(z,y)$ and $G_{c}(z)$ is presented in Algorithm 1, which is derived from the proof of Theorem 5.
+The algorithm uses a Luenberger observer to estimate the exosystem state $\theta(t)$ (cf.
+line $4$), and a parameter feedback algorithm is then applied to the estimated exosystem state to regulate the gradient to zero (precisely, $G_{c}(z)$ is designed following the approach of Theorem 1 – cf.
+line $3$ of Algorithm 1).
+Notice that the algorithm uses the particular choice for $\sigma(\theta)$ being the identity operator (see Remark 12).
+Note that line $2$ of the algorithm guarantees that the equilibrium of (22) is locally exponentially stable, as required by Theorem 4, and that the existence of such $L$ is guaranteed by Assumption 3.
+Data: $s(\theta),$ $\nabla_{x}f(x,\theta),$ $G_{c}(\theta)$ satisfying (23b), Jacobian matrices $Q$ and $S$ in (21)
+1 $n_{c}\leftarrow p$;
 2 $L\leftarrow$ any matrix such that $S-LQ$ is Hurwitz;
@@
 Algorithm 1 Gradient-feedback algorithm design
-Remark 5.
-Instead of a Luenberger observer, alternative dynamic observers could be considered in Line 4 of the algorithm to achieve different asymptotic or transient properties of the resulting gradient-feedback optimization algorithm; we leave the investigation of alternative state observer algorithms as the scope of future works.
+Theorem 6 (Correctness of Algorithm 1).
+Suppose Assumptions 13 hold. Then, the control algorithm (10) designed using Algorithm 1 asymptotically tracks a critical trajectory of (1).
+Remark 13 (Alternative observer choices).
+Instead of a Luenberger observer, alternative dynamic observers could be considered in Line 4 of the algorithm to achieve different asymptotic or transient properties of the resulting gradient-feedback optimization algorithm; we leave the investigation of alternative state observer algorithms to future works.
 $\square$
-We illustrate the applicability of (13) on a quadratic problem in the following example.
+We illustrate the algorithm design procedure on a quadratic problem in the following example.
 Example 3.
-Consider the quadratic problem studied in Example 3, and assume that the exosystem follows the linear model $\dot{\theta}=S\theta$ for some matrix $S\in{}^{p\times p}.$ According to Theorem 5, an optimization algorithm given by
+Consider the quadratic problem studied in Example 2, and assume that the exosystem follows the linear model $\dot{\theta}=S\theta$ for some matrix $S\in{}^{p\times p}$ as in Assumption 2.
+Further, suppose that $Q$ is such that the unobservable subspace $\mathcal{N}=\cap_{i\geq 0}\ker(QS^{i})$ is the empty set $\mathcal{N}=\emptyset,$ so that Assumption 3 is satisfied; notice that this assumption holds, for example, when $n=p$ and $Q$ is invertible.
+According to Theorem 5, an optimization algorithm given by
 $\dot{z}=A_{c}z+B_{c}y,$ $x=G_{c}z,$ $y=Rx+Q\theta,$ (29)
@@
 $\Sigma S$ $=A_{c}\Sigma,$ (30a)
-00 0 $=(RG_{c}\Sigma+Q),$ (18b)
-for all $\theta$ in a neighborhood of the origin. When this condition holds, the application of Algorithm 1 gives
-$A_{c}=S,$ $B_{c}=L,$ $G_{c}=-R^{\dagger}Q,$
-where $L$ is any matrix such that $S-LQ$ is Hurwitz; notice that this choice satisfies (18) with $\Sigma=I.$square
-IV-B The optimization algorithm as a copy of the exosystem
-As discussed after (16), the state of any optimization algorithm that achieves asymptotic tracking and that of the exosystem must be related by a change of coordinates.
-In the following, we show that $\sigma(\theta)$ is indeed an injective map, and it can thus be interpreted as a change of coordinates.
-We begin with the following result.
-Proposition 6.
+00 0 $=RG_{c}\Sigma+Q.$ (30b)
+Next, we apply Algorithm 1 to design an algorithm that solves this problem.
+The mapping $H_{c}(\theta)=-R^{\dagger}Q\theta$ zeros the gradient (see Example 2).
+Following Algorithm 1, we set $n_{c}=p$, choose matrix $L$ such that $S-LQ$ is Hurwitz, define $G_{c}(z)=H_{c}(z)=-R^{\dagger}Qz$, and then select $A_{c}$ and $B_{c}$ such that $F_{c}(z,y)=A_{c}z+B_{c}y$ satisfies:
+$F_{c}(z,y)$ $=s(z)+L(y-\nabla_{x}f(H_{c}(z),z)),$
+$=Sz+L(y-(RH_{c}(z)+Qz)),$
+$=Sz+Ly,$
+which gives $A_{c}=S$ and $B_{c}=L.$ In summary, a dynamic gradient-feedback algorithm is (29) with $A_{c}=S$, $B_{c}=L$, and $G_{c}=-R^{\dagger}Q$, where $L$ is any matrix such that $S-LQ$ is Hurwitz.
+With this choice, conditions (30) are satisfied with $\Sigma=I$ (see Remark 12).
+Finally, the autonomous model (22) becomes $\dot{z}(t)=(S-LQ)z(t)$, which is asymptotically stable, as required by Theorem 4.
+$\square$
+4.3 The optimization algorithm as a copy of the exosystem
+As shown by (27), the state of optimization algorithms that achieve asymptotic tracking and that of the exosystem are related by a change of coordinates.
+In this section, we show that $\sigma(\theta)$ is an injective map and thus is a well-posed change of coordinates.
+For Algorithm 1, we have the following.
+Proposition 7 (Injectivity of $\sigma(\theta)$ from Algorithm 1).
 Let the assumptions of Theorem 5 hold and let $F_{c}(z,y)$ and $G_{c}(z)$ be obtained by Algorithm 1. Then, there exists an injective map $\sigma(\theta)$ such that (27) holds. $\square$
 Proof.
-The claim follows by noting that $F_{c}(z,y)$ and $G_{c}(z)$ obtained by Algorithm 1 satisfy (13) with $\sigma$ the identity operator on $\Theta$. ∎
-The interpretation of Proposition 6 is that, for any solution returned by Algorithm 1, $F_{c}(z,y)$ incorporates a copy of $s(\theta);$ precisely,
-$F_{c}(\sigma(\theta),0)=s(\theta).$ (19)
-The above feature is not a result of using Algorithm 1. Indeed, every gradient feedback algorithm has this feature, as shown next.
-Proposition 7.
+The claim follows by noting that $F_{c}(z,y)$ and $G_{c}(z)$ obtained by Algorithm 1 satisfy (23) with $\sigma$ the identity. ∎
+It follows from Proposition 7 that, for any optimization method returned by Algorithm 1, there exists an injective mapping that, locally, maps the state of the exosystem into the state of the optimization algorithm.
+This property is not exclusive to Algorithm 1; in fact, any algorithm that achieves exact asymptotic tracking possesses the same characteristic, as demonstrated next.
+Proposition 8 (Injectivity of any $\sigma(\theta)$).
 Let the assumptions of Theorem 5 hold and assume that $s(\theta)$ is $C^{\infty}.$ For any $F_{c}(z,y)$ and $G_{c}(z)$ that achieve asymptotic tracking, there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin such that $\sigma(\theta)$ is injective in $\Theta_{\circ}$$\square$
 Proof.
-By contradiction, assume that $F_{c}(z,y)$ and $G_{c}(z)$ satisfy (13), but $\sigma(\theta)$ is not injective at the origin; namely, there exists nonzero $\theta^{\prime}\in\Theta_{\circ}$ such that $\sigma(\theta^{\prime})=z_{0}^{\star}.$ From (13b), we have
-$0=\nabla_{x}f(G_{c}(\sigma(\theta^{\prime})),\theta^{\prime}).$
+By contradiction, assume that $F_{c}(z,y)$ and $G_{c}(z)$ satisfy (23), but $\sigma(\theta)$ is not injective at the origin; namely, there exists nonzero $\theta^{\prime}\in\Theta_{\circ}$ such that $\sigma(\theta^{\prime})=z_{0}^{\star}.$ Then from (23b), $0=\nabla_{x}f(G_{c}(\sigma(\theta^{\prime})),\theta^{\prime})$.
 Moreover, from Theorem 4,
 $\frac{\partial\sigma(\theta^{\prime})}{\partial\theta}s(\theta^{\prime})$ $=F_{c}(\sigma(\theta^{\prime}),0)=0.$ (31)
-For $C^{\infty}$ vector fields $h_{1}(x)$ and $h_{2}(x)$, we let
-$L_{h_{1}}(h_{2})(x)=\frac{\partial h_{2}(x)}{\partial x}h_{1}(x).$
-By application of (20), we have
+For $C^{\infty}$ vector fields $h_{1}(x)$ and $h_{2}(x)$, we let $L_{h_{1}}(h_{2})(x)=\frac{\partial h_{2}(x)}{\partial x}h_{1}(x).$ By application of (31), we have
 $0=L_{s}(\nabla_{x}f)(x,\theta^{\prime})=L_{s}(L_{s}(\nabla_{x}f))(x,\theta^{\prime})=\dots.$
-Hence, the matrix
+Hence, the matrix of Lie derivatives of the measurable signal, which, intuitively, characterizes the information content of the gradient signal:
 $\begin{bmatrix}L_{s}(\nabla_{x}f)(x,\theta)\\ L_{s}(L_{s}(\nabla_{x}f))(x,\theta)\\ \vdots\end{bmatrix},$
-is not invertible at $\theta=\theta^{\prime}.$ By [35, Thm. 3.13], the system (2) is not weakly observable, thus violating Assumption 4. ∎
-V Fidelity of the internal model and tracking accuracy
-In this section, we illustrate through examples how an imprecise knowledge of the internal model impacts the tracking accuracy.
-For tractability, we will focus on the quadratic problem (12) (see Examples 3 and 4).
-Suppose that an imprecise knowledge of the internal model is available; namely, there exists $\Delta\in{}^{n_{c}\times p}$ such that (18) is modified to:
-$\Sigma S+\Delta$ $=A_{c}\Sigma,$ (21a)
-00 0 $=(RG_{c}\Sigma+Q).$ (21b)
+is not invertible at $\theta=\theta^{\prime}.$ By [53, Thm. 3.13], the system (2) is not weakly observable, thus violating Assumption 3. ∎
+5 Fidelity of the internal model vs accuracy
+In Section 4, we have shown that a necessary condition for exact asymptotic tracking is that the algorithm incorporates an internal model of the temporal variability of the problem.
+Departing from this finding, in this section we tackle the following question: how will the algorithm perform when the temporal variability of the problem is known only approximately?
+For illustration purposes, we focus on the quadratic problem (19) (see Examples 2 and 3).
+Suppose that an imprecise internal model is available; that is, there exists $\Delta\in{}^{n_{c}\times p}$ such that (30) is modified to777 To explicitly model uncertainty in the matrix $S,$ equation (32) could be rewritten as $\Sigma(S+\Delta_{s})=A_{c}\Sigma.$ By comparison, the two approaches are equivalent with the substitution $\Delta=\Sigma\Delta_{s}.$ :
+$\Sigma S+\Delta=A_{c}\Sigma\qquad\text{and}\qquad 0=RG_{c}\Sigma+Q.$ (32)
 To analyze the asymptotic behavior of $y(t),$ it is useful to define the auxiliary variable $\tilde{z}(t)=z(t)-\Sigma\theta(t).$ Using (32) and (29), the dynamics of $\tilde{z}$ follow the model:
-$\dot{\tilde{z}}=(A_{c}+B_{c}RG_{c})\tilde{z}+\Delta\theta.$ (22a)
-The corresponding gradient signal in terms of $\tilde{z}$ is:
-$y=RG_{c}\tilde{z}.$ (22b)
-Assuming that $(A_{c}+B_{c}RG_{c})$ is Hurwitz stable (see Theorem 4), the Final Value Theorem gives:
+$\dot{\tilde{z}}=(A_{c}+B_{c}RG_{c})\tilde{z}+\Delta\theta.$ (33)
+The corresponding gradient signal in terms of $\tilde{z}$ is $y=RG_{c}\tilde{z}$. Assuming that $(A_{c}+B_{c}RG_{c})$ is Hurwitz stable (see Theorem 4), the Final Value Theorem gives:
 $y_{\infty}$ $:=\lim_{t\rightarrow\infty}y(t)$ (34)
 […]
-As a first illustrative scenario, suppose $S=0;$ namely, that the exosystem states are constants at all times.
-Notice that, since $S=0,$ from (18a), $A_{c}$ incorporates an internal model of $S$ if and only if $A_{c}\Sigma=0;$ precisely, if and only if
-$\text{dim}(\ker(A_{c}))\geq p.$
-In this case, assuming that the perturbed $A_{c}$ is such that $(A_{c}+B_{c}RG_{c})$ remains Hurwitz stable, from (23), we have:
 […]
-As expected, when $\Delta=0,$ $A_{c}$ incorporates an exact internal model, and $y_{\infty}=0$ for any $\theta(0).$ On the other hand, when $\Delta\neq 0,$
 […]
 namely, bounded errors in $\Delta$ result in a bounded $y_{\infty}.$
-Interestingly, analogous continuity properties may not hold when the Poisson stability assumption is dropped. Consider an instance of (12) with $n=p=2,$ $R=Q=I,$ and
-$S=\begin{bmatrix}0&1\\ 0&0\end{bmatrix}.$
-Suppose the algorithm (17) is used, with $G_{c}=I,B_{c}=I,$
-$A_{c}=\begin{bmatrix}-\epsilon_{1}&1\\ 0&-\epsilon_{2}\end{bmatrix},$
-where $0\leq\epsilon_{1}<1,$ $10\leq\epsilon_{2}<$ 1 .
-To avoid trivial cases, we will assume $\theta_{1}(0)$ and $\theta_{2}(0)$ are nonzero.
-When $\epsilon_{1}=\epsilon_{2}=0,$ (21) holds with $\Sigma=-I$ and $\Delta=0$, in which case $y_{\infty}=0.$ On the other hand, when $\epsilon_{1},\epsilon_{2}\neq 0,$ from (23):
-$y_{\infty}=\lim_{s\rightarrow 0}\begin{bmatrix}\frac{\epsilon_{1}}{s-\alpha_{1}}&\frac{\epsilon_{1}}{s(s-\alpha_{1})}+\frac{\epsilon_{2}}{(s-\alpha_{1})(s-\alpha_{2})}\\ 0&\frac{\epsilon_{2}}{s-\alpha_{2}},\end{bmatrix}\begin{bmatrix}\theta_{1}(0)\\ \theta_{2}(0)\end{bmatrix}$
-where $\alpha_{i}=1-\epsilon_{i}$ for $i\in\{1,2\}.$ Since $\epsilon_{1}\neq 0$ and $\theta_{2}(0)\neq 0$, we conclude that $|y(t)|\rightarrow\infty$ as $t\rightarrow\infty$, i.e., asymptotic tracking is not attained.
-In conclusion, the tracking accuracy depends on the fidelity of the internal model as well as the asymptotic behavior of the exosystem.
-VI Extensions to constrained problems
-We can extend our results to study constrained time-varying optimization problems by searching for a stationary point of the Lagrangian.
-Consider the equality-constrained problem
+Interestingly, analogous continuity properties may not hold when the trajectories of (2) are not bounded (thus violating Assumption 2).
+Consider an instance of (19) with $n=p=2,$ $R=Q=I,$ and $S=\left[\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\right]$.
+The solutions of (2) with this exosystem are a linear combination of modes $e^{0t}$ and $te^{0t},$ and thus grow unbounded at a linear rate.
+Suppose algorithm (29) is used with $G_{c}=I$, $B_{c}=I,$ and $A_{c}=\left[\begin{smallmatrix}-\epsilon_{1}&1\\ 0&-\epsilon_{2}\end{smallmatrix}\right]$, where $\epsilon_{1},\epsilon_{2}\in[0,1)$.
+When $\epsilon_{1}=\epsilon_{2}=0,$ (32) holds with $\Sigma=-I$ and $\Delta=0$, in which case $y_{\infty}=0$ and exact asymptotic tracking is achieved.
+On the other hand, when $\epsilon_{1},\epsilon_{2}\neq 0,$ (32) holds with $\Delta=\left[\begin{smallmatrix}\epsilon_{1}&0\\ 0&\epsilon_{2}\end{smallmatrix}\right]$ and, from (34), we obtain:
+$y_{\infty}=\lim_{s\rightarrow 0}\begin{bmatrix}\frac{\epsilon_{1}}{s-\alpha_{1}}&\frac{\epsilon_{1}}{s(s-\alpha_{1})}+\frac{\epsilon_{2}}{(s-\alpha_{1})(s-\alpha_{2})}\\ 0&\frac{\epsilon_{2}}{s-\alpha_{2}}\end{bmatrix}\begin{bmatrix}\theta_{1}(0)\\ \theta_{2}(0)\end{bmatrix},$
+where $\alpha_{i}=1-\epsilon_{i}$, $i\in\{1,2\}.$ Since $\epsilon_{1}\neq 0$, and provided that $\theta_{2}(0)\neq 0$, the above identity implies that $\|y(t)\|\rightarrow\infty$ as $t\rightarrow\infty$, i.e., the asymptotic tracking error grows unbounded.
+This example illustrates that the tracking accuracy depends on the fidelity of the internal model but also the asymptotic behavior of the exosystem.
+6 Extensions
+In this section, we discuss possible extensions to constrained optimization problems and to discrete-time algorithms.
+6.1 Constrained problems
+Our results can be extended to constrained time-varying optimization. Consider the equality-constrained problem
 minimize $f(x,\theta(t))$
-subject to $h_{i}(x,\theta(t))=0,\quad i=1,\ldots,m$
+subject to $h_{i}(x,\theta(t))=0,\quad i=1,\ldots,m,$
 where the constraint functions $h_{i}(x,\theta(t))$ depend on the parameter vector. The associated Lagrangian function is
-$L(x,\lambda,\theta(t))=f(x,\theta(t))+\sum_{i=1}^{m}\lambda_{i}h_{i}(x,\theta(t))$
+$L(x,\lambda,\theta(t))=f(x,\theta(t))+\sum_{i=1}^{m}\lambda_{i}h_{i}(x,\theta(t)),$
 […]
-$L(x,\bar{\lambda},\theta(t))\leq L(x,\lambda,\theta(t))\leq L(\bar{x},\lambda,\theta(t))$
-for all pairs $(\bar{x},\bar{\lambda})\in{}^{n}\times{}^{m}$.
-For any such saddle-point, $x$ is primal optimal, $\lambda$ is dual optimal, and the optimal duality gap is zero.
+$L(x,\bar{\lambda},\theta(t))\leq L(x,\lambda,\theta(t))\leq L(\bar{x},\lambda,\theta(t)).$
+For any such saddle-point, if strong duality holds, $x$ is primal optimal, $\lambda$ is dual optimal, and the optimal duality gap is zero.
 Moreover, the gradient of the Lagrangian (assuming it exists) is zero at any saddle-point.
-It follows from the derivations in the previous sections that the gradient-feedback and parameter-feedback algorithms can be directly applied to seek a stationary point of the Lagrangian function by replacing (in (1)) the variable $x$ with the extended decision variable $\tilde{x}=(x,\lambda)$ and by letting $f(\tilde{x},\theta)=L(x,\lambda,\theta).$ Notice that, if the critical point computed by (4a) is also a saddle-point, then it is also a solution to the equality-constrained problem; see [36, Ch.
+It follows from the derivations in the previous sections that the gradient-feedback and parameter-feedback algorithms can be directly applied to seek a stationary point of the Lagrangian function by replacing (in (1)) the variable $x$ with the extended decision variable $\tilde{x}=(x,\lambda)$ and by letting $f(\tilde{x},\theta)=L(x,\lambda,\theta).$ Notice that, if the critical point computed by (10b) is also a saddle-point and strong duality holds, then it is also a solution to the equality-constrained problem.
+When strong duality does not hold, however, such a saddle-point may not correspond to an optimizer [54, Ch.
 5].
-VII Simulation results
+6.2 Discrete-time algorithms
+In practice, optimization algorithms are often implemented on digital machines and thus, in these cases, implementing (10) requires discretizations.
+In these cases, a possible solution is to pursue a design by emulation [55], which consists in converting the continuous-time algorithm (10) into a discrete-time one by approximation (that is, by approximating the time-derivative by a discrete difference using, e.g., Backward/Forward Difference, Tustin, ZOH methods, etc.).
+Then, the continuous-time guarantees given here can be translated into discrete-time ones (using standard tools from emulation [55] or sampled-data systems [56]), provided that the sampling time is sufficiently small.
+In these cases, exact tracking established here will continue to hold at the discrete time instants, while tracking will be inexact in-between sampling times and the error can be bounded in terms of the sampling frequency using the explicit bounds in [55].
+An alternative method that could be pursued is that of discretizing the temporal variability $\theta(t)$, thus leading to a discrete-time optimization problem as in (4).
+For this alternative formulation, the critical trajectories are discrete signals and, in this case, it would be natural to seek a fully-discrete alternative to (10) and pursue an analysis in discrete time.
+Such a design is beyond the scope of this work and will be left as the focus of future works.
+7 Simulation results
 In this section, we illustrate our results and optimization design method through a set of numerical simulations.
+For implementation on digital machines, all algorithms have been discretized by using the explicit Runge-Kutta-Fehlberg method of order $4(5)$ with adaptive stepsize [57].
 Figure 3: Simulation results illustrating the performance of an optimization algorithm synthesized using Algorithm 1 for the quadratic instance (19) of (1).
-See Example 3 and Section VII-A for a discussion.
+See Example 2 and Section 7.1 for a discussion.
 (Top) Illustration of the temporal variability of the parameter $\theta(t)$ and of $z(t).$ (Second from top) $z(t)$ is an estimator for $\theta(t),$ and thus $z(t)\rightarrow\theta(t)$ as $t\rightarrow\infty.$ (Third from top) The proposed control algorithm is successful in regulating the gradient feedback signal $y(t)$ to zero asymptotically.
 (Bottom) Illustration that $x(t)\rightarrow x^{\star}(t)$ as $t\rightarrow\infty.$
-VII-A Optimization design for quadratic costs
+7.1 Optimization design for quadratic costs
 […]
-$F_{c}(z,y)$ $=s(z)+L(y-\nabla_{x}f(H_{c}(z),z)),$
-$=Sz+L(y-(RG_{c}z+Qz)),$
-$=Sz+Ly.$
-The simulation results are presented in Figure 3.
-It follows from our choice of $F_{c}(z,y),$ that the optimization state $z(t)$ converges to the parameter vector $\theta(t)$ (see the top two plots in Fig.
-3).
-Moreover, it follows from our choice of $G_{c}(z)$ that $y(t)\rightarrow 0$ and thus $x(t)\rightarrow x^{\star}(t)$ as $t\rightarrow\infty$ (see the bottom two plots in Fig.
-3).
+7.2 Nonlinear example
+We now illustrate our algorithm design on a nonlinear time-varying optimization problem from [30, 34], and we compare the results with several prediction-correction algorithm variants.
+Consider minimizing the function:
+$f(x,\theta)=\tfrac{1}{2}(x-\theta)^{2}+\kappa\log(1+\exp(\mu x)),$
+with parameters $\kappa=7.5$ and $\mu=1.5$, where the time-varying parameter $\theta(t)$ satisfies the exosystem dynamics $\dot{\theta}=S\theta$ where $S=\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]$.
+We compare the error $\|x(t)-x^{\star}(t)\|$ over time for several algorithms in Fig.
+4.
+The objective function satisfies the conditions in Theorem 3; for our simulations, we computed numerically a mapping that zeros the gradient locally.
+In Fig.
+4(a), we show the error for our algorithm with observer pole locations $(-\eta,-2\eta)$ for $\eta\in\{1,5,20\}$ (shown in dark to light blue) along with the corresponding rate $e^{-\eta t}$ (dashed black).
+For comparison, we also show the error of the prediction-correction (PC) algorithm (8) along with two of its finite-time (FT) variants [34, Equations (9) and (20)] using the parameter values in [34, Section VI.A].
+Instead of computing $H_{c}(\theta)$ numerically, we can use its linear approximation in (20).
+Doing so yields the results in Fig.
+4(b) for various size exosystem initial conditions $\theta(0)=\zeta\theta_{0}$ for $\zeta\in\{1,0.1,0.01\}$.
+The linear approximation does not zero the gradient exactly, which results in a finite steady-state error whose size depends on the size of the limit set of the exosystem (which corresponds to the size of the initial conditions since the exosystem dynamics are sinusoidal).
+(a) Error over time for the proposed dynamic gradient-feedback (DGF) algorithm with observer pole locations $(-\eta,-2\eta)$ for $\eta\in\{1,5,20\}$ (dark to light blue) along with the prediction-correction (PC) algorithm and two finite-time (FT) variants.
+(b) Error over time for the proposed dynamic gradient-feedback algorithm using the linear approximation to $H_{c}(\theta)$ in (20) with different size initial conditions $\theta(0)=\zeta\theta_{0}$ for $\zeta\in\{1,0.1,0.01\}$.
+When the initial conditions are small, the exosystem state stays near the equilibrium, in which case $H_{c}$ approximately zeros the gradient on the limit set of the exosystem.
+Figure 4: Dynamic gradient-feedback algorithm applied to the nonlinear time-varying optimization problem from Section 7.2.
+7.3 Application to solve the dynamic traffic assignment problem in transportation
 Figure 5: (Left) Areal view of the highway system between the cities of Wavre and Brussels, Belgium. (Right) Graph utilized to model the portion of traffic network of interest.
-VII-B Application to solve the dynamic traffic assignment problem in transportation
-We next illustrate the applicability of the framework in solving the dynamic traffic assignment problem in roadway transportation [37]; intuitively, the objective is to decide how traffic flows split among the available paths of a network to minimize the drivers’ travel time to destination.
+Figure 6: Illustration of the performance of the optimization method synthesized using Algorithm 1 to solve the dynamic traffic assignment problem (7.3).
+See Section 7.3 for a discussion.
+The bottom figure shows that, from a zero initial condition, the algorithm is capable of computing a Wardrop equilibrium in about $1.5$ hours, and then of tracking this equilibrium.
+Notice that, a faster rate of convergence could be obtained by shifting the eigenvalues of $S-LQ.$
+We next illustrate the applicability of the framework in solving the dynamic traffic assignment problem in roadway transportation [58]; intuitively, the objective is to decide how traffic flows are split among the available paths of a network to minimize the drivers’ travel time to destination.
 We model a roadway transportation network using a static flow model [58], described by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E}),$ with edges $i\in\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ (modeling traffic roads) and nodes $\mathcal{V}$ (modeling traffic junctions).
 For $i\in\mathcal{E},$ we denote by $i^{+}\in\mathcal{V}$ and $i^{-}\in\mathcal{V}$ its origin and destination nodes, respectively.
-We assume that an exogenous, time-varying, inflow of traffic $\theta(t)$ enters the network at a certain origin node, denoted by $o\in\mathcal{V},$ and exits at a certain destination node, denoted by $d\in\mathcal{V};$ for simplicity, we assume that there is only one origin-destination pair, but this is without loss of generality [37].
-We describe the network state using a vector $x\in\mathbb{R}_{\geq 0}^{|\mathcal{E}|}$ (where $|\mathcal{E}|$ denotes the number of edges) whose entries $x_{i}$ describe the amount of inflow $\theta$ routed through road $i.$ To each link $i,$ we associate a function $\ell_{i}(x_{i})$ describing the latency (or travel time) of road $i$.
+We assume that an exogenous, time-varying, inflow of traffic $\theta(t)$ enters the network at a certain origin node, denoted by $o\in\mathcal{V},$ and exits at a certain destination node, denoted by $d\in\mathcal{V};$ for simplicity, we assume that there is only one origin-destination pair, but this is without loss of generality [^8] When the network has multiple origin-destination pairs, an optimization problem of the form (7.3) will be associated with each origin-destination pair.
+Because the optimization problems are independent from one another [58], each of them can be solved independently using the methods illustrated here..
+We describe the network state using a vector $x\in\mathbb{R}_{\geq 0}^{|\mathcal{E}|}$ (where $|\mathcal{E}|$ denotes the number of edges) whose entries $x_{i}$ describe the amount of inflow $\theta$ routed through road $i.$ To each link $i,$ we associate a function $\ell_{i}(x_{i})$ describing the latency (or travel time) of the road $i$.
 For our simulations, we consider the network topology in Fig.
-4, and choose the latency functions as follows:
+5, and choose the latency functions as follows:
 $\ell_{1}(x_{1})$ $=x_{1},$ $\ell_{2}(x_{2})$ $=10x_{2},$ $\ell_{3}(x_{3})$ $=x_{3},$
 $\ell_{4}(x_{4})$ $=5x_{4},$ $\ell_{5}(x_{5})$ $=x_{5}.$
-According to Wardrop’s first principle [37, pp.
-31], transportation networks operate at a condition where travelers select their path to minimize their travel time to destination.
-Mathematically, a Wardrop’s equilibrium is the optimizer of the following optimization problem:
+According to Wardrop’s first principle [58, pp.
+31], transportation networks operate at a condition where travelers select their path to minimize their travel time to their destination.
+Mathematically, Wardrop’s equilibrium is the optimizer of the following optimization problem:
 […]
@@
 $x_{i}\geq 0,\ \forall i\in\mathcal{E},$ (35)
-where $\delta_{v}(\theta),$ $v\in\mathcal{V},$ is defined as:
-$\delta_{v}(\theta)=\begin{cases}\theta,&\text{if }v=o,\\ -\theta,&\text{if }v=d,\\ 0,&\text{otherwise}.\end{cases}$
-The loss function in (VII-B) is used to model travelers who will switch to a different path if it has shorter travel time to destination, while the first constraint in (VII-B) describes the network topology, namely, that traffic flows are conserved at each node.
-Notice that (VII-B) is a time-varying optimization problem, where the temporal variability originates from the dependence of the constraint on $\theta(t)$, which describes the inflow of vehicles at the origin and outflow at the destination, measured in vehicles per hour.
+where $\delta_{v}(\theta)$ for $v\in\mathcal{V}$ is defined as $\delta_{o}(\theta)=\theta$ for the origin node, $\delta_{d}(\theta)=-\theta$ for the destination node, and $\delta_{v}(\theta)=0$ for all other nodes.
+The loss function in (7.3) is used to model travelers who will switch to a different path if it has a shorter travel time to destination, while the first constraint in (7.3) describes the network topology, namely, that traffic flows are conserved at each node.
+Notice that (7.3) is a time-varying optimization problem, where the temporal variability originates from the dependence of the constraint on $\theta(t)$, which describes the inflow of vehicles at the origin and outflow at the destination, measured in vehicles per hour.
 For our simulations, we assume that the network inflow is sinusoidal:
 $\theta(t)=\theta_{0}-\theta_{1}\cos(\omega_{1}t+\phi_{1})-\theta_{2}\cos(\omega_{2}t+\phi_{2}),$ (36)
-where $\theta_{0},\theta_{1},\theta_{2},\omega_{1},\omega_{2},\phi_{1},\phi_{2}\in\mathbb{R}_{>0},$ satisfy $\theta_{0}>\theta_{1},$ $\theta_{0}>\theta_{2},$ and $\omega_{2}>\omega_{1}.$ The model (25) states that the network inflow is the sum of a constant term, $\theta_{0},$ a slowly-varying sinusoid with angular frequency $\omega_{1}$ and a quickly-varying sinusoid with angular frequency $\omega_{2}.$ The low-frequency sinusoid is used here to describe slowly-varying (e.g., hourly) traffic demands, while the high-frequency sinusoid is used to model sudden (e.g., at the minute-level) variations in traffic demand.
+where $\theta_{0},\theta_{1},\theta_{2},\omega_{1},\omega_{2},\phi_{1},\phi_{2}\in\mathbb{R}_{>0},$ satisfy $\theta_{0}>\theta_{1},$ $\theta_{0}>\theta_{2},$ and $\omega_{2}>\omega_{1}.$ The model (36) states that the network inflow is the sum of a constant term, $\theta_{0},$ a slowly-varying sinusoid with angular frequency $\omega_{1}$ and a quickly-varying sinusoid with angular frequency $\omega_{2}.$ The low-frequency sinusoid is used here to describe slowly-varying (e.g., hourly) traffic demands, while the high-frequency sinusoid is used to model sudden (e.g., at the minute level) variations in traffic demand.
 For our simulations, we let $\theta_{0}=3$ veh/h, $\theta_{1}=1$ veh/h, $\theta_{2}=0.1$ veh/h, $\omega_{1}=0.1$ rad/hour, $\omega_{2}=\sqrt{50}$ rad/hour, and $\phi_{1}=\phi_{2}=0,$ see Fig.
 6 (top).
-We applied Algorithm 1 to derive an optimization algorithm to solve the traffic assignment problem (VII-B). For the synthesis, we utilized the internal model $s(z)=Sz,$ where
-$S$ $=\begin{bmatrix}0&1&0&0&0\\ -\omega_{1}^{2}&0&0&0&0\\ 0&0&0&1&0\\ 0&0&-\omega_{2}^{2}&0&0\\ 0&0&0&0&0\end{bmatrix}.$
+We applied Algorithm 1 to derive an optimization algorithm to solve the traffic assignment problem (7.3). For the synthesis, we utilized the internal model $s(z)=Sz,$ where
+$S=\text{diag}\biggl{(}\begin{bmatrix}0&1\\ -\omega_{1}^{2}&0\end{bmatrix},\begin{bmatrix}0&1\\ -\omega_{2}^{2}&0\end{bmatrix},0\biggr{)}.$
 Notice that knowledge of $\theta_{0},\theta_{1},\theta_{2},\phi_{1},\phi_{2}$ is not required to synthesize the optimization algorithm – only the frequencies $\omega_{1},\omega_{2}$ are required to be known.
-To seek a solution to the constrained problem (VII-B), consider the Lagrangian function:
+To seek a solution to the constrained problem (7.3), consider the Lagrangian function:
 […]
-$\quad+\sum_{v\in\mathcal{V}}\lambda_{v}\left(\delta_{v}(\theta(t))-\sum_{\begin{subarray}{c}j\in\mathcal{E}:\\ j^{-}=v\end{subarray}}x_{j}+\sum_{\begin{subarray}{c}j\in\mathcal{E}:\\ j^{+}=v\end{subarray}}x_{j}\right),$
 where $\lambda:=(\lambda_{1},\dots,\lambda_{|\mathcal{V}|})$ is the vector of Lagrange multipliers.
-We applied Algorithm 1 to the optimization problem (1) with $f(\tilde{x},\theta)=L(x,\lambda,\theta)$ (see Section VI); the inequality constraints in (VII-B) have been accounted for by projecting $x(t)$ onto the feasible set.
+We applied Algorithm 1 to the optimization problem (1) with $f(\tilde{x},\theta)=L(x,\lambda,\theta)$ (see Section 6); the inequality constraints in (7.3) have been accounted for by projecting $x(t)$ onto the feasible set.
 Here, matrix $L$ has been chosen so that the eigenvalues of $S-LQ$ are uniformly distributed in the open real interval $(-1,-2)$.
 Notice that, since the latencies are strictly increasing, the Lagrangian is strongly convex-strongly concave [58], and thus the problem admits a unique critical point that is a saddle point.
-It follows that our algorithm is guaranteed to converge to a minimizer of (VII-B).
-Simulation for this problem are presented in Fig.
-5.
+It follows that our algorithm is guaranteed to converge to a minimizer of (7.3).
+Simulations for this problem are presented in Fig.
+6.
 The simulation shows that $z(t)\rightarrow\theta(t)$ as $t\rightarrow\infty$ (see Fig.
-5-second figure) and $y(t)\rightarrow 0,$ which implies that the algorithm successfully computes the critical point (see Fig.
-5-third figure).
+6-second figure) and $y(t)\rightarrow 0$ (up to machine precision) which implies that the algorithm successfully computes a critical point (see Fig.
+6-third figure).
 Finally, the bottom figure of Fig.
 6 illustrates the rate of convergence of the algorithm, which, as expected, is governed by the placement of the eigenvalues of the observer.
-Figure 5: Simulation results illustrating the performance of the optimization algorithm synthesized using Algorithm 1 to solve the dynamic traffic assignment problem (VII-B).
-See Section VII-B for a discussion.
-The bottom figure show that, from a zero initial condition, the algorithm is capable of computing a Wardrop equilibrium in about $1.5$ hours, and then of tracking this equilibrium.
-Notice that, a faster rate of convergence could be obtained by shifting the eigenvalues of $S-LQ.$
-VIII Conclusions
-We showed that the problem of designing optimization algorithms for time-varying optimization problems can be reconducted to the output regulation problem of nonlinear multivariable systems.
-This connection allowed us to prove the internal model principle of time-varying optimization, which states that asymptotic tracking can be achieved only if the optimization algorithm incorporates a reduplicated model of the temporal variability of the problem.
-Further, we showed that asymptotic tracking can be achieved under more relaxed assumptions (namely, assumption (A2)) than what is normally imposed in the literature (see assumption (A3)).
+8 Conclusions
+We showed that the problem of designing algorithms for time-varying optimization problems can be reformulated as a nonlinear multivariate output regulation problem.
+This connection allowed us to prove the internal model principle of time-varying optimization, which states that (when the algorithm has access only to functional evaluations of the gradient) asymptotic tracking can be achieved only if the algorithm incorporates a reduplicated model of the temporal variability of the problem.
+On the other hand, when the time-varying parameters are measurable, asymptotic tracking can be achieved by seeking a mapping that zero the gradient.
+Importantly, our results show that asymptotic tracking can be achieved under more relaxed assumptions than what is normally imposed in the literature.
 Moreover, our algorithm structure is novel in the literature, and it relies on the use of an observer for the temporal variability of the problem.
-This work opens the opportunity for several directions of future work, including an investigation at discrete-time and possibly the application in feedback optimization.
+This work opens the opportunity for several directions of future work, including the relaxation of the convexity and smoothness assumptions, an investigation of discrete-time problems, and applications to feedback optimization.
 We now summarize relevant facts in center manifold theory from [19]; see also [59]. Consider the nonlinear system:
@@
 Let $E^{-},E^{\circ},$ and $E^{+}$ be the (generalized) real eigenspaces of $F$ associated with eigenvalues of $F$ lying on the open left half plane, the imaginary axis, and the open right half plane, respectively.
-Note that $E^{\circ},E^{-},E^{+}$ have dimension $n^{\circ},n^{-},n^{+}$, respectively and that each of these spaces is invariant under the flow of $\dot{x}=Fx.$ If the linear mapping $F$ is viewed as a representation of the differential (at $x=0$) of the nonlinear mapping $f$, its domain is the tangent space $T_{0}U$ to $U$ at $x=0$, and the three subspaces in question can be viewed as subspaces of $T_{0}U$ satisfying $T_{0}U=E^{\circ}\oplus E^{-}\oplus E^{+}.$ We refer to [39, Sec.
+Note that $E^{\circ},E^{-},E^{+}$ have dimension $n^{\circ},n^{-},n^{+}$, respectively and that each of these spaces is invariant under the flow of $\dot{x}=Fx.$ If the linear mapping $F$ is viewed as a representation of the differential (at $x=0$) of the nonlinear mapping $f$, its domain is the tangent space $T_{0}U$ to $U$ at $x=0$, and the three subspaces in question can be viewed as subspaces of $T_{0}U$ satisfying $T_{0}U=E^{\circ}\oplus E^{-}\oplus E^{+}.$ We refer to [60, Sec.
 A.II] for a precise definition of $C^{k}$ manifolds; loosely speaking, a set $S\subset U$ is a $C^{k}$ manifold it can be locally represented as the graph of a $C^{k}$ function.
 Definition 4 (Locally invariant manifold).
-A $C^{k}$ manifold $S$ of $U$ is said to be locally invariant for (26) if, for each $x_{\circ}\in S$, there exists $t_{1}<0<t_{2}$ such that the integral curve $x(t)$ of (26) satisfying $x(0)=x_{\circ}$ is such that $x(t)\in S$ for all $t\in\left(t_{1},t_{2}\right)$.
+A $C^{k}$ manifold $S$ of $U$ is locally invariant for (37) if, for each $x_{\circ}\in S$, there exists $t_{1}<0<t_{2}$ such that the integral curve $x(t)$ of (37) satisfying $x(0)=x_{\circ}$ satisfiest $x(t)\in S$ for all $t\in(t_{1},t_{2})$.
 $\square$
@@
 $\dot{y}=f_{y}(\theta,y),$ $\dot{\theta}=f_{\theta}(\theta,y),$ (38)
-a curve $y=\pi(\theta)$ is an invariant manifold for (27) if the solution of (27) with $\theta(0)=\theta_{\circ}$ and $y(0)=\pi(\theta_{\circ})$ lies on the curve $y=\pi(\theta)$ for $t$ in a neighborhood of $0$ The notion of invariant manifold is useful as, under certain assumptions, it allow us to reduce the analysis of (26) to the study of a reduced system in the variable $\theta$ only.
+a curve $y=\pi(\theta)$ is an invariant manifold for (38) if the solution of (38) with $\theta(0)=\theta_{\circ}$ and $y(0)=\pi(\theta_{\circ})$ lies on the curve $y=\pi(\theta)$ for $t$ in a neighborhood of $0$ The notion of invariant manifold is useful as, under certain assumptions, it allows us to reduce the analysis of (37) to the study of a reduced system in the variable $\theta$ only.
 The remainder of this section is devoted to formalizing this fact.
@@
 $\square$
-Returning to the decomposition (27), intuitively, the invariant manifold $y=\pi(\theta)$ is said to be a center manifold when all orbits of $y$ decay to zero and those of $\theta$ neither decay nor grow exponentially.
-In what follows, we will assume that all eigenvalues of $F$ have nonpositive real part, i.e.,
-$n^{+}=0.$ (28)
-When (28) holds, it is always possible choose coordinates in $U$ such that (26) reads as:
-$\dot{y}=Ay+g(y,\theta)$ (29a)
-$\dot{\theta}=B\theta+h(y,\theta)$ (29b)
+Intuitively, the invariant manifold $y=\pi(\theta)$ is a center manifold for (38) when all orbits of $y$ decay to zero and those of $\theta$ neither decay nor grow exponentially.
+In what follows, we will assume that all eigenvalues of $F$ have nonpositive real part, i.e., $n^{+}=0$. When this holds, it is always possible to choose coordinates in $U$ such that (37) is
+$\dot{y}=Ay+g(y,\theta),$ (39a)
+$\dot{\theta}=B\theta+h(y,\theta),$ (39b)
 where $A$ is an $n^{-}\times n^{-}$ matrix having all eigenvalues with negative real part, $B$ is an $n^{\circ}\times n^{\circ}$ matrix having all eigenvalues with zero real part, and the functions $g$ and $h$ are $C^{k}$ functions vanishing at $(y,\theta)=(0,0),$ together with all their first-order derivatives.
@@
 Theorem 9 (Center manifold existence theorem).
-Assume that (28) holds. There exist a neighborhood $V\subset{}^{n^{0}}$ of 00 and a class $C^{k-1}$ mapping $\pi:V\rightarrow{}^{n^{-}}$ such that the set
+Assume that $n^{+}=0$. There exists a neighborhood $V\subset{}^{n^{0}}$ of 00 and a class $C^{k-1}$ mapping $\pi:V\rightarrow{}^{n^{-}}$ such that the set
 $S=\{(y,\theta)\in{}^{n^{-}}\times V:y=\pi(\theta)\},$
@@
 as deduced by differentiating with respect to time any solution $(y(t),\theta(t))$ of (39) on the manifold $y(t)=\pi(\theta(t))$.
-In other words, any center manifold for (29) can equivalently be described as the graph of a mapping $y=\pi(\theta)$ satisfying the partial differential equation (31), with the constraints specified by (30).
+In other words, any center manifold for (39) can equivalently be described as the graph of a mapping $y=\pi(\theta)$ satisfying the partial differential equation (41), with the constraints (40).
 Remark 14.
-Theorem 8 shows existence but not uniqueness of a center manifold.
+Theorem 9 shows existence but not the uniqueness of a center manifold.
 Moreover, (i) if $g$ and $h$ are $C^{k},k\in\mathbb{N}_{>0},$ then (39) admits a $C^{k-1}$ center manifold; (ii) if $g$ and $h$ are $C^{\infty}$ functions, then (39) has a $C^{k}$ center manifold for any finite $k$, but not necessarily a $C^{\infty}$ center manifold.
 $\square$
-The next result shows that any $y$-trajectory of (29), starting sufficiently close to the origin, converges, as time tends to infinity, to a trajectory that belongs to the center manifold.
+The next result shows that any $y$-trajectory of (39), starting sufficiently close to the origin converges, as time tends to infinity, to a trajectory that belongs to the center manifold.
 Theorem 10.
-Assume that (28) holds and suppose $y=\pi(\theta)$ is a center manifold for (29) at $(0,0)$.
+Assume that $n^{+}=0$ and suppose $y=\pi(\theta)$ is a center manifold for (39) at $(0,0)$.
 Let $(y(t),\theta(t))$ be a solution of (39).
@@
 $\|y(t)-\pi(\theta(t))\|\leq Me^{-Kt}\|y(0)-\pi(\theta(0))\|.$square
-From the above discussion, any trajectory of (29) starting at a point $y^{\circ}=\pi\left(\theta^{\circ}\right)$ of a center manifold satisfies:
-$y(t)=\pi(\zeta(t)),$ $\theta(t)=\zeta(t),$
-where $\zeta(t)$ is any solution of
-$\dot{\zeta}=B\zeta+h(\pi(\zeta),\zeta),$ (32)
-satisfying the initial condition $\zeta(0)=\theta_{\circ}$.
-This decomposition allows us to predict the asymptotic behavior of (29) by studying the asymptotic behavior of a reduced-order system, namely, (32).
-This is formalized in the following result.
+From the above discussion, any trajectory of (39) starting at a point $y^{\circ}=\pi\left(\theta^{\circ}\right)$ of a center manifold satisfies $y(t)=\pi(\zeta(t))$ and $\theta(t)=\zeta(t)$, where $\zeta(t)$ is any solution of
+$\dot{\zeta}=B\zeta+h(\pi(\zeta),\zeta),\qquad\zeta(0)=\theta_{\circ}.$ (42)
+This decomposition allows us to predict the asymptotic behavior of (39) by studying the asymptotic behavior of the reduced-order system (42). This is formalized in the following result.
 Theorem 11.

### v3

@@
 We refer to this concept as the internal model principle of time-varying optimization.
-Our analysis relies on showing that time-varying optimization problems can be recast as output regulation problems and, by using tools from center manifold theory, we establish necessary and sufficient conditions for exact asymptotic tracking.
-As a result, these findings enable the design of new algorithms for time-varying optimization.
+Our analysis relies on showing that designing optimization algorithms for time-varying problems can be reformulated as an output regulation problem.
+By leveraging tools from center manifold theory, we derive necessary and sufficient conditions for exact asymptotic tracking.
+These results, in turn, pave the way for the development of new optimization algorithms.
 We demonstrate the effectiveness of the approach through numerical experiments on both synthetic problems and the dynamic traffic assignment problem from traffic control.
@@
 Historically, discrete-time algorithms for time-varying optimization have been proposed and studied first since they emerged as a natural extension of their time-invariant counterparts, allowing for cost functions and constraints that may change over time [6].
-These approaches build on the classical perspective on mathematical optimization, which seeks to construct methods to determine optimizers and consist of iterative procedures implemented on digital devices.
+These approaches build on the classical perspective on optimization, which seeks to construct methods to determine optimizers and consist of iterative procedures implemented on digital devices.
 In recent years, however, numerous optimization algorithms have been formulated and studied as continuous-time processes [7, 8, 9], mainly due to the wide availability of tools for Ordinary Differential Equations (ODEs) that can facilitate their analysis [9].
+Figure 1: Architecture of the gradient-feedback design scheme studied in this work.
+An optimization algorithm is to be designed (bottom block), having access to gradient evaluations of the loss to be minimized (top right block), and generating a sequence of exploration points $x(t)$ at which the gradient shall be evaluated.
+The loss function varies with time, where the temporal variability $\theta(t)$ is assumed to be unmeasurable and generated by an exosystem (top left block).
+Shaded blocks emphasize the presence of dynamics.
 Motivated by these recent developments, in this paper, we study time-varying convex optimization problems and focus on the use of continuous-time dynamics to track exactly optimal solutions.
 Several approaches have already been developed for this purpose [10, 11, 12, 13], yet all these techniques implicitly require full knowledge of the temporal variability of the problem [13].
-Unfortunately, in most practical applications, having such knowledge is impractical, either because the temporal variability enters the optimization in the form of exogenous disturbances that are unknown and cannot be measured (see, e.g., [14, 15]), or simply because it is unrealistic to ask for a noiseless model of how the problem changes with time.
+Unfortunately, in most practical applications, having such knowledge is impractical, either because the temporal variability enters the optimization in the form of exogenous disturbances that are unknown and cannot be measured (see, e.g., [14, 15]), or simply because it is unrealistic to ask for a noiseless model of the problem’s variability.
 Departing from this, in this paper we pose the following question: is it possible to track (exactly and asymptotically) a minimizer of a time-varying optimization problem without any knowledge of the temporal variability of the optimization?
@@
 1 illustrates the architecture of the gradient-feedback design scheme studied in this paper.
-Figure 1: Architecture of the gradient-feedback design scheme studied in this work.
-An optimization algorithm is to be designed (bottom block), having access only to gradient evaluations of the loss function to be minimized (top right block), and generating a sequence of exploration points $x(t)$ at which the gradient shall be evaluated.
-The loss function to be optimized varies with time, where the temporal variability $\theta(t)$ is assumed to be unmeasurable and generated by an exosystem (top left block).
-Shaded blocks emphasize the presence of dynamics.
 Related works.
@@
 The first class consists of methods that do not utilize any model of the temporal variability of the problem [20, 21, 22, 23]; instead, they seek to solve a sequence of static problems.
-Several established approaches belong to this class, including the online gradient descent method [24] and the online Newton step algorithm [25] – see [26] and references therein.
-Because the temporal variability of the problem is unknown (or ignored), algorithms in this class can make a decision only after each variation has been observed, thus incurring a certain ‘regret.’ Mathematically, these techniques are capable of reaching only a neighborhood of an optimizer [13], and exact tracking is out of reach in general for these approaches.
+Several established approaches belong to this class, including the online gradient descent method [24] and the online Newton step algorithm [25]—see [26] and references therein.
+Because the temporal variability of the problem is unknown (or ignored), algorithms in this class can make a decision only after each variation has been observed, thus incurring a certain regret.
+Mathematically, these techniques are capable of reaching only a neighborhood of an optimizer [13], and exact tracking is out of reach in general for these approaches.
+This may not be surprising, as it has been recognized recently that these algorithms can be interpreted as basic integral controllers [27].
 Moreover, suitable assumptions need to be made to provide convergence guarantees for these methods; we refer to [13, 26] and references therein for a detailed discussion.
 In contrast, the second class of methods uses a model of the temporal evolution of the problem to exactly track the optimal trajectories [28, 2, 29, 30].
-Particularly celebrated is the prediction-correction algorithm (see [10], the recent work [11], and the discrete implementations [30]), whereby at each time a prediction step is used to anticipate how the optimizer will evolve over time, and a correction step is used to seek a solution to each instantaneous optimization problem.
+Particularly celebrated is the prediction-correction algorithm (see [10], the recent work [11], and the discrete implementations [31]), whereby, at each time, a prediction step is used to anticipate how the optimizer will evolve over time, and a correction step is used to seek a solution to each instantaneous optimization problem.
 We refer to [13] for a recent overview of the topic.
@@
 Initial works can be traced back to the 1970s [16, 39, 18] focusing on linear systems, and later extended to nonlinear systems using both local [40, 41] and global [42] approaches for the analysis.
-In recent years, the field has received new attention using modern methods, see, e.g., [42, 43, 44] and the recent tutorial [45].
-To the best of our knowledge, this is the first work in the literature that establishes a connection between output regulation and optimization algorithms design.
+In recent years, the field has received new attention using modern methods; see, e.g., [43, 44, 45] and the recent tutorial [46].
+To the best of our knowledge, this is the first work in the literature that establishes a connection between the problem of time-varying optimization and that of output regulation.
 Contributions.
@@
 Organization.
-Section 2 presents the problem of interest, Section 3 the parameter-feedback problem (where the temporal variability of the problem can be explicitly measured), and Section 4 the gradient-feedback problem (where the algorithm has access only to first-order functional evaluations).
+Section 2 presents the problem of interest, Section 3 the parameter-feedback problem (where the temporal variability of the problem can be explicitly measured), and Section 4 the gradient-feedback problem (where the algorithm has access only to first-order function evaluations).
 Section 5 discusses the tracking accuracy in relationship to the fidelity of the internal model, Section 6 presents extensions to constrained optimization and discrete-time algorithms, Section 7 validates numerically the results, and Section 8 illustrates our conclusions.
 Finally, in the Appendix, we summarize basic facts on center manifold theory used in the paper.
-[…]
+Notation.
+We denote by $\mathbb{S}^{n}$ the space of $n\times n$ symmetric real matrices.
+Given an open set $U,$ we say that $f:U\rightarrow\real$ is of differentiability class $C^{k}$ if it has a $k\textsuperscript{th}$ derivative that is continuous in $U.$ The gradient of $f(x,\theta):\real^{n}\times\Theta\rightarrow\real,\Theta\subseteq\real^{p},$ with respect to ${x}\in\real^{n}$ is denoted by $\nabla_{{x}}f({x},\theta):\real^{n}\times\Theta\rightarrow\real^{n}$.
+The partial derivatives of $\nabla_{{x}}f({x},\theta)$ with respect to x{x} and $\theta$ are denoted by $\nabla_{{xx}}f({x},\theta):\real^{n}\times\Theta\rightarrow\mathbb{S}^{n}$ and $\nabla_{{x}\theta}f({x},\theta):\real^{n}\times\Theta\rightarrow\real^{n\times p}$, respectively.
 2 Problem setting
@@
 We consider the time-varying optimization problem:
-$\min_{x\in{}^{n}}f(x,\theta(t)),$ (1)
-where $t\in\mathbb{R}_{\geq 0}$ denotes time and $f:{}^{n}\times\Theta\rightarrow\real,$ $\Theta\subseteq{}^{p},$ is a loss function that is parametrized by the time-varying parameter vector $\theta:\mathbb{R}_{\geq 0}\rightarrow\Theta.$ We make the following standard assumptions on the loss.
+$\min_{x\in\real^{n}}f(x,\theta(t)),$ (1)
+where $t\in\mathbb{R}_{\geq 0}$ denotes time and $f:\real^{n}\times\Theta\rightarrow\real,$ $\Theta\subseteq\real^{p},$ is a loss function that is parametrized by the time-varying parameter vector $\theta:\mathbb{R}_{\geq 0}\rightarrow\Theta.$ We make the following standard assumptions on the loss.
 Assumption 1 (Properties of the objective function).
-The map $x\mapsto f(x,\theta)$ is convex and $x\mapsto\nabla_{x}f(x,\theta)$ is Lipschitz continuous in n, for each $\theta\in\Theta.$square
+The map $x\mapsto f(x,\theta)$ is convex and $x\mapsto\nabla_{x}f(x,\theta)$ is Lipschitz continuous in n, for each $\theta\in\Theta.$ $\square$
 Convexity and smoothness are standard assumptions in optimization [26], which have been widely used in works on related problems [8, 29, 14, 15, 10, 11, 12, 13, 24, 33, 34, 30, 32].
@@
 Assumption 2 (Class of temporal variabilities).
-There exists a smooth (i.e., $C^{\infty}$) vector field $s:\Theta\rightarrow{}^{p}$ and $\theta(0)\in\Theta$ such that the parameter vector $\theta(t)$ satisfies
+There exists a smooth (i.e., $C^{\infty}$) vector field $s:\Theta\rightarrow\real^{p}$ and $\theta(0)\in\Theta$ such that the parameter vector $\theta(t)$ satisfies
 $\dot{\theta}(t)=s(\theta(t)),$ (2)
@@
 In line with [40, 39], we call the autonomous system (2) the exosystem.
-For simplicity of the presentation, we assume that $\Theta$ is some neighborhood of the origin of p.
-We put no restrictions on the size of this neighborhood (which is, e.g., allowed to be the entire space $\Theta={}^{p}$), and thus on the size of $\theta(t)$ nor on its temporal variation.
-Moreover, there is no restriction with asking that $\Theta$ contains the origin because, if $\theta(t)$ takes values in the neighborhood of any other point, such a point can be shifted to the origin through a time-invariant change of variables without altering the solutions of (1).
+For simplicity of the presentation, we assume that $\Theta$ is some neighborhood of the origin of p, corresponding to the requirement that $\theta=0$ is an equilibrium of (2).
+We put no restrictions on the size of this neighborhood (which is, e.g., allowed to be the entire space $\Theta=\real^{p}$), and thus on the size of $\theta(t)$ nor on its temporal variation.
+Moreover, there is no restriction with asking that $\theta=0$ is an equilibrium because, if the equilibrium is located at any other point, the latter can be shifted to the origin through a time-invariant change of variables that does not alter the critical points of (1) or the validity of Assumption 1.
 Remark 1 (Parametrization in time-varying optimization).
 In (1), the time variation is captured implicitly by the parameter $\theta(t)$. A related, yet slightly more general, problem is:
-$\min_{x\in{}^{n}}f_{0}(x,t),$ (3)
+$\min_{x\in\real^{n}}f_{0}(x,t),$ (3)
 where the dependency on time is explicit.
@@
 $\square$
-In what follows, we say that $x^{\star}(t),$ with $x^{\star}:\mathbb{R}_{\geq 0}\rightarrow{}^{n}$, is a critical trajectory of (1) if it satisfies:
+In what follows, we say that $x^{\star}(t),$ with $x^{\star}:\mathbb{R}_{\geq 0}\rightarrow\real^{n}$, is a critical trajectory of (1) if it satisfies:
 $0=\nabla_{x}f(x^{\star}(t),\theta(t)),\quad\forall t\in\mathbb{R}_{\geq 0}.$ (5)
@@
 Continuity of the critical trajectories can also be ensured under standard assumptions: for example, by Berge’s theorem [47], by requiring that $f(x,\theta)$ is continuous in $\theta$.
-In our analysis, we will use a critical point with $\theta$ at rest, which is a constant vector $x^{\star}_{\circ}\in{}^{n}$, defined implicitly as:
+In our analysis, we will use a critical point with $\theta$ at rest, which is a constant vector $x^{\star}_{\circ}\in\real^{n}$, defined implicitly as:
 $0=\nabla_{x}f(x^{\star}_{\circ},0).$ (6)
@@
 where $\eta>0$ models the algorithm’s step size.
-It is known [48] that this class of algorithms is capable of computing (asymptotically) critical points of (1) only when $\theta(t)$ is a constant signal; in all other cases, the algorithm converges to a neighborhood of a critical point and convergence is inexact.
+It is known [49] that this class of algorithms is capable of computing (asymptotically) critical points of (1) only when the latter are constant; in all other cases, the algorithm converges to a neighborhood of a critical point and convergence is inexact.
 We refer the reader to Remark 11 for an insightful interpretation of this fundamental limitation.
@@
 Variations of this algorithm have also been proposed to achieve fixed-time or finite-time convergence [35].
-We conclude by noting that implementing (8) requires (i) the cost to be twice continuously differentiable and strongly convex, (ii) knowledge of the maps $s(\theta)$, $\nabla_{x}f(x,\theta)$, $\nabla_{xx}f(x,\theta)$, and $\nabla_{x\theta}f(x,\theta)$, and (iii) knowledge of $\theta(t).$square
+We conclude by noting that implementing (8) requires (i) the cost to be twice continuously differentiable and strongly convex, (ii) knowledge of the maps $s(\theta)$, $\nabla_{x}f(x,\theta)$, $\nabla_{xx}f(x,\theta)$, and $\nabla_{x\theta}f(x,\theta)$, and (iii) knowledge of $\theta(t).$ $\square$
 2.2 Problem statement
-In line with the literature [26, 10, 11, 34, 12], we focus on gradient-type algorithms for solving (1), which are algorithms that have access to first-order oracles of the cost; that is, functional evaluations of the map:
+In line with the literature [26, 10, 11, 35, 12], we focus on gradient-type algorithms for solving (1), which are algorithms that have access to first-order oracles of the cost; that is, function evaluations of the map:
 $(t,x)\mapsto\nabla_{x}f(x,\theta(t)).$ (9)
 Remark 6 (Evaluating the gradient).
-In the optimization literature [26], functional evaluations of (9) are generally obtained through two main approaches: (i) the algorithm has access to both the analytical form of $\nabla_{x}f(x,\theta)$ and $\theta(t)$ at each time $t$ (either because this signal is known or measurable), or (ii) an oracle returns evaluations of (9) (computed, e.g., through numerical differentiation).
+In the optimization literature [26], function evaluations of (9) are generally obtained through two main approaches: (i) the algorithm has access to both the analytical form of $\nabla_{x}f(x,\theta)$ and $\theta(t)$ at each time $t$ (either because this signal is known or measurable), or (ii) an oracle returns evaluations of (9) (computed, e.g., through numerical differentiation).
 These alternatives will be discussed more in detail in (O1)–(O4), shortly below.
 $\square$
-We let the optimization algorithm be described by a dynamic internal state $z(t),$ taking values on an open subset $\mathcal{Z}\subseteq{}^{n_{c}},n_{c}\in\mathbb{N}_{>0}$.
-The optimization algorithm generates a sequence of points $x(t)\in{}^{n}$ (called exploration signal), and has access to functional evaluations of (9) at these points (called gradient feedback signal):
+We let the optimization algorithm be described by a dynamic internal state $z(t),$ taking values on an open subset $\mathcal{Z}\subseteq\real^{n_{c}},n_{c}\in\mathbb{N}_{>0}$.
+The optimization algorithm generates a sequence of points $x(t)\in\real^{n}$ (called exploration signal), and has access to function evaluations of (9) at these points (called gradient feedback signal):
 $y(t)=\nabla_{x}f(x(t),\theta(t)).$ (10a)
@@
 $\dot{z}(t)$ $=F_{c}(z(t),y(t)),$ $x(t)$ $=G_{c}(z(t)),$ (10b)
-where $F_{c}:\mathcal{Z}\times{}^{n}\rightarrow{}^{n_{c}}$ and $G_{c}:\mathcal{Z}\rightarrow{}^{n}$ are functions to be designed.
+where $F_{c}:\mathcal{Z}\times\real^{n}\rightarrow\real^{n_{c}}$ and $G_{c}:\mathcal{Z}\rightarrow\real^{n}$ are functions to be designed.
 In the remainder, we refer to (10) as a dynamic gradient-feedback optimization algorithm.
@@
 For simplicity of the presentation, we will require that $F_{c}(z,y)$ and $G_{c}(z)$ are such that [^3] Notice that this is without loss of generality since $F_{c}(z,y)$ and $G_{c}(z)$ are to be designed and $x_{\circ}^{\star}$ is known through (6).:
-00 0 $=F_{c}(z_{\circ}^{\star},0)\qquad\text{and}\qquad x_{\circ}^{\star}=G_{c}(z_{\circ}^{\star}),$ (11)
-for some locally unique $z_{\circ}^{\star}\in{}^{n_{c}}$. Together with (6), this ensures that $z_{\circ}^{\star}$ is an equilibrium of (10) when the gradient feedback signal $y(t)$ is equal to zero.
+$0$ $=F_{c}(z_{\circ}^{\star},0)\qquad\text{and}\qquad x_{\circ}^{\star}=G_{c}(z_{\circ}^{\star}),$ (11)
+for some locally unique $z_{\circ}^{\star}\in\real^{n_{c}}$.
+Together with (6), this ensures that $z_{\circ}^{\star}$ is an equilibrium of (10) when the gradient feedback signal $y(t)$ is equal to zero.
 The dynamics of algorithm (10), coupled with the exosystem (2), have the form of a nonlinear autonomous system:
@@
 Table 1: Summary of required knowledge and known guarantees by various established approaches.
-The symbol ()($\checkmark$)( ✓ ) is used when (O1) follows from (E1) and (O2).
+The symbol $(\checkmark)$ is used when (O1) follows from (E1) and (O2).
 See Remarks 4, 5, and 7 for discussions.
+Notice that particular assumptions on the optimization (1) and the exosystem (2) may differ across works.
 Approximate (E1) (E2) (O1) (O2) (O3) (O4) (O5)
@@
 $x(t)=H_{c}(\theta(t)),$ (14)
-where $H_{c}:\Theta\rightarrow{}^{n}$ is a mapping to be designed.
+where $H_{c}:\Theta\rightarrow\real^{n}$ is a mapping to be designed.
 Because of the explicit dependence on $\theta(t),$ we will refer to (14) to as a parameter-feedback optimization algorithm.
@@
 In analogy with Definition 1, we will say that the parameter-feedback algorithm (14) exactly asymptotically tracks a critical trajectory of (1) if there exists a neighborhood $\Theta_{s}\subset\Theta$ of the origin such that, for each initial condition $\theta(0)\in\Theta_{s}$, the solution of (15) satisfies $y(t)\to 0$ as $t\to\infty$.
-For the algorithm (14) considered in this section, Problems 2.2 and 1 are reformulated as follows.
+For the algorithm (14) considered in this section, Problems 2.2 and 1 are reformulated, respectively, as follows.
 Problem 2 (Minimal knowledge for parameter feedback).
@@
 Definition 2 (Mapping zeroing the gradient).
-We say that a mapping $H_{c}:\Theta\to{}^{n}$ zeros the gradient at the point $\theta\in\Theta$ if
-00 0 $=\nabla_{x}f(H_{c}(\theta),\theta).$ (16)
+We say that a mapping $H_{c}:\Theta\to\real^{n}$ zeros the gradient at the point $\theta\in\Theta$ if
+$0$ $=\nabla_{x}f(H_{c}(\theta),\theta).$ (16)
 Moreover, we say that $H_{c}$ zeros the gradient on a set $\Theta_{\circ}\subseteq\Theta$ if (16) holds for all $\theta\in\Theta_{\circ}$. $\square$
 Definition 3 (Limit point and limit set).
-A point $\theta_{\omega}\in\Theta$ is a limit point with respect to the initialization $\theta_{\circ}\in\Theta$ if there exists a sequence $\{t_{i}\}_{i\in\mathbb{N}_{\geq 0}}$ with $t_{i}\to\infty$ as ${i\to\infty}$ such that the exosystem (2) with $\theta(0)=\theta_{\circ}$ satisfies $\theta({t_{i}})\to\theta_{\omega}$ as ${i\to\infty}$.
-Let $\Omega(\theta_{\circ})$ denote the set of all limit points (i.e., for all sequences $t_{i}$) of (2) with respect to the initialization $\theta_{\circ}\in\Theta.$ Given $\Theta_{\circ}\subseteq\Theta,$ the set $\Omega(\Theta_{\circ}):=\cup_{\theta_{\circ}\in\Theta_{\circ}}\Omega(\theta_{\circ})$ is called the limit set with respect to initializations in $\Theta_{\circ}$ [49$\square$
+[…]
 Intuitively, $\Omega(\Theta_{\circ})$ denotes the set of all limit points (equilibria, limit cycles, etc.) that can be reached by the exosystem when initialized at points in $\Theta_{\circ}$.
@@
 Therefore, we have the following.
-Answer to Problem 2.
-When a parameter-feedback algorithm that achieves exact tracking exists, the minimal knowledge needed to design such an algorithm is (O2).
-Notice that assumption (E1) is implicitly required to execute the algorithm (14).
+Answer to Problem 2: When a parameter-feedback algorithm exists, the minimal knowledge needed to design such an algorithm is (O2).
+Notice that assumption (E1) is implicitly required for (14) to be implementable.
 We illustrate the applicability of the result and the necessity of the provided condition in the following example.
@@
 This graphical observation can be formalized with the aid of the implicit function theorem [48], as described next.
-Define $F(x,\theta):=\nabla_{x}f(x,\theta)$ and notice that $F$ is continuously differentiable with $F(x^{\star}_{\circ,2},0)=0.$ By the implicit function theorem, there exists a neighborhood $\Theta_{\circ}$ of $x^{\star}_{\circ,2}$ and a function $H_{c}:\Theta_{\circ}\rightarrow{}^{n}$ such that $F(H_{c}(\theta),\theta)=0$ in $\Theta_{\circ},$ provided that $\left.\frac{\partial F}{\partial x}\right|_{(x,\theta)=(x^{\star}_{\circ,2},0)}\neq 0.$ By inspection, it is immediate to see that the latter condition is satisfied.
+Define $F(x,\theta):=\nabla_{x}f(x,\theta)$ and notice that $F$ is continuously differentiable with $F(x^{\star}_{\circ,2},0)=0.$ By the implicit function theorem, there exists a neighborhood $\Theta_{\circ}$ of $x^{\star}_{\circ,2}$ and a function $H_{c}:\Theta_{\circ}\rightarrow\real^{n}$ such that $F(H_{c}(\theta),\theta)=0$ in $\Theta_{\circ},$ provided that $\left.\frac{\partial F}{\partial x}\right|_{(x,\theta)=(x^{\star}_{\circ,2},0)}\neq 0.$ By inspection, it is immediate to see that the latter condition is satisfied.
 $\square$
 We conclude by observing that the proof of Theorem 1 is constructive in that it provides an explicit procedure to construct parameter-feedback algorithms.
-Answer to Problem 3.
-When a parameter-feedback algorithm exists, a procedure to design it requiring minimal knowledge (O2) is to first determine the limit set $\Omega(\Theta_{\circ})$ for the given set of initial conditions $\Theta_{\circ}$ and then find a mapping $H_{c}(\theta)$ that zeros the gradient on the limit set, in which case the parameter-feedback algorithm is given by (14).
+Answer to Problem 3: When a parameter-feedback algorithm exists, a procedure to design it requiring minimal knowledge is given by: 1) determine the limit set $\Omega(\Theta_{\circ})$ for the given set of initial conditions $\Theta_{\circ}$; 2) find a mapping $H_{c}(\theta)$ that zeros the gradient on the limit set; and, 3) the parameter-feedback algorithm is given by (14).
 We conclude by illustrating the parameter-feedback design procedure on a quadratic problem.
@@
 $f(x(t),\theta(t))=\tfrac{1}{2}x(t)^{\mathsf{T}}Rx(t)+x(t)^{\mathsf{T}}Q\theta(t),$ (19)
-with matrices $R\in\mathbb{S}^{n}$ and $Q\in{}^{n\times p}$. This loss is convex and Lipschitz smooth and thus satisfies Assumption 1. In this case, the signal we wish to regulate to zero is:
+with matrices $R\in\mathbb{S}^{n}$ and $Q\in\real^{n\times p}$. This loss is convex and Lipschitz smooth and thus satisfies Assumption 1. In this case, the signal we wish to regulate to zero is:
 $y(t)=\nabla_{x}f(x(t),\theta(t))=Rx(t)+Q\theta(t).$
 For arbitrary $\theta,$ this problem admits a critical point if and only if $\operatorname{Im}Q\subseteq\operatorname{Im}R$, in which case $x_{\circ}^{\star}$ is unique.
-Applying Theorem 1 amounts to finding a linear transformation $H_{c}\in{}^{n\times p}$ such that $0=(RH_{c}+Q)\theta$ for all $\theta$ in a neighborhood of the origin.
+Applying Theorem 1 amounts to finding a linear transformation $H_{c}\in\real^{n\times p}$ such that $0=(RH_{c}+Q)\theta$ for all $\theta$ in a neighborhood of the origin.
 Assuming $\operatorname{Im}Q\subseteq\operatorname{Im}R$, we can choose $H_{c}=-R^{\dagger}Q$, where $R^{\dagger}$ is the pseudo-inverse of $R$.
@@
 This property holds because the mapping $H_{c}(\theta)$ derived here zeros the gradient globally (and not only in some neighborhood of the origin).
-In fact, this feature holds for the any optimization problem for which the gradient of the cost admits a mapping that zeros the gradient globally.
+More generally, this feature will hold for any problem for which the gradient of the cost admits a mapping that zeros the gradient globally.
 A set of sufficient conditions to check for this property are given in Theorem 3.
@@
     the Hessian $\left.\nabla^{2}_{xx}f(x,\theta)\right|_{x=x^{\star}_{\circ},\theta=0}$ is positive definite.
-Then, there exists a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient on some neighborhood of the origin of $\Theta$. $\square$
+Then, there exists a $C^{0}$ mapping $H_{c}:\Theta\rightarrow\real^{n}$ that zeros the gradient on some neighborhood of the origin of $\Theta$. $\square$
 Proof.
@@
 $\det\left[\frac{\partial F(x,\theta)}{\partial x}\right]_{(x,\theta)=(x_{\circ}^{\star},0)}\neq 0.$
-Then, the result follows from the implicit function theorem (see, e.g., [50, Thm.
+Then, the result follows from the implicit function theorem (see, e.g., [51, Thm.
 1]) applied to the first-order optimality conditions $0=F(x,\theta)$ at the point $(x,\theta)=(x^{\star}_{\circ},0)$.
@@
 is a first-order approximation of a mapping that zeros the gradient on a neighborhood of the origin. $\square$
-Although the conditions in Proposition 2 are immediate to verify, the convergence claims of Theorem 1 and Proposition 2 are of local nature, namely, $y(t)\rightarrow 0$ is ensured provided that $\theta(0)$ is sufficiently close to the origin.
+Although the conditions in Proposition 2 are immediate to verify, the convergence claims of Theorem 1 and Proposition 2 are of local nature; namely, $y(t)\rightarrow 0$ is ensured provided that $\theta(0)$ is sufficiently close to the origin.
 The following result provides a sufficient condition for global convergence.
@@
   1. (i)
-    the loss function $f$ is $C^{1}$ on ×${}^{n}\times\Theta$,
+    the loss function $f$ is $C^{1}$ on $\real{}^{n}\times\Theta$,
   2. (ii)
@@
   3. (iii)
-    the Hessian $\nabla^{2}_{xx}f(x,\theta)$ is positive definite on ×${}^{n}\times\Theta$,
+    the Hessian $\nabla^{2}_{xx}f(x,\theta)$ is positive definite on $\real{}^{n}\times\Theta$,
   4. (iv)
-    the mappings $\nabla_{xx}^{2}f$ and $\frac{\partial\nabla_{xx}^{2}f}{\partial x}$ are $C^{0}$ on ×${}^{n}\times\Theta.$
-Then, there exits a $C^{0}$ mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient everywhere in $\Theta.$square
+    the mappings $\nabla_{xx}^{2}f$ and $\frac{\partial\nabla_{xx}^{2}f}{\partial x}$ are $C^{0}$ on $\real{}^{n}\times\Theta.$
+Then, there exits a $C^{0}$ mapping $H_{c}:\Theta\rightarrow\real^{n}$ that zeros the gradient everywhere in $\Theta.$ $\square$
 Proof.
-Define $F(x,\theta)=\nabla_{x}f(x,\theta)$ and note that, under the stated assumptions, $F$ is $C^{0}$ on ×${}^{n}\times\Theta$, the mapping $x\mapsto F(x,\theta)$ is $C^{2}$ on n for all $\theta\in\Theta$,
-$\det\left[\frac{\partial F(x,\theta)}{\partial x}\right]\neq 0,\quad\forall(x,\theta)\in{}^{n}\times\Theta,$
-and the mappings $\frac{\partial F}{\partial x}$ and $\frac{\partial^{2}F}{\partial x^{2}}$ are $C^{0}$ in ×${}^{n}\times\Theta$.
-Hence, assumptions (B1)–(B5) of a global version of the implicit function theorem [50, Thm.
-2] (see also [51, Thm.
+Define $F(x,\theta)=\nabla_{x}f(x,\theta)$ and note that, under the stated assumptions, $F$ is $C^{0}$ on $\real{}^{n}\times\Theta$, the mapping $x\mapsto F(x,\theta)$ is $C^{2}$ on n for all $\theta\in\Theta$,
+$\det\left[\frac{\partial F(x,\theta)}{\partial x}\right]\neq 0,\quad\forall(x,\theta)\in\real^{n}\times\Theta,$
+and the mappings $\frac{\partial F}{\partial x}$ and $\frac{\partial^{2}F}{\partial x^{2}}$ are $C^{0}$ in $\real{}^{n}\times\Theta$.
+Hence, assumptions (B1)–(B5) of a global version of the implicit function theorem [51, Thm.
+2] (see also [52, Thm.
 6]) are satisfied for the first-order optimality conditions $0=F(x,\theta)$, and the claim follows.
 ∎
-Theorem 3 shows that, under additional continuity assumptions on the loss function and when the Hessian of the loss is positive definite everywhere, there exists a mapping $H_{c}:\Theta\rightarrow{}^{n}$ that zeros the gradient everywhere in $\Theta.$
+Theorem 3 shows that, under additional continuity assumptions on the loss function and when the Hessian of the loss is positive definite everywhere, there exists a mapping $H_{c}:\Theta\rightarrow\real^{n}$ that zeros the gradient everywhere in $\Theta.$
 By combination of Theorem 3 and Theorem 1, it follows that under the assumptions of Theorem 3, the parameter-feedback algorithm $x(t)=H_{c}(\theta(t))$ ensures that $y(t)\rightarrow 0$ as $t\rightarrow\infty$ for all initial conditions $\theta(0)\in\Theta$.
@@
 The pair $(Q,S)$ is detectable. $\square$
-The detectability property of Assumption 3 is equivalent to the existence of a local exponential observer for all (non-asymptotically stable) modes of $\theta(t)$ based on measurements $y(t)$ [52, Cor.
+The detectability property of Assumption 3 is equivalent to the existence of a local exponential observer for all (non-asymptotically stable) modes of $\theta(t)$ based on measurements $y(t)$ [53, Cor.
 3.4].
-Undetectability corresponds to a redundant description of the exogenous signal: if some modes of $\theta(t)$ are undetectable, these modes do not influence the gradient (and, as a consequence, the critical points of (1)), and thus they can be removed from the optimization problem (1) without altering its critical points.
+Undetectability corresponds to a redundant description of the exogenous signal: if some modes of $\theta(t)$ are undetectable, these modes do not influence the gradient (at $(x_{\circ}^{\star},0)$) and, as a consequence, the critical points of (1); therefore, they can be removed from the optimization problem (1) without altering its critical points.
 In other words, Assumption 3 ensures that (2) does not contain redundancies.
@@
 $\left.\frac{\partial\sigma(\theta)}{\partial\theta}\right|_{\theta=\theta_{\omega}}s(\theta_{\omega})$ $=F_{c}(\sigma(\theta_{\omega}),0),$ (23a)
-00 0 $=\nabla_{x}f(G_{c}(\sigma(\theta_{\omega})),\theta_{\omega}),$ (23b)
+$0$ $=\nabla_{x}f(G_{c}(\sigma(\theta_{\omega})),\theta_{\omega}),$ (23b)
 at all limit points $\theta_{\omega}\in\Omega(\Theta_{\circ})$.
@@
 We now prove that (23) implies $\lim_{t\rightarrow\infty}y(t)=0$}.
-It follows from (23) and Theorem 10 that the graph of the mapping $z=\sigma(\theta)$ (i.e., $(\sigma(\theta),\theta)$) is a center manifold for the coupled dynamics (12).
-Moreover, such manifold is locally attractive; namely, $z(t)\rightarrow\sigma(\theta(t))$ as $t\rightarrow\infty.$ Then, the fulfillment of (23b) guarantees that the right-hand side of (25) $y(t)\rightarrow 0.$ The conclusion then follows by iterating the (If) part of the proof of Theorem 1.
+It follows from (23) that the graph of the mapping $z=\sigma(\theta)$ (i.e., $(\sigma(\theta),\theta)$) is a center manifold for the coupled dynamics (12).
+Hence, by Theorem 10, such manifold is locally attractive; namely, $z(t)\rightarrow\sigma(\theta(t))$ as $t\rightarrow\infty.$ Then, the fulfillment of (23b) guarantees that the right-hand side of (25) $y(t)\rightarrow 0.$ The conclusion then follows by iterating the (If) part of the proof of Theorem 1.
 ∎
@@
 $x(t)=G_{c}(\sigma(\theta)),$ (26)
-is a parameter-feedback optimization algorithm for (1). Finally, we note that the existence of an exponentially stable system as in (22) is always guaranteed under Assumption 3 [52, Sec. 3].
+is a parameter-feedback optimization algorithm for (1). Finally, we note that the existence of an exponentially stable system as in (22) is always guaranteed under Assumption 3 [53, Sec. 3].
 The following interpretation follows from our findings.
 Remark 10 (The internal model principle).
-We refer to condition (23a) as the internal model principle of time-varying optimization, as it encapsulates the requirement that an optimization algorithm must incorporate an internal model of the exosystem to achieve exact asymptotic tracking.
+We refer to condition (23a) as the internal model principle of time-varying optimization, as it encapsulates the requirement that an optimization algorithm must incorporate an internal model of the exosystem to achieve exact tracking.
 Note that the use of a copy of the temporal variability of the optimization problem is explicit in the prediction-correction algorithm (see (7)).
@@
 Remark 11 (Internal model-based interpretation of basic gradient flow methods).
-Recall that the gradient-flow algorithm (4) can be viewed as an instance of (10a) with $F_{c}(z,y)$ and $G_{c}(z)$ as in Remark 7.
-By direct substitution into (23), it is immediate to see that this algorithm satisfies (23a) with $s(\theta)=0$ and $\sigma(\theta)$ arbitrary.
-Since $s(\theta)=0$ describes the internal model of a constant signal, it follows from Theorem 4 that these algorithms achieve exact asymptotic tracking if and only if $\theta(t)$ is a constant signal.
-This observation is in line with the inexact convergence properties of these algorithms typically given for these algorithms [48, 26, 14$\square$
-Since by, Theorem 4, the problem of exact asymptotic tracking can be reduced to that of finding an optimization algorithm that is algebraically equivalent to the exosystem and a mapping that zeros the gradient, we have the following.
-Answer to Problem 2.2.
-When an algorithm as in (10) that achieves exact asymptotic tracking exists, the minimal knowledge needed to design such an algorithm is given by assumptions (E2) and (O2).
+[…]
+Since Theorem 4 shows that the problem of exact asymptotic tracking can be reduced to finding an optimization algorithm that is algebraically equivalent to the exosystem together with a mapping that zeros the gradient, we conclude the following.
+Answer to Problem 2.2: When an algorithm as in (10) that achieves exact asymptotic tracking exists, the minimal knowledge needed to design such an algorithm is given by assumptions (E2) and (O2).
 Notice that assumption (O1) is implicitly required to execute the algorithm (10).
-By comparison with Answer to Problem 2, it follows that relaxing the knowledge of $\theta(t)$ (i.e., relaxing (E1)) comes at the cost of requiring knowledge of the exosystem (i.e., requiring (E2)).
+By comparison with the Answer to Problem 2, it follows that relaxing the knowledge of $\theta(t)$ (i.e., relaxing (E1)) comes at the cost of requiring knowledge of the exosystem (i.e., requiring (E2)).
 By Theorem 4, the exosystem state $\theta$ and that of the optimization $z$ must be related, everywhere in $\Theta_{\circ},$ by
@@
 Remark 12 (Special case with $\sigma$ being the identity operator).
-An important special case is obtained when $\sigma$ is the identity operator on $\Theta$; in this case, (23) simplifies to $s(\theta)=F_{c}(\theta,0)$, which states that the controller vector field $F_{c}(z,y)$ must coincide with that of the exosystem $s(\theta)$ in $\Omega(\Theta_{\circ}).$ In this case, (27) gives $z(t)=\theta(t);$ namely, the controller state $z(t)$ and that of the exosystem $\theta(t)$ coincide in $\Omega(\Theta_{\circ}).$square
+An important special case is obtained when $\sigma$ is the identity operator on $\Theta$; in this case, (23) simplifies to $s(\theta)=F_{c}(\theta,0)$, which states that the controller vector field $F_{c}(z,y)$ must coincide with that of the exosystem $s(\theta)$ in $\Omega(\Theta_{\circ}).$ In this case, (27) gives $z(t)=\theta(t);$ namely, the controller state $z(t)$ and that of the exosystem $\theta(t)$ coincide in $\Omega(\Theta_{\circ}).$ $\square$
 While Theorem 4 provides a full characterization of all gradient-feedback algorithms that achieve tracking, it remains to address under what conditions such an algorithm is guaranteed to exist.
@@
 Suppose Assumptions 13 hold.
-Then, there exists a gradient-feedback algorithm, such that $z=z^{\star}_{\circ}$ is exponentially stable for (22), that solves Problem 1 if and only if there exists a mapping $H_{c}:\Theta\to{}^{n}$ that zeros the gradient on the limit set of (2) with respect to its initializations.
+Then, there exists a gradient-feedback algorithm, such that $z=z^{\star}_{\circ}$ is exponentially stable for (22), that solves Problem 1 if and only if there exists a mapping $H_{c}:\Theta\to\real^{n}$ that zeros the gradient on the limit set of (2) with respect to its initializations.
 $\square$
@@
 (If) We will prove this claim by constructing a gradient-feedback algorithm that achieves $y(t)\rightarrow 0$ as $t\to\infty$.
-First, notice that by Assumption 3, there exists a matrix $L$ such that $S-LQ$ has eigenvalues in $\mathbb{C}^{-}$.
+First, notice that by Assumption 3, there exists a matrix $L$ such that SLQS-LQ has eigenvalues in $\mathbb{C}^{-}$.
 Consider the algorithm (10) with $n_{c}=p$ and [^6] Note that the second argument of the gradient is evaluated at the algorithm state $z$ instead of the parameter vector $\theta$ as the latter is unknown.
@@
 where $H_{c}(z)$ is as in (16).
-The Jacobian of $\dot{z}(t)=F_{c}(z(t),y(t))$ with respect to $z$ is given by $S-LQ$ (two terms of the form $LRM$ with opposite sign cancel out in forming the Jacobian: one from the gradient term $\nabla_{x}f(H_{c}(z),z)$ and the other one from $y$); since $S-LQ$ has eigenvalues in $\mathbb{C}^{-}$, $z=z^{\star}_{\circ}$ is exponentially stable for (22).
+The Jacobian of $\dot{z}(t)=F_{c}(z(t),y(t))$ with respect to $z$ is given by SLQS-LQ (two terms of the form $LRM$ with opposite sign cancel out in forming the Jacobian: one from the gradient term $\nabla_{x}f(H_{c}(z),z)$ and the other one from $y$); since SLQS-LQ has eigenvalues in $\mathbb{C}^{-}$, $z=z^{\star}_{\circ}$ is exponentially stable for (22).
 The claim thus follows by Theorem 4 with $\sigma$ the identity operator on $\Theta$.
 ∎
-Interestingly, the conditions for the existence of a gradient-feedback algorithm and those for the existence of a parameter-feedback algorithm (cf.
+Interestingly, the conditions for the existence of a gradient-feedback algorithm and those for a parameter-feedback algorithm (cf.
 Section 3) are identical.
-This is not surprising, since a parameter-feedback algorithm is assumed to have access to $\theta(t)$ while a gradient-feedback algorithm needs to measure $\theta(t)$ indirectly through $y(t).$ More precisely, as stated by (27), the dynamic state of the controller $z(t)$ acts as an alternative representation (i.e., in different coordinates) of the exosystem state $\theta(t),$ while the exploration signal $x(t)$ acts as a parameter-feedback algorithm (see (26)).
+This should not be surprising: on the one hand, a parameter-feedback algorithm assumes access to $\theta(t)$; on the other hand, a gradient-feedback algorithm must infer $\theta(t)$ through $y(t)$, but (see (27)) the dynamic state of the controller $z(t)$ acts as a copy (possibly in different coordinates) of $\theta(t)$.
+In this sense, both algorithm classes rely on essentially the same information, and it is therefore natural that the conditions for their existence coincide.
 4.2 Design of dynamic gradient feedback algorithms
@@
 The algorithm uses a Luenberger observer to estimate the exosystem state $\theta(t)$ (cf.
-line $4$), and a parameter feedback algorithm is then applied to the estimated exosystem state to regulate the gradient to zero (precisely, $G_{c}(z)$ is designed following the approach of Theorem 1 – cf.
-line $3$ of Algorithm 1).
+line $4$), and a parameter feedback algorithm is applied to the estimated exosystem state to regulate the gradient to zero; precisely, $G_{c}(z)$ is designed following the approach of Theorem 1 (see line $3$ of Algorithm 1).
 Notice that the algorithm uses the particular choice for $\sigma(\theta)$ being the identity operator (see Remark 12).
@@
 1 $n_{c}\leftarrow p$;
-2 $L\leftarrow$ any matrix such that $S-LQ$ is Hurwitz;
+2 $L\leftarrow$ any matrix such that SLQS-LQ is Hurwitz;
 3 $G_{c}(z)\leftarrow H_{c}(z)$;
@@
 Example 3.
-Consider the quadratic problem studied in Example 2, and assume that the exosystem follows the linear model $\dot{\theta}=S\theta$ for some matrix $S\in{}^{p\times p}$ as in Assumption 2.
+Consider the quadratic problem studied in Example 2, and assume that the exosystem follows the linear model $\dot{\theta}=S\theta$ for some matrix $S\in\real^{p\times p}$ as in Assumption 2.
 Further, suppose that $Q$ is such that the unobservable subspace $\mathcal{N}=\cap_{i\geq 0}\ker(QS^{i})$ is the empty set $\mathcal{N}=\emptyset,$ so that Assumption 3 is satisfied; notice that this assumption holds, for example, when $n=p$ and $Q$ is invertible.
@@
 $\dot{z}=A_{c}z+B_{c}y,$ $x=G_{c}z,$ $y=Rx+Q\theta,$ (29)
-where $A\in{}^{n_{c}\times n_{c}},$ $B_{c}\in{}^{n_{c}\times n},$ $G_{c}\in{}^{n\times n_{c}},$ achieves asymptotic tracking if and only if there exists a linear transformation $\Sigma\in{}^{n_{c}\times p}$ such that:
+where $A\in\real^{n_{c}\times n_{c}},$ $B_{c}\in\real^{n_{c}\times n},$ $G_{c}\in\real^{n\times n_{c}},$ achieves asymptotic tracking if and only if there exists a linear transformation $\Sigma\in\real^{n_{c}\times p}$ such that:
 $\Sigma S$ $=A_{c}\Sigma,$ (30a)
-00 0 $=RG_{c}\Sigma+Q.$ (30b)
+$0$ $=RG_{c}\Sigma+Q.$ (30b)
 Next, we apply Algorithm 1 to design an algorithm that solves this problem.
 The mapping $H_{c}(\theta)=-R^{\dagger}Q\theta$ zeros the gradient (see Example 2).
-Following Algorithm 1, we set $n_{c}=p$, choose matrix $L$ such that $S-LQ$ is Hurwitz, define $G_{c}(z)=H_{c}(z)=-R^{\dagger}Qz$, and then select $A_{c}$ and $B_{c}$ such that $F_{c}(z,y)=A_{c}z+B_{c}y$ satisfies:
+Following Algorithm 1, we set $n_{c}=p$, choose matrix $L$ such that SLQS-LQ is Hurwitz, define $G_{c}(z)=H_{c}(z)=-R^{\dagger}Qz$, and then select $A_{c}$ and $B_{c}$ such that $F_{c}(z,y)=A_{c}z+B_{c}y$ satisfies:
 $F_{c}(z,y)$ $=s(z)+L(y-\nabla_{x}f(H_{c}(z),z)),$
@@
 $=Sz+Ly,$
-which gives $A_{c}=S$ and $B_{c}=L.$ In summary, a dynamic gradient-feedback algorithm is (29) with $A_{c}=S$, $B_{c}=L$, and $G_{c}=-R^{\dagger}Q$, where $L$ is any matrix such that $S-LQ$ is Hurwitz.
+which gives $A_{c}=S$ and $B_{c}=L.$ In summary, a dynamic gradient-feedback algorithm is (29) with $A_{c}=S$, $B_{c}=L$, and $G_{c}=-R^{\dagger}Q$, where $L$ is any matrix such that SLQS-LQ is Hurwitz.
 With this choice, conditions (30) are satisfied with $\Sigma=I$ (see Remark 12).
@@
 4.3 The optimization algorithm as a copy of the exosystem
-As shown by (27), the state of optimization algorithms that achieve asymptotic tracking and that of the exosystem are related by a change of coordinates.
-In this section, we show that $\sigma(\theta)$ is an injective map and thus is a well-posed change of coordinates.
-For Algorithm 1, we have the following.
+As shown in the previous section, the state of the optimization algorithm and that of the exosystem must be related, at all points of the limit set, by (27).
+In this section, we show that $\sigma(\theta)$ is injective, and therefore defines a well-posed change of coordinates.
+We begin by establishing the following result.
 Proposition 7 (Injectivity of $\sigma(\theta)$ from Algorithm 1).
@@
 Proposition 8 (Injectivity of any $\sigma(\theta)$).
-Let the assumptions of Theorem 5 hold and assume that $s(\theta)$ is $C^{\infty}.$ For any $F_{c}(z,y)$ and $G_{c}(z)$ that achieve asymptotic tracking, there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin such that $\sigma(\theta)$ is injective in $\Theta_{\circ}$$\square$
+Let the assumptions of Theorem 5 hold, and assume that $s(\theta)$ is of class $C^{\infty}$.
+Then, for any algorithm of the form (10) that achieves asymptotic tracking, there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin such that $\sigma(\theta)$ is injective on $\Theta_{\circ}\square$
 Proof.
@@
 $\begin{bmatrix}L_{s}(\nabla_{x}f)(x,\theta)\\ L_{s}(L_{s}(\nabla_{x}f))(x,\theta)\\ \vdots\end{bmatrix}$,
-is not invertible at $\theta=\theta^{\prime}.$ By [53, Thm. 3.13], the system (2) is not weakly observable, thus violating Assumption 3. ∎
+is not invertible at $\theta=\theta^{\prime}.$ By [54, Thm. 3.13], the system (2) is not weakly observable, thus violating Assumption 3. ∎
 5 Fidelity of the internal model vs accuracy
-In Section 4, we have shown that a necessary condition for exact asymptotic tracking is that the algorithm incorporates an internal model of the temporal variability of the problem.
+In Section 4, we showed that a necessary condition for exact asymptotic tracking is that the algorithm incorporates an internal model of the temporal variability of the problem.
 Departing from this finding, in this section we tackle the following question: how will the algorithm perform when the temporal variability of the problem is known only approximately?
 For illustration purposes, we focus on the quadratic problem (19) (see Examples 2 and 3).
-Suppose that an imprecise internal model is available; that is, there exists $\Delta\in{}^{n_{c}\times p}$ such that (30) is modified to777 To explicitly model uncertainty in the matrix $S,$ equation (32) could be rewritten as $\Sigma(S+\Delta_{s})=A_{c}\Sigma.$ By comparison, the two approaches are equivalent with the substitution $\Delta=\Sigma\Delta_{s}.$ :
+Suppose that an imprecise internal model is available; that is, there exists $\Delta\in\real^{n_{c}\times p}$ such that (30) is modified to [^7] To explicitly model uncertainty in the matrix $S,$ equation (32) could be rewritten as $\Sigma(S+\Delta_{s})=A_{c}\Sigma.$ By comparison, the two approaches are equivalent with the substitution $\Delta=\Sigma\Delta_{s}.$ :
 $\Sigma S+\Delta=A_{c}\Sigma\qquad\text{and}\qquad 0=RG_{c}\Sigma+Q.$ (32)
@@
 $L(x,\lambda,\theta(t))=f(x,\theta(t))+\sum_{i=1}^{m}\lambda_{i}h_{i}(x,\theta(t)),$
-[…]
+where $\lambda_{i}$ is the Lagrange multiplier associated with the $i\textsuperscript{th}$ equality constraint.
+A pair $(x,\lambda)\in\real^{n}\times\real^{m}$ is said to be a saddle-point of the Lagrangian if, for all $(\bar{x},\bar{\lambda})\in\real^{n}\times\real^{m}$,
 $L(x,\bar{\lambda},\theta(t))\leq L(x,\lambda,\theta(t))\leq L(\bar{x},\lambda,\theta(t)).$
@@
 It follows from the derivations in the previous sections that the gradient-feedback and parameter-feedback algorithms can be directly applied to seek a stationary point of the Lagrangian function by replacing (in (1)) the variable $x$ with the extended decision variable $\tilde{x}=(x,\lambda)$ and by letting $f(\tilde{x},\theta)=L(x,\lambda,\theta).$ Notice that, if the critical point computed by (10b) is also a saddle-point and strong duality holds, then it is also a solution to the equality-constrained problem.
-When strong duality does not hold, however, such a saddle-point may not correspond to an optimizer [54, Ch.
+When strong duality does not hold, a saddle-point may not correspond to an optimizer [55, Ch.
 5].
 6.2 Discrete-time algorithms
-In practice, optimization algorithms are often implemented on digital machines and thus, in these cases, implementing (10) requires discretizations.
+In practice, optimization algorithms are often implemented on digital machines and thus require discretization of (10).
+It is worth noting that this scenario corresponds to designing a discrete-time algorithm to track a continuous-time optimizer (cf.
+(1) and (5)).
 In these cases, a possible solution is to pursue a design by emulation [56], which consists in converting the continuous-time algorithm (10) into a discrete-time one by approximation (that is, by approximating the time-derivative by a discrete difference using, e.g., Backward/Forward Difference, Tustin, ZOH methods, etc.).
-Then, the continuous-time guarantees given here can be translated into discrete-time ones (using standard tools from emulation [55] or sampled-data systems [56]), provided that the sampling time is sufficiently small.
-In these cases, exact tracking established here will continue to hold at the discrete time instants, while tracking will be inexact in-between sampling times and the error can be bounded in terms of the sampling frequency using the explicit bounds in [55].
-An alternative method that could be pursued is that of discretizing the temporal variability $\theta(t)$, thus leading to a discrete-time optimization problem as in (4).
-For this alternative formulation, the critical trajectories are discrete signals and, in this case, it would be natural to seek a fully-discrete alternative to (10) and pursue an analysis in discrete time.
-Such a design is beyond the scope of this work and will be left as the focus of future works.
+Then, the continuous-time guarantees given here can be translated into discrete-time ones using standard tools from emulation [56], provided that the sampling time is sufficiently small.
+In these cases, exact tracking established above will hold at the discrete time instants, while tracking will be inexact in-between sampling times and the error can be bounded in terms of the sampling frequency using the explicit bounds in [56].
+An alternative approach that could be pursued is that of discretizing the temporal variability $\theta(t)$, thus leading to a discrete-time optimization problem as in (4).
+For this alternative formulation, the critical trajectories are instead discrete signals and, in this case, it would be natural to seek a discrete algorithm, in place of (10), and pursue an analysis in discrete time.
+Such a design is beyond the scope of this work and is left as the focus of future works.
 7 Simulation results
@@
 4(a), we show the error for our algorithm with observer pole locations $(-\eta,-2\eta)$ for $\eta\in\{1,5,20\}$ (shown in dark to light blue) along with the corresponding rate $e^{-\eta t}$ (dashed black).
-For comparison, we also show the error of the prediction-correction (PC) algorithm (8) along with two of its finite-time (FT) variants [34, Equations (9) and (20)] using the parameter values in [34, Section VI.A].
+For comparison, we also show the error of the prediction-correction (PC) algorithm (8) along with two of its finite-time (FT) variants [35, Equations (9) and (20)] using the parameter values in [35, Section VI.A].
 Instead of computing $H_{c}(\theta)$ numerically, we can use its linear approximation in (20).
@@
 The bottom figure shows that, from a zero initial condition, the algorithm is capable of computing a Wardrop equilibrium in about $1.5$ hours, and then of tracking this equilibrium.
-Notice that, a faster rate of convergence could be obtained by shifting the eigenvalues of $S-LQ.$
+Notice that, a faster rate of convergence could be obtained by shifting the eigenvalues of SLQ.S-LQ.
 We next illustrate the applicability of the framework in solving the dynamic traffic assignment problem in roadway transportation [58]; intuitively, the objective is to decide how traffic flows are split among the available paths of a network to minimize the drivers’ travel time to destination.
@@
 Mathematically, Wardrop’s equilibrium is the optimizer of the following optimization problem:
-[…]
+$\min_{x\in\real^{|\mathcal{E}|}}$~~~ $~~~\sum_{i\in\mathcal{E}}\int_{0}^{x_{i}}\ell_{i}(s)ds$
 subject to: $\sum_{j\in\mathcal{E}:j^{-}=v}x_{j}-\sum_{j\in\mathcal{E}:j^{+}=v}x_{j}=\delta_{v}(\theta(t)),\ \forall v\in\mathcal{V},$
@@
 $S=\text{diag}\biggl(\begin{bmatrix}0&1\\ -\omega_{1}^{2}&0\end{bmatrix},\begin{bmatrix}0&1\\ -\omega_{2}^{2}&0\end{bmatrix},0\biggr).$
-Notice that knowledge of $\theta_{0},\theta_{1},\theta_{2},\phi_{1},\phi_{2}$ is not required to synthesize the optimization algorithm – only the frequencies $\omega_{1},\omega_{2}$ are required to be known.
+Notice that knowledge of $\theta_{0},\theta_{1},\theta_{2},\phi_{1},\phi_{2}$ is not required to synthesize the optimization algorithm—only the frequencies $\omega_{1},\omega_{2}$ are required to be known.
 To seek a solution to the constrained problem (7.3), consider the Lagrangian function:
-[…]
+$L(x,\lambda,\theta(t)):=\sum_{i\in\mathcal{E}}\int_{0}^{x_{i}}\ell_{i}(s)ds\\ +\sum_{v\in\mathcal{V}}\lambda_{v}\biggl(\delta_{v}(\theta(t))-\sum_{j\in\mathcal{E}:j^{-}=v}x_{j}+\sum_{j\in\mathcal{E}:j^{+}=v}x_{j}\biggr),$
 where $\lambda:=(\lambda_{1},\dots,\lambda_{|\mathcal{V}|})$ is the vector of Lagrange multipliers.
 We applied Algorithm 1 to the optimization problem (1) with $f(\tilde{x},\theta)=L(x,\lambda,\theta)$ (see Section 6); the inequality constraints in (7.3) have been accounted for by projecting $x(t)$ onto the feasible set.
-Here, matrix $L$ has been chosen so that the eigenvalues of $S-LQ$ are uniformly distributed in the open real interval $(-1,-2)$.
+Here, matrix $L$ has been chosen so that the eigenvalues of SLQS-LQ are uniformly distributed in the open real interval (1,2)(-1,-2).
 Notice that, since the latencies are strictly increasing, the Lagrangian is strongly convex-strongly concave [58], and thus the problem admits a unique critical point that is a saddle point.
@@
 We showed that the problem of designing algorithms for time-varying optimization problems can be reformulated as a nonlinear multivariate output regulation problem.
-This connection allowed us to prove the internal model principle of time-varying optimization, which states that (when the algorithm has access only to functional evaluations of the gradient) asymptotic tracking can be achieved only if the algorithm incorporates a reduplicated model of the temporal variability of the problem.
-On the other hand, when the time-varying parameters are measurable, asymptotic tracking can be achieved by seeking a mapping that zero the gradient.
+This connection allowed us to prove the internal model principle of time-varying optimization, which states that (when the algorithm has access only to function evaluations of the gradient) asymptotic tracking can be achieved only if the algorithm incorporates a reduplicated model of the temporal variability of the problem.
+On the other hand, when the time-varying parameters are measurable, asymptotic tracking can be achieved by seeking a mapping that zeros the gradient.
 Importantly, our results show that asymptotic tracking can be achieved under more relaxed assumptions than what is normally imposed in the literature.
@@
 We now summarize relevant facts in center manifold theory from [19]; see also [59]. Consider the nonlinear system:
-$\dot{x}=f(x)$ (37)
+$\dot{x}=f(x),$ (37)
 where $f$ is a $C^{k}$ vector field defined on an open subset $U$ of n, and let $x_{\circ}\in U$ be an equilibrium point for $f$, i.e., $f(x_{\circ})=0$.
@@
 Definition 4 (Locally invariant manifold).
-A $C^{k}$ manifold $S$ of $U$ is locally invariant for (37) if, for each $x_{\circ}\in S$, there exists $t_{1}<0<t_{2}$ such that the integral curve $x(t)$ of (37) satisfying $x(0)=x_{\circ}$ satisfiest $x(t)\in S$ for all $t\in(t_{1},t_{2})$.
+A $C^{k}$ manifold $S$ of $U$ is locally invariant for (37) if, for each $x_{\circ}\in S$, there exists $t_{1}<0<t_{2}$ such that the integral curve $x(t)$ of (37), such that $x(0)=x_{\circ}$, satisfies $x(t)\in S$ for all $t\in(t_{1},t_{2})$.
 $\square$
@@
 $\dot{y}=f_{y}(\theta,y),$ $\dot{\theta}=f_{\theta}(\theta,y),$ (38)
-a curve $y=\pi(\theta)$ is an invariant manifold for (38) if the solution of (38) with $\theta(0)=\theta_{\circ}$ and $y(0)=\pi(\theta_{\circ})$ lies on the curve $y=\pi(\theta)$ for $t$ in a neighborhood of $0$ The notion of invariant manifold is useful as, under certain assumptions, it allows us to reduce the analysis of (37) to the study of a reduced system in the variable $\theta$ only.
+a curve $y=\pi(\theta)$ is an invariant manifold for (38) if the solution of (38) with $\theta(0)=\theta_{\circ}$ and $y(0)=\pi(\theta_{\circ})$ lies on the curve $y=\pi(\theta)$ for $t$ in a neighborhood of $0.$ The notion of invariant manifold is useful as, under certain assumptions, it allows us to reduce the analysis of (37) to the study of a reduced system in the variable $\theta$ only.
 The remainder of this section is devoted to formalizing this fact.
@@
 Theorem 9 (Center manifold existence theorem).
-Assume that $n^{+}=0$. There exists a neighborhood $V\subset{}^{n^{0}}$ of 00 and a class $C^{k-1}$ mapping $\pi:V\rightarrow{}^{n^{-}}$ such that the set
-$S=\{(y,\theta)\in{}^{n^{-}}\times V:y=\pi(\theta)\},$
+Assume that $n^{+}=0$. There exists a neighborhood $V\subset\real^{n^{0}}$ of 0 and a class $C^{k-1}$ mapping $\pi:V\rightarrow\real^{n^{-}}$ such that the set
+$S=\{(y,\theta)\in\real^{n^{-}}\times V:y=\pi(\theta)\},$
 is a center manifold for (39). $\square$
@@
 There exists a neighborhood $U^{\circ}$ of $(0,0)$ and real numbers $M>0$ and $K>0$ such that, if $(y(0),\theta(0))\in U^{\circ}$, then for all $t\geq 0$,
-$\|y(t)-\pi(\theta(t))\|\leq Me^{-Kt}\|y(0)-\pi(\theta(0))\|.$square
+$\|y(t)-\pi(\theta(t))\|\leq Me^{-Kt}\|y(0)-\pi(\theta(0))\|.$ $\square$
 From the above discussion, any trajectory of (39) starting at a point $y^{\circ}=\pi\left(\theta^{\circ}\right)$ of a center manifold satisfies $y(t)=\pi(\zeta(t))$ and $\theta(t)=\zeta(t)$, where $\zeta(t)$ is any solution of